### v2

# Graded Projection Recursion (GPR): A Framework for Controlling Bit-Complexity of Algebraic Packing

Abstract
We present Graded Projection Recursion (GPR), a framework for converting certain blocked recursive algorithms into model-honest (i.e., reflects full bit complexity) near-quadratic procedures under bounded intermediate budgets. Part I gives a proof-complete integral specification of a three-band packing identity and a two-round middle-band extractor, and shows how per-node centering and sqrt-free dyadic $\ell_{2}$ normalization (recursive invariant amplification) gives a sufficient packing base $\beta(\sigma)$ that grows linearly with the scaling depth $\sigma$ (i.e., logarithmic bit-growth) in exact arithmetic.
On fixed-width hardware, exact evaluation of the packed recursion generally requires either an extended-precision path or digit-band (slice) staging that emulates $b(n)$ bits of mantissa precision using $w$-bit words, incurring only polylogarithmic overhead; This leads to a soft-quadratic bit cost $\tilde{O}(n^{2})$ (e.g., $\Theta(n^{2}\log^{2}n)$ when $b(n)=\Theta(\log n)$ in the basic $2\times 2$ recursion). Part II introduces execution-format comparators (e.g., IEEE-754), a drift ledger, and a decision-invariance theorem that supports commensurate-accuracy claims in floating arithmetic (and that cleanly accounts for any staged/truncated auxiliary drift). Part III provides case-study reductions (LUP/solve/det/inv, LDL, blocked QR, SOI/SPD functions, GSEVP, dense LP/SDP IPM kernels, Gaussian process regression, and representative semiring problems) showing how to export the kernel advantage without reintroducing uncontrolled intermediate growth. Part IV abstracts admissible packings and extractors via a master condition and an easily checkable BWBM sufficient condition, and sketches extensions to multilinear/multigraded kernels and non-rounding extractors (e.g., CRT and semiring bucket projections).
Keywords: Graded Projection Recursion (GPR). Bit-Complexity. Algebraic Packing. Matrix Multiplication. Uniform Bit Model. Word-RAM model. Divide-and-Conquer Algorithms. Commensurate-Accuracy.
1 Executive summary: why GPR avoids degree-mixing and bit-growth obstructions
Many packing-based schemes attempt to accelerate bilinear computations by embedding multiple quantities into a single multiplication, separating them by a large base parameter (e.g., $\beta$), and decoding a target coefficient/band. A recurring pitfall, visible in Winograd/APA-style analyses, is a degree-mixing obstruction: as recursion proceeds, terms intended to remain in well-separated bands interact, forcing either (1) depth-dependent bases $\beta(m)$ or (2) growing precision/word size, which can erase the putative speedup in any model-honest (bounded-budget) complexity measure.
GPR avoids this obstruction by combining two ideas:
  1. 1.
    Band extraction with slack. Packed products are designed to admit a stable three-band decomposition
    $Z_{v}~=~\beta U_{v}+T_{v}+L_{v}/\beta,$ (1)
    and a two-round extractor (Part I) recovers $T_{v}$ exactly whenever a gap condition holds with slack. Part IV abstracts this as an extractor whose correctness follows from a checkable sufficient condition (a bounded-width/bounded-mixing inequality) that is established by the per-node projection discipline (or verified via the master condition/BWBM in Part IV), rather than assumed as an input restriction. (We do not assume band boundedness as an input promise. We either prove it from the projection invariant (Part I and Part II) or state it as a verifiable admissibility condition (Part IV) that must be checked for a new kernel.)
  2. 2.
    Recursive invariant amplification (graded projection). At every recursion node, we re-establish a local invariant (a “centered manifold” and a normalization) so that: (1) the target and tail bands remain uniformly bounded relative to the classical block bound, and (2) the extractor’s sufficient condition holds uniformly across depth. In the MM instantiation, the projection is centering plus dyadic $\ell_{2}$ normalization, giving dimension-free dot-product bounds and a sufficient base $\beta$ that adapts to the lattice resolution.
The obstruction, made precise
In a generic bilinear recursion, a packed leaf computation produces many “leaf terms” corresponding to different ways of pairing subblocks. If these terms can drift between graded bands as they pass down the recursion, then the off-band magnitude cannot be bounded uniformly without increasing $\beta$ or precision with depth. This is the degree-mixing obstruction in model-honest settings.
The evasion mechanism: first-mismatch elimination
GPR enforces a compositional band structure: at every node, the only active degrees are confined to a constant-width window (e.g., $\{-1,0,1\}$ in the MM three-band design). In this setting, every leaf term has a well-defined first mismatch node: the first node where it pairs a “cross” block rather than a “matched” block for the target. The local packing identity forces such a first mismatch contribution into an off-band at that node, and the local extractor removes it once the gap condition holds. Consequently, only fully matched terms survive to the root. See Definition 5.4 and Corollary 5.5 for a formal statement of the mechanism: non-matched leaf terms are eliminated locally at their first mismatch node, preventing degree-mixing accumulation across depth.
Consequences
This perspective gives three immediate, reusable consequences:
  1. 1.
    Depth-uniform boundedness. If graded projection gives a depth-uniform bound on the target/tail bands (Part IV master condition or BWBM), then a choice of extractor parameters dependent only on depth (e.g., $\beta\propto 2^{\sigma}$) suffices.
  2. 2.
    Model-honest near-quadratic complexity. If graded projection gives depth-uniform extraction parameters and bounded intermediates, the recursion has the structural recurrence $4T(n/2)+O(n^{2})$ (one packed multiply per quadrant plus $O(n^{2})$ projection/packing work per level). In a unit-cost scalar model this gives $\Theta(n^{2}\log n)$ arithmetic work. In a fixed-word implementation, if packed intermediates require $b(n)$ bits of effective mantissa precision, then digit-band (slice) staging on $w$-bit words introduces an explicit factor $K(n):=\lceil b(n)/w\rceil$ (see Theorem 14.1), giving a soft-quadratic format/bit cost $\tilde{O}(n^{2})$ (typically $\Theta(n^{2}\log^{2}n)$ when $b(n)=\Theta(\log n)$ and $w$ is fixed).
  3. 3.
    Robustness and feasibility. Recursive normalization bounds intermediates throughout the recursion tree. This enlarges feasible regions under fixed intermediate budgets (and, in float mode, localizes drift to ledger-bounded decision sites).
[…]
Remark 1.2 (No-cheating model note).
We do not claim “magic compression” (i.e., exact unbounded-precision behavior inside a single fixed-width word). All principal claims are model-honest in one of the following regimes:
  1. 1.
    Fixed-format (float) regime. All scalars are executed in a fixed format $\mathsf{F}$ (e.g., IEEE-754) and claims are comparator-based: we prove commensurate-accuracy and decision invariance under a drift ledger (Part II).
  2. 2.
    Staged-precision regime. When a target accuracy or exactness requirement demands $b(n)$ bits of effective mantissa precision, we emulate those bits on fixed-width hardware via digit-band (slice) staging. The resulting overhead is an explicit polylogarithmic factor (e.g., $K(n):=\lceil b(n)/w\rceil$ on $w$-bit words), and the associated truncation effects are either eliminated (exact staged mode) or budgeted explicitly as $\varepsilon_{\mathrm{aux}}$ in the drift ledger (float staged mode).
Instances that overflow the declared exponent/word budgets leave the stated success regime, exactly as in classical implementations.
How to read the rest
  • Part I proves the mechanism end-to-end for matrix multiplication in exact dyadic lattice semantics.
  • Part II formulates baseline comparators and drift-ledger conditions for float-GPR.
  • Part III gives case studies showing how GPR-MM kernels export to other algorithms.
  • Part IV abstracts admissibility conditions (master condition/BWBM) and extractor generalizations.
Part I Integral kernel (exact dyadic lattice semantics)
2 Scope and design goals
Definition 2.1 (Near-quadratic).
For our purposes, near-quadratic means $O(n^{2}\log n)$ unless stated otherwise. The $\log n$ factor comes from recursion depth in the $2\times 2$ blocked recursion. All other per-level overheads are $O(n^{2})$.
2.1 What the kernel does
  • Exact specification of $C=AB$ in dyadic lattice semantics, with exact realization obtained either by staged precision (slice emulation) or, in fixed-format mode, by commensurate-accuracy execution (Part II).
  • Near-quadratic arithmetic recurrence $4T(n/2)+O(n^{2})$.
  • Eliminate bit-growth using recursive invariant amplification (centering + dyadic $\ell_{2}$ normalization).
2.2 What is intentionally deferred
  • Float/commensurate-accuracy comparator model (Part II).
  • Semiring algorithms and graph/language case studies (Part III).
2.3 Proof dependency map (Part I)
The proof spine for the integral kernel has the following dependency chain:
  1. 1.
    Rounding primitives:
    • Lemma 4.1 (rounding stability) and
    • Lemma 4.2 (exactness on $D^{-2}\mathbb{Z}$).
  2. 2.
    Extraction:
    • Definition 5.1 (sieve),
    • Lemma 5.2 (sieve exactness),
    • Lemma 5.3 (master gap).
  3. 3.
    Packing:
    • Definition 5.6 (quadrant packing),
    • Proposition 5.7 (three-band identity).
  4. 4.
    Normalization (recursive invariant amplification):
    • Definitions 6.16.2 and
    • Proposition 6.3 (dyadic $\ell_{2}$ scaling),
    • Lemma 6.5 (dimension-free dot products).
  5. 5.
    Sufficient base:
    • Theorem 7.1.
  6. 6.
    Reconstruction and correctness:
    • Lemma 8.1 (centering identity),
    • Theorem 8.5 (kernel correctness).
  7. 7.
    Complexity:
    • Theorem 9.2 and
    • Theorem 9.3.
3 Execution model and representations
3.1 Dyadic lattice and scaled coefficient lattice
Definition 3.1 (Dyadic lattice).
Fix $D\in\mathbb{N}$. Define $\mathcal{L}_{D}:=D^{-1}\mathbb{Z}[1/2]$. All matrices in Part I are over $\mathcal{L}_{D}$ entrywise.
Definition 3.2 (Scaled coefficient lattice).
Recursive invariant amplification introduces dyadic scalings at each node. To maintain exactness at a recursion node with cumulative scaling exponent $\sigma\in\mathbb{Z}_{\geq 0}$, we define the local scaled coefficient lattice:
$\mathcal{L}_{\mathrm{coeff}}(\sigma)~:=~D^{-2}2^{-\sigma}\mathbb{Z}.$ (2)
The rounding operator relative to this lattice is defined via the lifted integers:
$\frac{1}{D^{2}2^{\sigma}}\left\lfloor D^{2}2^{\sigma}z\right\rceil~:=~\frac{1}{D^{2}2^{\sigma}}\left\lfloor D^{2}2^{\sigma}z\right\rceil.$ (3)
If $z\in\mathcal{L}_{\mathrm{coeff}}(\sigma)$, then $\frac{1}{D^{2}2^{\sigma}}\left\lfloor D^{2}2^{\sigma}z\right\rceil=z$ exactly.
Definition 3.3 (Digit-Band Staging).
In the fixed-word model, elements of $\mathcal{L}_{\mathrm{coeff}}(\sigma)$ are represented by a tuple of $K(n)\propto\sigma$ fixed-width words (digits) and a shared exponent offset. This allows exact representation of the lattice sites despite the dynamic scaling, incurring the polylogarithmic overhead described in Theorem 9.3.
3.2 Exact centering model
Definition 3.4 (Exact dyadic means).
At a recursion node of dimension $m=2^{t}$, define means by dyadic division:
$\mu~:=~\frac{1}{m}\sum_{i=1}^{m}a_{i}.$ (4)
Because $\mathcal{L}_{D}=D^{-1}\mathbb{Z}[1/2]$ is closed under division by powers of two, $\mu\in\mathcal{L}_{D}$, and the division by $m$ is implemented exactly as an exponent-offset update.
4 Rounding primitives
Lemma 4.1 (Margin-based rounding stability).
Let $\left\lfloor\cdot\right\rceil:\mathbb{R}\to\mathbb{Z}$ denote rounding to the nearest integer with any fixed deterministic tie-breaking convention. Fix $m\in\mathbb{Z}$ and $\delta\in(0,\frac{1}{2}]$. If $r,e\in\mathbb{R}$ satisfy:
$\lvert r-m\rvert \displaystyle\leq$ $\frac{1}{2}-\delta$ (5)
$\lvert e\rvert$ $<$ $\delta,$ (6)
then $\left\lfloor r\right\rceil=m$ and $\left\lfloor r+e\right\rceil=m$. In particular, $\left\lfloor r+e\right\rceil=\left\lfloor r\right\rceil$.
Proof.
From $\lvert r-m\rvert\leq\frac{1}{2}-\delta$ we have
$m-\left(\frac{1}{2}-\delta\right)~\leq~r~\leq~m+\left(\frac{1}{2}-\delta\right),$ (7)
so $r$ lies in the closed interval $m-\frac{1}{2}+\delta,m+\frac{1}{2}-\delta]$. Because $\lvert e\rvert<\delta$, we obtain:
$m-\frac{1}{2}~<~r+e~<~m+\frac{1}{2}.$ (8)
Therefore $r+e$ lies strictly between the two half-integer boundaries $m\pm\frac{1}{2}$, so its nearest integer is uniquely $m$, independent of tie-breaking, so $\left\lfloor r+e\right\rceil=m$. Similarly, because $r\in[m-\frac{1}{2}+\delta$, $m+\frac{1}{2}-\delta]\subset(m-\frac{1}{2}$, $m+\frac{1}{2})$, the nearest integer to $r$ is also uniquely $m$, so $\left\lfloor r\right\rceil=m$. ∎
Lemma 4.2 (Exactness on the coefficient lattice).
Let $\frac{1}{D^{2}}\left\lfloor D^{2}z\right\rceil:=\frac{1}{D^{2}}\left\lfloor D^{2}z\right\rceil$ be rounding to the coefficient lattice $D^{-2}\mathbb{Z}$. If $z\in D^{-2}\mathbb{Z}$, then $\frac{1}{D^{2}}\left\lfloor D^{2}z\right\rceil=z$.
Proof.
Because $z\in D^{-2}\mathbb{Z}$, we can write $z=t/D^{2}$ for some $t\in\mathbb{Z}$, then $D^{2}z=t$ is an integer, so $\left\lfloor D^{2}z\right\rceil=\left\lfloor t\right\rceil=t$. Therefore:
$\frac{1}{D^{2}}\left\lfloor D^{2}z\right\rceil~=~\frac{1}{D^{2}}\left\lfloor D^{2}z\right\rceil~=~\frac{1}{D^{2}}t~=~z.$ (9)
5 Packing and extraction
5.1 Two-round sieve on scaled lattices
Definition 5.1 (Scaled Sieve).
Fix integer base $\beta>1$ and scaling exponent $\sigma\geq 0$. Define the scaled sieve operator:
$\mathcal{S}_{\beta,\sigma}(z)~:=~\frac{1}{D^{2}2^{\sigma}}\left\lfloor D^{2}2^{\sigma}z\right\rceil-\beta\frac{1}{D^{2}2^{\sigma}}\left\lfloor D^{2}2^{\sigma}z/\beta\right\rceil.$ (10)
Lemma 5.2 (Sieve exactness on scaled lattice).
Fix base $\beta\geq 2$ and scaling $\sigma$ and let:
$z~=~\beta U+T+\frac{L}{\beta}$ (11)
with $U,T\in\mathcal{L}_{\mathrm{coeff}}(\sigma)$ and $L\in\mathbb{R}$. Assume the gap inequalities hold relative to the lattice resolution:
$\lvert D^{2}2^{\sigma}L/\beta\rvert$ $<$ $\frac{1}{2},$ (12)
$\lvert D^{2}2^{\sigma}(T/\beta+L/\beta^{2})\rvert$ $<$ $\frac{1}{2},$ (13)
then $\mathcal{S}_{\beta,\sigma}(z)=T$.
Proof.
Multiply the identity for $z$ by the scaling factor $M:=D^{2}2^{\sigma}$. Because $U,T\in\mathcal{L}_{\mathrm{coeff}}(\sigma)$, the quantities $MU$ and $MT$ are integers:
$Mz~=~\beta(MU)+(MT)+\frac{ML}{\beta}.$ (14)
The gap inequalities imply the noise terms are strictly bounded by $1/2$. Thus, nearest-integer rounding gives:
$\left\lfloor Mz/\beta\right\rceil~=~MU\quad\text{and}\quad\left\lfloor Mz\right\rceil~=~\beta(MU)+MT.$ (15)
Dividing by $M$ recovers $U$ and $\beta U+T$. The sieve subtraction $(\beta U+T)-\beta U$ gives $T$ exactly. ∎
Lemma 5.3 (Master gap condition (scaled)).
If the single inequality
$D^{2}2^{\sigma}\left(\lvert L/\beta\rvert+\lvert T/\beta\rvert+\lvert L/\beta^{2}\rvert\right)~\leq~\frac{1}{2}-\delta$ (16)
holds, then the gap conditions of Lemma 5.2 hold with slack $\delta$, and $\mathcal{S}_{\beta,\sigma}(z)=T$.
Proof.
Immediate by applying the triangle inequality to the gap conditions. ∎
Definition 5.4 (First mismatch node (MM leaf term)).
Consider expanding the root output entry of the recursive $2\times 2$ blocked multiplication into a sum of leaf terms (indexed by choices of block indices at each recursion level). A leaf term is fully matched if at every recursion node it uses the matched pair contributing to the middle band (i.e., it contributes to the true quadrant product). Otherwise, it has a first mismatch node: the highest node on its root-to-leaf path at which it first uses a cross pair (i.e., is forced into an off-band by the local packing identity).
Corollary 5.5 (First-mismatch elimination at the first mismatch node).
Assume at every recursion node the local three-band identity of Proposition 5.7 holds and the sieve extractor returns the middle band exactly whenever the master gap condition holds (Lemmas 5.3 and 5.2), then any leaf term that is not fully matched is eliminated at its first mismatch node and does not contribute to the root output.
Proof.
At the first mismatch node, the local packing identity forces that term into one of the off-bands (the $\beta^{+1}$ or $\beta^{-1}$ band), so it does not appear in the middle band coefficient $T$ for that node. Exact local extraction returns only the middle band and discards off-band contributions. Therefore that term is eliminated at the first mismatch node and cannot propagate upward to the root. ∎
5.2 Quadrant packing identity
Definition 5.6 (Quadrant packing).
Let $A,B\in\mathcal{L}_{D}^{n\times n}$ with $n$ even, and partition into $2\times 2$ blocks
$A$ $=$ $\begin{pmatrix}A_{11}&A_{12}\\ A_{21}&A_{22}\end{pmatrix}$, (17)
$B$ $=$ $\begin{pmatrix}B_{11}&B_{12}\\ B_{21}&B_{22}\end{pmatrix}$, (18)
where each block is $(n/2)\times(n/2)$. Fix an integer base $\beta>1$. For each quadrant $(i,j)\in\{1,2\}^{2}$, define packed blocks:
$X_{ij}$ $:=$ $A_{i1}+\beta^{-1}A_{i2},$ (19)
$Y_{ij}$ $:=$ $B_{1j}+\beta B_{2j}.$ (20)
Proposition 5.7 (Three-band decomposition and band dictionary).
Let $A,B\in\mathcal{L}_{D}^{n\times n}$ be partitioned into $2\times 2$ blocks, and fix an integer base $\beta>1$. For each $(i,j)\in\{1,2\}^{2}$, define $(X_{ij},Y_{ij})$ by Quadrant packing (Definition 5.6), then:
$X_{ij}Y_{ij}~=~\beta G_{ij}+T_{ij}+\frac{1}{\beta}L_{ij},$ (21)
where the bands are:
$T_{ij}$ $:=$ $A_{i1}B_{1j}+A_{i2}B_{2j}~=~(AB)_{ij},$ (22)
$G_{ij}$ $:=$ $A_{i1}B_{2j},$ (23)
$L_{ij}$ $:=$ $A_{i2}B_{1j}.$ (24)
Proof.
By definition:
$X_{ij}$ $=$ $A_{i1}+\beta^{-1}A_{i2},$ (25)
$Y_{ij}$ $=$ $B_{1j}+\beta B_{2j}.$ (26)
Expand distributively:
$X_{ij}Y_{ij}$ $=$ $(A_{i1}+\beta^{-1}A_{i2})(B_{1j}+\beta B_{2j})$ (27)
$=$ $A_{i1}B_{1j}+\beta A_{i1}B_{2j}+\beta^{-1}A_{i2}B_{1j}+A_{i2}B_{2j}.$ (28)
Grouping terms by powers of $\beta$ gives:
$X_{ij}Y_{ij}~=~\beta(A_{i1}B_{2j})+\left(A_{i1}B_{1j}+A_{i2}B_{2j}\right)+\beta^{-1}(A_{i2}B_{1j}),$ (29)
which matches the claimed $\beta G_{ij}+T_{ij}+L_{ij}/\beta$ form. Finally, the middle band is exactly the $(i,j)$ block of $AB$ by standard block multiplication. ∎
Remark 5.8 (Why degree-mixing does not accumulate).
A leaf term that first uses a cross pair at some node is forced into an off-band at that node and is removed by exact local extraction (Corollary 5.5).
5.3 Worked micro-example (sanity check)
This example illustrates (1) three-band separation and (2) exact sieve extraction on a single scalar entry.
Scalar three-band and extraction.
Take $D=1$, choose $\delta=\frac{1}{8}$, and let $\beta=32$. Let $U=7$, $T=11$, L=9L=-9, and form:
$z~=~\beta U+T+L/\beta,$ (30)
then:
$\left|\frac{L}{\beta}\right|~=~\frac{9}{32}<\frac{1}{2},$ (31)
$\left|\frac{T}{\beta}+\frac{L}{\beta^{2}}\right|~\leq~\frac{11}{32}+\frac{9}{32^{2}}~<~\frac{1}{2},$ (32)
so Lemma 5.2 applies and $\mathcal{S}_{\beta}(z)=T$.
One quadrant packing identity.
For the $(1,1)$ quadrant, define $X_{11}=A_{11}+\beta^{-1}A_{12}$ and $Y_{11}=B_{11}+\beta B_{21}$, then Proposition 5.7 gives:
$X_{11}Y_{11}~=~\beta(A_{11}B_{21})+\underbrace{(A_{11}B_{11}+A_{12}B_{21})}_{(AB)_{11}}+\beta^{-1}(A_{12}B_{11}),$ (33)
so extracting the middle band recovers $(AB)_{11}$ exactly.
6 Recursive invariant amplification
6.1 Dyadic $\ell_{2}$ normalization without square roots
Definition 6.1 (Lifted squared norm).
Let $x\in\mathcal{L}_{D}^{m}$ be a real vector whose entries lie in the dyadic lattice $\mathcal{L}_{D}$. Define its lifted squared norm by:
$S(x)~:=~\sum_{r=1}^{m}\left(Dx_{r}\right)^{2}\in\mathbb{Z}_{\geq 0}.$ (34)
Equivalently, if $x_{r}=(w_{r}/D)2^{e_{r}}$ with $w_{r}\in\mathbb{Z}$, then $Dx_{r}=w_{r}2^{e_{r}}$ is dyadic and $S(x)$ is the exact squared $\ell_{2}$ norm of the lifted vector $Dx$.
Definition 6.2 (Dyadic scaling exponent).
Let $x\in\mathcal{L}_{D}^{m}$ and let $S(x)$ be its lifted squared norm from Definition 6.1. Define the dyadic scaling exponent by:
$k(x)~:=~\min\left\{k\in\mathbb{Z}_{\geq 0}:S(x)~\leq~2^{2k}\right\},$ (35)
with the convention $k(0)=0$ (equivalently, $S(0)=0$ so $k(0)=0$ is the minimum). Define the normalized vector:
$x^{\prime}~:=~2^{-k(x)}x.$ (36)
Proposition 6.3 (Implementability of $k(x)$).
Let $x\in\mathcal{L}_{D}^{m}$ and define $S(x)\in\mathbb{Z}_{\geq 0}$ and $k(x)\in\mathbb{Z}_{\geq 0}$ as in Definitions 6.16.2, then:
  1. 1.
    (Deterministic $\ell_{2}$ bound.) With $x^{\prime}:=2^{-k(x)}x$, we have $\lVert Dx^{\prime}\rVert_{2}^{2}\leq 1$, and therefore $\lVert x^{\prime}\rVert_{2}\leq 1$.
  2. 2.
    (No square roots needed.) If $S(x)>0$ and $\ell:=\lfloor\log_{2}S(x)\rfloor$ (bitlength minus one), let $k_{0}:=\lfloor\ell/2\rfloor$, then $k(x)$ is either $k_{0}$ or $k_{0}+1$, i.e.:
    $k(x)~=~\begin{cases}k_{0},&\text{if}~S(x)\leq 2^{2k_{0}},\\ k_{0}+1,&\text{otherwise.}\end{cases}$ (37)
    Equivalently, $k(x)=\left\lceil\frac{1}{2}\log_{2}S(x)\right\rceil$ with the convention $k(0)=0$.
  3. 3.
    (Offset-shift scaling.) Because $\mathcal{L}_{D}=D^{-1}\mathbb{Z}[1/2]$ is closed under multiplication by powers of two, the map $x\mapsto 2^{\pm k}x$ is an exact exponent-offset update in dyadic semantics (no mantissa growth is required).
Remark 6.4 (Computing $k(x)$ is a fixed-width operation in the intended model).
In integral mode, the squared-norm reduction is computed in a fixed-width accumulator (or under a fixed staging budget) and $k(x)$ is obtained from its bitlength. In float mode, $k(x)$ can be computed from the exponent field of a normal-range float. In either case this is $O(1)$ per row/column in the fixed-precision execution model.
Proof.
  1. 1.
    By Definition 6.1, $S(x)=\lVert Dx\rVert_{2}^{2}$. By Definition 6.2, $S(x)\leq 2^{2k(x)}$, so:
    $\lVert D(2^{-k(x)}x)\rVert_{2}^{2}~=~2^{-2k(x)}\lVert Dx\rVert_{2}^{2}~=~2^{-2k(x)}S(x)\leq 1.$ (38)
    Because $\lVert x^{\prime}\rVert_{2}=\frac{1}{D}\lVert Dx^{\prime}\rVert_{2}$, we also have $\lVert x^{\prime}\rVert_{2}\leq 1$.
  2. 2.
    If $S(x)=0$, then $x=0$ in the lifted coordinates and $k(x)=0$ by convention. Assume $S(x)>0$ and let $\ell=\lfloor\log_{2}S(x)\rfloor$, then:
    $2^{\ell}~\leq~S(x)~<~2^{\ell+1}.$ (39)
    Letting $k_{0}=\lfloor\ell/2\rfloor$ implies $2k_{0}\leq\ell<2k_{0}+2$, so:
    $2^{2k_{0}}~\leq~2^{\ell}\leq S(x)~<~2^{\ell+1}\leq 2^{2k_{0}+2}.$ (40)
    So $S(x)\in(2^{2k_{0}-1},2^{2k_{0}+2})$, and the smallest $k$ with $S(x)\leq 2^{2k}$ is either $k_{0}$ (if $S(x)\leq 2^{2k_{0}}$) or $k_{0}+1$ (otherwise). This is exactly the stated rule, and it matches $k(x)=\lceil\frac{1}{2}\log_{2}S(x)\rceil$.
  3. 3.
    By Definition 3.1 of $\mathcal{L}_{D}=D^{-1}\mathbb{Z}[1/2]$, multiplying any entry by $2^{\pm k}$ remains in $\mathcal{L}_{D}$ exactly.In the mantissa/exponent-offset representation, this is implemented by updating only the exponent offset by $\pm k$.
Lemma 6.5 (Dimension-free dot products).
Let $A^{\prime},B^{\prime}\in\mathcal{L}_{D}^{m\times m}$. Assume every row of $A^{\prime}$ has $\ell_{2}$ norm at most $1$ and every column of $B^{\prime}$ has $\ell_{2}$ norm at most $1$:
$\|A^{\prime}_{i,:}\|_{2}$ $\leq$ $1~\text{for all}~i,$ (41)
$\|B^{\prime}_{:,j}\|_{2}$ $\leq$ $1~\text{for all}~j,$ (42)
then every entry of the product $C^{\prime}:=A^{\prime}B^{\prime}$ satisfies:
$|C^{\prime}_{ij}|~\leq~1~\text{for all}~i,j.$ (43)
Proof.
Fix $i,j$. The $(i,j)$ entry is the dot product of the $i$-th row of $A^{\prime}$ and the $j$-th column of $B^{\prime}$:
$C^{\prime}_{ij}~=~\sum_{r=1}^{m}A^{\prime}_{ir}B^{\prime}_{rj}~=~\langle A^{\prime}_{i,:},B^{\prime}_{:,j}\rangle.$ (44)
By the Cauchy–Schwarz inequality:
$|C^{\prime}_{ij}|$ $=$ $|\langle A^{\prime}_{i,:},B^{\prime}_{:,j}\rangle|$ (45)
$\leq$ $\|A^{\prime}_{i,:}\|_{2}\|B^{\prime}_{:,j}\|_{2}$ (46)
$\leq$ $1\cdot 1~=~1.$ (47)
7 Sufficient-base theorem
Theorem 7.1 (Sufficient Base $\beta(\sigma)$).
Fix $\delta\in(0,\frac{1}{2}]$. Assume that at a recursion node with cumulative scaling $\sigma$, the normalized operands satisfy the dimension-free dot-product bounds $|T_{ij}|\leq 1$ and $|L_{ij}|\leq 1$.
If the packing base $\beta$ satisfies the resolution-dependent condition:
$D^{2}2^{\sigma}\left(\frac{2}{\beta}+\frac{1}{\beta^{2}}\right)~\leq~\frac{1}{2}-\delta,$ (48)
then the master gap condition (Lemma 5.3) holds, and $\mathcal{S}_{\beta,\sigma}(X_{ij}Y_{ij})=T_{ij}$ exactly.
Proof.
Substitute the bounds $|T_{ij}|\leq 1$ and $|L_{ij}|\leq 1$ into the master gap inequality (Eq 16). The LHS becomes bounded by $D^{2}2^{\sigma}(2/\beta+1/\beta^{2})$. If condition (48) holds, this upper bound is $\leq 1/2-\delta$, ensuring the gap is open and extraction is exact. ∎
Remark 7.2 (Why a growing base is not “cheating”).
Condition (48) implies that $\beta$ must grow roughly in proportion to $2^{\sigma}$. Because $\sigma$ represents the cumulative normalization depth, $\sigma\approx\log n$. Consequently, the required bit-width of $\beta$ grows logarithmically ($O(\log n)$). This is not a violation of the model; rather, it is the precise origin of the digit-band staging factor $K(n)$ (Theorem 9.3). The algorithm exchanges the uncontrolled magnitude growth of classical methods for a strictly controlled, logarithmic growth in precision (staging depth), giving the soft-quadratic complexity bound.
8 Kernel algorithm and correctness
8.1 Centering correction identity
Lemma 8.1 (Bilinear centering identity).
Let $A,B\in\mathbb{R}^{n\times n}$ and let $\alpha,\beta\in\mathbb{R}^{n}$, and define the centered matrices:
$A_{0}$ $:=$ $A-\mathbf{1}\alpha^{\top},$ (49)
$B_{0}$ $:=$ $B-\beta\mathbf{1}^{\top},$ (50)
where $\mathbf{1}\in\mathbb{R}^{n}$ is the all-ones vector, then:
$AB~=~A_{0}B_{0}+\mathbf{1}(\alpha^{\top}B_{0})+(A_{0}\beta)\mathbf{1}^{\top}+(\alpha^{\top}\beta)\mathbf{1}\mathbf{1}^{\top}.$ (51)
Proof.
From the definitions, $A=A_{0}+\mathbf{1}\alpha^{\top}$ and $B=B_{0}+\beta\mathbf{1}^{\top}$. Multiply and expand:
$AB$ $=$ $(A_{0}+\mathbf{1}\alpha^{\top})(B_{0}+\beta\mathbf{1}^{\top})$ (52)
$=$ $A_{0}B_{0}+A_{0}(\beta\mathbf{1}^{\top})+(\mathbf{1}\alpha^{\top})B_{0}+(\mathbf{1}\alpha^{\top})(\beta\mathbf{1}^{\top}).$ (53)
Now simplify each term:
$A_{0}(\beta\mathbf{1}^{\top})$ $=$ $(A_{0}\beta)\mathbf{1}^{\top},$ (54)
$(\mathbf{1}\alpha^{\top})B_{0}$ $=$ $\mathbf{1}(\alpha^{\top}B_{0}),$ (55)
and:
$(\mathbf{1}\alpha^{\top})(\beta\mathbf{1}^{\top})$ $=$ $\mathbf{1}(\alpha^{\top}\beta)\mathbf{1}^{\top}$ (56)
$=$ $(\alpha^{\top}\beta)\mathbf{1}\mathbf{1}^{\top}.$ (57)
Substituting gives the stated identity. ∎
Definition 8.2 (Normalization policy (row/column choice)).
At each recursion node, after centering we enforce the following $\ell_{2}$ bounds:
$\|A^{\prime}_{i,:}\|_{2}$ $\leq$ $1~\text{for all rows}~i,$ (58)
$\|B^{\prime}_{:,j}\|_{2}$ $\leq$ $1~\text{for all columns}~j.$ (59)
We refer to this as the row(A)-col(B) policy. It implies the entrywise bound $|(A^{\prime}B^{\prime})_{ij}|\leq 1$ by Cauchy–Schwarz, which is the invariant used to justify a constant packing base $\beta_{\star}$.
Remark 8.3 (What is centered vs what is normalized).
Centering enforces mean constraints. Normalization enforces norm constraints. These operations generally do not commute, and we cannot in general enforce zero-mean rows and columns and unit row and column norms simultaneously. Throughout, we use the row(A)–col(B) policy (Definition 8.2) specifically to guarantee the entrywise bound $|(A^{\prime}B^{\prime})_{ij}|\leq 1$ needed for a constant $\beta_{\star}$.
Mean constraints propagate only in the directions explicitly centered: if $B^{\prime}1=0$ (zero row means of $B^{\prime}$), then $(A^{\prime}B^{\prime})1=0$ (zero row means of the product). if $1^{\top}A^{\prime}=0$ (zero column means of $A^{\prime}$), then $1^{\top}(A^{\prime}B^{\prime})=0$ (zero column means of the product).
Normalization policy is fixed (do not over-constrain).
Throughout the manuscript we use the row(A)-col(B) normalization policy (Definition 8.2):
$\|A^{\prime}_{i,:}\|_{2}$ $\leq$ $1~\text{for all rows}~i,$ (60)
$\|B^{\prime}_{:,j}\|_{2}$ $\leq$ $1~\text{for all columns}~j.$ (61)
This choice is deliberate because it implies the entrywise bound $|(A^{\prime}B^{\prime})_{ij}|\leq 1$, which is the invariant needed to justify a constant packing base $\beta_{\star}$.
  • Do not attempt to enforce all row/column norm constraints simultaneously. In general you cannot simultaneously maintain (1) zero-mean rows and columns and (2) unit row and column norms for both operands, because row and column scalings interact and do not commute with centering. Implement exactly the policy above.
  • Centering is separate from normalization. Centering enforces mean constraints (row and/or column means) and is inverted by the bilinear identity (Lemma 8.1). Normalization enforces $\ell_{2}$ bounds and is inverted by applying the stored dyadic exponents, i.e., are treat as separate ordered steps: center first, then normalize.
  • What propagates. If centering enforces $B^{\prime}1=0$ (zero row means of $B^{\prime}$), then $(A^{\prime}B^{\prime})1=0$ (zero-mean rows propagate to the product). If centering enforces $1^{\top}A^{\prime}=0$ (zero column means of $A^{\prime}$), then $1^{\top}(A^{\prime}B^{\prime})=0$ (zero-mean columns propagate). The norm policy is chosen independently to guarantee bounded dot products.
In implementations, the dyadic exponents should be stored as two length-$n$ integer vectors (row exponents for $A^{\prime}$ and column exponents for $B^{\prime}$), and apply them only as exponent-offset updates (bit shifts), never as mantissa multiplications.
8.2 Implementation contract (integral kernel)
Reductions may use a wider accumulator.
The fixed-word model constrains the base word width for stored scalars (mantissas plus exponent offsets). In staged-precision regimes, one logical scalar may be represented by a small vector of fixed-width word “digits” (slice staging). Reductions such as means and squared norms may be computed using a wider accumulator type. Only the stored results must satisfy the fixed-word invariant.
The integral kernel’s guarantees rely on the following implementation contract:
  1. 1.
    Dyadic means are exact: Node dimensions are powers of two. Division by $m=2^{t}$ is implemented as an exponent-offset update.
  2. 2.
    No square roots: Dyadic $\ell_{2}$ normalization uses $S(x)=\sum(Dx_{r})^{2}$ and $k(x)$ via bitlength comparisons (Proposition 6.3). Scaling by $2^{\pm k}$ is an exponent-offset update.
  3. 3.
    Sufficient base: The packing base $\beta$ is chosen to satisfy Theorem 7.1 given the node’s scaling $\sigma$ (e.g., via the allocated staging budget).
  4. 4.
    Shifts do not touch mantissas: Multiplication by $2^{\pm t}$ and by $\beta_{\star}^{\pm 1}$ must be implemented as exponent updates, not as integer multiplication on mantissas.
  5. 5.
    Sieve sites are the only discrete extraction decisions: The only places where rounding/thresholding affects correctness are the sieve/extractor decision sites governed by Lemmas 5.2 and 5.3.
8.3 Kernel pseudocode
Input: $A,B\in\mathcal{L}_{D}^{n\times n}$, with $n$ a power of two
Output: $C=AB$ exactly
if $n=1$ then
 return ABAB
(1) Center. Compute the exact column-mean vector $\alpha\in\mathcal{L}_{D}^{n}$ of $A$ and exact row-mean vector $\beta\in\mathcal{L}_{D}^{n}$ of $B$ using dyadic division:
$\alpha_{j}$ $:=$ $\frac{1}{n}\sum_{i=1}^{n}A_{ij},$
$\beta_{i}$ $:=$ $\frac{1}{n}\sum_{j=1}^{n}B_{ij}.$
Set $A_{0}:=A-\mathbf{1}\alpha^{\top}$ and $B_{0}:=B-\beta\mathbf{1}^{\top}$.
(2) Dyadic $\ell_{2}$ normalize (sqrt-free). For each row $i$ of $A_{0}$, compute $k_{i}:=k(A_{0,i,:})$ via Definitions 6.16.2, and scale:
$A^{\prime}_{i,:}~:=~2^{-k_{i}}A_{0,i,:}.$
For each column $j$ of $B_{0}$, compute $\ell_{j}:=k(B_{0,:,j})$ and scale:
$B^{\prime}_{:,j}~:=~2^{-\ell_{j}}B_{0,:,j}.$
Store the exponents $\{k_{i}\}_{i=1}^{n}$ and $\{\ell_{j}\}_{j=1}^{n}$ for the undo step.
(3) Pack, recurse, and extract quadrants. Partition $A^{\prime},B^{\prime}$ into $2\times 2$ blocks of size $(n/2)\times(n/2)$. Choose base $\beta$ satisfying Theorem 7.1 given current scaling $\sigma$. foreach $(i,j)\in\{1,2\}^{2}$ do
   Form packed blocks (Definition 5.6):
$X_{ij}$ $:=$ $A^{\prime}_{i1}+\beta^{-1}A^{\prime}_{i2},$
$Y_{ij}$ $:=$ $B^{\prime}_{1j}+\beta B^{\prime}_{2j}.$
$Z_{ij}\leftarrow\textsc{KernelMM}(X_{ij},Y_{ij})$ $T_{ij}\leftarrow\mathcal{S}_{\beta,\sigma}(Z_{ij})$
 // Entrywise sieve (Definition 5.1)
Assemble $T$ from the four blocks $T_{ij}$.
(4) Undo normalization and un-center. Undo the dyadic scalings to recover $A_{0}B_{0}$ from $T$:
$(A_{0}B_{0})_{ij}~=~2^{k_{i}+\ell_{j}}T_{ij}~~~\text{(entrywise)},$
and then recover $C=AB$ using the bilinear centering identity (Lemma 8.1):
$C~:=~A_{0}B_{0}+\mathbf{1}(\alpha^{\top}B_{0})+(A_{0}\beta)\mathbf{1}^{\top}+(\alpha^{\top}\beta)\mathbf{1}\mathbf{1}^{\top}.$
return $C$
Algorithm 1 KernelMM(A,B)(A,B) (dyadic lattice semantics)
Remark 8.4 (Order of operations).
The order is fixed: center, then normalize, then pack and recurse. Undo steps are applied in reverse order. This is required because centering and row(A)-col(B) normalization do not commute in general.
Theorem 8.5 (Kernel correctness).
Assume KernelMM is implemented as in the Kernel pseudocode with:
  • Exact dyadic centering means at every node (Definition 3.4),
  • Dyadic $\ell_{2}$ normalization via $k(x)$ at every node (Definitions 6.1-6.2),
  • A sufficient base $\beta$ satisfying Theorem 7.1,
  • Sieve extraction $\mathcal{S}_{\beta,\sigma}$ utilizing the local scaling exponent $\sigma$.
  • Lattice alignment: The digit-band staging provides sufficient precision $b(n)$ such that the normalized coefficients are exactly representable in $\mathcal{L}_{\mathrm{coeff}}(\sigma)$ (i.e., no mantissa bits are dropped below the lattice resolution).
Therefore, for all $A,B\in\mathcal{L}_{D}^{n\times n}$ with $n$ a power of two, KernelMM returns $C=AB$ exactly in dyadic lattice semantics.
Proof.
We prove the claim by induction on $n$.
Base case ($n=1$).
The algorithm returns the scalar product $AB$, which is exact.
Inductive step.
Assume the theorem holds for size $n/2$. Consider inputs $A,B\in\mathcal{L}_{D}^{n\times n}$.
  1. 1.
    Centering and normalization: The algorithm forms centered matrices $A_{0},B_{0}$ using exact dyadic means, and then produces normalized matrices $A^{\prime},B^{\prime}$ by dyadic $\ell_{2}$ row/column scaling. By Proposition 6.3, these scalings are exact in $\mathcal{L}_{D}$. By Lemma 6.5, every entry of $A^{\prime}B^{\prime}$ is bounded in magnitude by $1$, so every dot product contributing to the band terms in Proposition 5.7 entrywise satisfies $\lvert T\rvert\leq 1$ and $\lvert L\rvert\leq 1$.
  2. 2.
    Packing identity and recursive products: Partition $A^{\prime},B^{\prime}$ into $2\times 2$ blocks and form packed blocks $(X_{ij},Y_{ij})$ by Definition 5.6. For each $(i,j)\in\{1,2\}^{2}$, Proposition 5.7 gives the three-band decomposition:
    $X_{ij}Y_{ij}~=~\beta G_{ij}+T_{ij}+\frac{1}{\beta}L_{ij},$ (62)
    where the middle band is the desired quadrant product $T_{ij}=(A^{\prime}B^{\prime})_{ij}$.
    The algorithm computes $Z_{ij}\leftarrow\textsc{KernelMM}(X_{ij},Y_{ij})$. Because $X_{ij},Y_{ij}\in\mathcal{L}_{D}^{(n/2)\times(n/2)}$, the induction hypothesis implies:
    $Z_{ij}~=~X_{ij}Y_{ij}$ (63)
    exactly.
  3. 3.
    Exact extraction: Because $\lvert T_{ij}\rvert\leq 1$ and $\lvert L_{ij}\rvert\leq 1$ entrywise, Theorem 7.1 ensures that the master gap condition (Lemma 5.3) holds entrywise for each packed product. Therefore, by the coefficient-lattice alignment hypothesis and Lemma 5.2, extraction applies: entrywise and gives:
    $\mathcal{S}_{\beta}(Z_{ij})~=~\mathcal{S}_{\beta}(X_{ij}Y_{ij})=T_{ij},$ (64)
    so the extracted blocks assembled by the algorithm equal the exact quadrants of $A^{\prime}B^{\prime}$, so the assembled matrix $T$ equals $A^{\prime}B^{\prime}$ exactly.
  4. 4.
    Undo normalization and un-centering: The algorithm then undoes the dyadic row/column scalings exactly (again by exponent-offset updates in $\mathcal{L}_{D}$), which recovers $A_{0}B_{0}$ exactly. It then recovers $AB$ from $A_{0}B_{0}$ using the bilinear centering identity (Lemma 8.1), which is an exact algebraic identity. Therefore the returned matrix exactly equals $AB$.
This completes the induction. ∎
9 Complexity and implementation model
9.1 Arithmetic recurrence
Lemma 9.1 (Per-node overhead is $O(m^{2})$ scalar primitives).
At a recursion node of dimension $m$, the non-recursive work of KernelMM consists of:
  1. 1.
    computing exact dyadic means (row/column sums and a division by $m=2^{t}$),
  2. 2.
    computing lifted squared norms $S(x)=\sum(Dx_{i})^{2}$ row-wise/column-wise and the corresponding exponents $k(x)$,
  3. 3.
    applying dyadic scaling by $2^{-k(x)}$ (exponent-offset updates),
  4. 4.
    packing/unpacking via $X_{ij}=A_{i1}+\beta^{-1}A_{i2}$ and $Y_{ij}=B_{1j}+\beta B_{2j}$,
  5. 5.
    applying the sieve entrywise to extract the middle bands,
  6. 6.
    and performing the rank-one centering correction.
Each of these steps takes $O(m^{2})$ scalar primitive operations (add/subtract/multiply by constants, and rounding calls).
Proof.
All steps are either (1) streaming passes over the $m^{2}$ entries (packing, scaling, sieve, correction) or (2) $m$ reductions each of length $m$ (means and squared norms), which also total $O(m^{2})$ scalar operations. The bitlength computation for each $k(x)$ is $O(1)$ per row/column once $S(x)$ is accumulated, so it contributes $O(m)$ additional work. ∎
Theorem 9.2 (Arithmetic complexity).
Let $T(n)$ denote the arithmetic time of KernelMM on $n\times n$ inputs, measured as the number of scalar ring operations (additions, multiplications by fixed constants such as powers of two, and nearest-integer rounding calls), assuming $n$ is a power of two, then:
$T(n)~=~4T(n/2)+O(n^{2}),$ (65)
and therefore:
$T(n)~=~\Theta(n^{2}\log n).$ (66)
Proof.
We now consider a single call of KernelMM on size $n$.
Recursive calls:
After centering/normalization and $2\times 2$ blocking, the algorithm forms four packed pairs $(X_{ij},Y_{ij})$ for $(i,j)\in\{1,2\}^{2}$ and makes exactly four recursive calls:
$Z_{ij}~\leftarrow~\textsc{KernelMM}(X_{ij},Y_{ij}),$ (67)
each on matrices of dimension $n/2$, which contributes $4T(n/2)$.
Nonrecursive work is $O(n^{2})$:
The nonrecursive work per node is $O(n^{2})$ by Lemma 9.1. The recursion makes $4$ subcalls on size $n/2$ blocks, so $T(n)=4T(n/2)+O(n^{2})$.
Solving the recurrence:
Here $a=4$, $b=2$, and the additive term is $f(n)=\Theta(n^{2})$. Because $n^{\log_{b}a}=n^{\log_{2}4}=n^{2}$, this is the Master Theorem critical case, which gives:
$T(n)~=~\Theta(n^{2}\log n).$ (68)
9.2 Bit/word complexity (fixed words, staged precision)
Theorem 9.3 (Bit/word complexity with digit-band staging).
Fix a word width $w$ for mantissa storage, and assume exponent offsets and loop indices fit in fixed-width machine integers. Let $b(n)$ denote the required effective mantissa precision (in bits) needed to represent packed temporaries and preserve the extractor slack $\delta$ at all decision sites for dimension $n$ under the chosen execution discipline.
If extended precision is disallowed, emulate $b(n)$-bit mantissas on $w$-bit words via digit-band (slice) staging with:
$K(n)~:=~\left\lceil\frac{b(n)}{w}\right\rceil$ (69)
words per scalar. Combining the arithmetic recurrence $T(n)=\Theta(n^{2}\log n)$ (Theorem 9.2) with slice staging, the total fixed-word cost satisfies:
$B(n)~=~\Theta\!\left(n^{2}\log n\cdot K(n)\right),$ (70)
i.e., the kernel is soft-quadratic whenever $b(n)=O(\log^{c}n)$ for some constant $c$. In particular, if $b(n)=\Theta(\log n)$ and $w$ is fixed, then $B(n)=\Theta(n^{2}\log^{2}n)=\tilde{O}(n^{2})$.
Remark 9.4 (How this resolves “catastrophic truncation”).
In a fixed-format implementation (e.g., IEEE-754 binary64), choosing a large packing base $\beta=2^{b}$ consumes about $b$ bits of significand headroom: lower-band content and/or low bits of the middle band may be truncated before extraction. Slice staging restores that headroom by representing each scalar as $K(n)$ fixed-width “digits” and executing the same GPR kernel on these digits. In float mode, any remaining staged/truncated effects are accounted for explicitly as an auxiliary drift term $\varepsilon_{\mathrm{aux}}$ in the drift ledger (Definition 12.9).
Remark 9.5 (Exponent offsets are not a hidden unbounded-precision model).
Exponent offsets are ordinary machine integers in this model. They are updated by adding node-dependent shifts. If an exponent offset overflows the chosen integer type, the instance leaves the declared execution regime, exactly as in classical implementations that rely on fixed-width exponents.
Proof.
Theorem 9.2 gives $\Theta(n^{2}\log n)$ scalar primitives (add/subtract, shifts by powers of two, packing updates, and extractor bookkeeping) across the recursion. Under digit-band staging, each such scalar primitive is implemented by $O(K(n))$ fixed-width word operations on the slice vectors. Therefore the total word cost is multiplied by $K(n)$, giving $B(n)=\Theta(n^{2}\log n\cdot K(n))$. ∎
Remark 9.6 (Staging constant factors (no “cheating”)).
In Theorem 9.3, the factor $K(n)=\lceil b(n)/w\rceil$ accounts for representing a $b(n)$-bit mantissa using $w$-bit words. Each staged primitive therefore costs $O(K(n))$ word operations. Concretely, multiword addition, subtraction, and rounding require linear passes over the $K(n)$ digits; multiplication by fixed powers of two is an exponent update; and carry propagation across digit boundaries is bounded by the fixed staging width. Thus the hidden constant in $B(n)=\Theta(n^{2}\log n\cdot K(n))$ depends on the efficiency of the digit-band implementation (e.g., how slice boundaries are handled) but remains independent of $n$.
9.3 Example: integer polynomial convolution as a GPR-admissible packing
Here we show that the three-band packing plus sieve extraction mechanism is not specific to matrix multiplication.
Problem.
Let $f,g\in\mathbb{Z}^{n}$ and define their length-(2n1)(2n-1) convolution $(f\ast g)\in\mathbb{Z}^{2n-1}$ as:
$(f\ast g)[t]~:=~\sum_{i+j=t}f[i]g[j]~~(t=0,\dots,2n-2).$ (71)
Assume $n$ is a power of two (pad with zeros otherwise). We view $\mathbb{Z}\subset\mathcal{L}_{D}$ for $D=1$ (or more generally embed entrywise into $\mathcal{L}_{D}$).
Dyadic normalization for vectors (same invariant as Part I.)
For a vector $x\in\mathcal{L}_{D}^{m}$, define $S(x)$ and $k(x)$ as in Definitions 6.1 and 6.2, and set $x^{\prime}:=2^{-k(x)}x$. Proposition 6.3 gives $\|x^{\prime}\|_{2}\leq 1$.
Lemma 9.7 (Dimension-free bound for convolution coefficients under dyadic normalization).
If $a,b\in\mathcal{L}_{D}^{m}$ satisfy $\|a\|_{2}\leq 1$ and $\|b\|_{2}\leq 1$, then every coefficient of $a\ast b$ satisfies:
$|(a\ast b)[t]|~\leq~1~\text{for all}~t.$ (72)
Proof.
Fix $t$. The coefficient $(a\ast b)[t]=\sum_{i}a[i]b[t-i]$ is a dot product of $a$ with a shifted/truncated version of $b$. Truncation cannot increase $\ell_{2}$ norm, so the relevant vector $b^{(t)}$ satisfies $\|b^{(t)}\|_{2}\leq\|b\|_{2}\leq 1$. By Cauchy–Schwarz, $|(a\ast b)[t]|\leq\|a\|_{2}\|b^{(t)}\|_{2}\leq 1$. ∎
Three-band packing.
Partition $f=(f_{0},f_{1})$ and $g=(g_{0},g_{1})$ into halves of length $n/2$, and then apply dyadic normalization to each half to obtain $f_{0}^{\prime},f_{1}^{\prime},g_{0}^{\prime},g_{1}^{\prime}$ with $\|f_{i}^{\prime}\|_{2}\leq 1$ and $\|g_{j}^{\prime}\|_{2}\leq 1$.
Fix an integer base $\beta>1$ and form packed vectors:
$F$ $:=$ $f_{0}^{\prime}+\beta^{-1}f_{1}^{\prime},$ (73)
$G$ $:=$ $g_{0}^{\prime}+\beta g_{1}^{\prime}.$ (74)
Proposition 9.8 (Three-band identity for convolution).
The packed convolution admits the three-band form:
$F\ast G~=~\beta U+T+\beta^{-1}L,$ (75)
where:
$T$ $:=$ $f_{0}^{\prime}\ast g_{0}^{\prime}+f_{1}^{\prime}\ast g_{1}^{\prime},$ (76)
$U$ $:=$ $f_{0}^{\prime}\ast g_{1}^{\prime},$ (77)
$L$ $:=$ $f_{1}^{\prime}\ast g_{0}^{\prime}.$ (78)
Proof.
Expanding distributively gives:
$(f_{0}^{\prime}+\beta^{-1}f_{1}^{\prime})\ast(g_{0}^{\prime}+\beta g_{1}^{\prime})~=~f_{0}^{\prime}\ast g_{0}^{\prime}+f_{1}^{\prime}\ast g_{1}^{\prime}+\beta(f_{0}^{\prime}\ast g_{1}^{\prime})+\beta^{-1}(f_{1}^{\prime}\ast g_{0}^{\prime}),$ (79)
and group by powers of $\beta$. ∎
Uniform gap and exact extraction.
By Lemma 9.7, coefficients are bounded. To maintain exactness on the refined lattice produced by normalization, we choose $\beta$ to satisfy the sufficient-base condition (Theorem 7.1) given the recursion depth.
Remark 9.9 (Connection to Part IV).
This is an instance of the master condition (Definition 19.1) with a constant-width exponent window and bounded mixing. The depth-uniform bound is supplied here by dyadic $\ell_{2}$ normalization, exactly as in the MM kernel.
Sufficient base and exact extraction
By Lemma 9.7, every coefficient of each convolution term in Proposition 9.8 has magnitude $\leq 1$. Thus the middle band $T:=f_{0}^{\prime}\ast g_{0}^{\prime}+f_{1}^{\prime}\ast g_{1}^{\prime}$ and the tail band $L:=f_{1}^{\prime}\ast g_{0}^{\prime}$ satisfy $|T|\leq 2$ and $|L|\leq 1$ coefficientwise.
Therefore, choosing $\beta$ so that the master gap inequality holds coefficientwise (Lemma 5.3) the master gap inequality holds coefficientwise (Lemma 5.3) gives exact middle-band recovery coefficientwise by Lemma 5.2. For example, any $\beta$ satisfying:
$D^{2}\left(\frac{|L|}{\beta}+\frac{|T|}{\beta}+\frac{|L|}{\beta^{2}}\right)\leq\frac{1}{2}-\delta$ (80)
with $|T|\leq 2$ and $|L|\leq 1$ suffices.
Theorem 9.10 (Correctness and complexity of GPR convolution).
There exists a sufficient base sequence $\beta(\sigma)$ (growing with recursion depth $\sigma$) such that the recursive packed convolution algorithm recovers $f\ast g$ exactly in dyadic lattice semantics. The arithmetic recurrence is:
T(n)=4T(n/2)+O(n),T(n)~=~4T(n/2)+O(n), (81)
giving an arithmetic cost of $\Theta(n^{2})$. Under the fixed-word model with digit-band staging, the bit/word cost is:
$B(n)~=~\Theta\!\left(n^{2}\cdot K(n)\right)~=~\tilde{O}(n^{2}),$ (82)
where $K(n)=\lceil b(n)/w\rceil=\Theta(\log n)$ accounts for the logarithmic growth of the required precision (base $\beta$) with depth.
Proof.
We prove correctness and complexity by induction:
Correctness.
At each node, the packed product $F\ast G$ admits the three-band form (Proposition 9.8). Dyadic normalization (Lemma 9.7) ensures the band coefficients are bounded in the normalized lattice $\mathcal{L}_{\mathrm{coeff}}(\sigma)$. By choosing $\beta$ to satisfy the sufficient-base condition (Theorem 7.1) for the current depth $\sigma$, the master gap inequality holds, and the scaled sieve $\mathcal{S}_{\beta,\sigma}$ recovers the middle band $T$ exactly. The algorithm performs $4$ subcalls (implicitly or explicitly to recover all components of the convolution from the packed forms), maintaining exactness throughout via the staged precision model.
Complexity.
The algorithm performs $4$ recursive calls on size $n/2$ (one for each quadrant product in the convolution algebraic expansion), plus linear time $O(n)$ for packing, normalization, and extraction. The arithmetic recurrence is $T(n)=4T(n/2)+O(n)$, which solves to $T(n)=\Theta(n^{2})$ by the Master Theorem. In the fixed-word model, maintaining exactness requires a mantissa width $b(n)$ that scales with the recursion depth $\sigma\approx\log n$. Using digit-band staging, each scalar operation costs $O(K(n))$ word operations, so the total bit/word complexity is $B(n)=\Theta(n^{2}\cdot K(n))=\Theta(n^{2}\log n)$. ∎
10 Prior related work and what is new here
The core algebraic trick behind our quadrant packing is not new: embed multiple quantities into a single multiplication by introducing a large parameter (a base $\beta$, or a formal variable $\lambda$), and then decodes a desired coefficient/band by an extractor (rounding, digit extraction, interpolation, CRT, etc.). Variants of this idea appear across several communities.
What is new in this note is not the existence of coefficient/band extraction, but the way recursive invariant amplification (centering plus dyadic $\ell_{2}$ normalization at every recursion node) makes coefficient extraction compatible with a model-honest intermediate-budget analysis: it gives a constant base $\beta_{\star}$, eliminates bit-growth of mantissas, and improves feasibility/robustness by bounding intermediates at every depth rather than only at the root.
10.1 APA and border-rank methods in fast matrix multiplication
In fast matrix multiplication theory, approximate (APA) bilinear algorithms introduce a formal parameter (often written $\lambda$ or $\varepsilon$) to separate contributions across different powers of that parameter and then recover a desired coefficient via an exactification step (interpolation/limit argument). This is similar to the work of Bini et al. on approximate MM [18] and in Bini’s later discussion relating exact and approximate bilinear algorithms [19]. A closely related tensor-language view is that such constructions upper-bound the border rank of the MM tensor, which is sufficient to derive asymptotic exponents (see [20] and Pan’s trilinear/field-extension methods [21]).
Our three-band identity:
$z~=~\beta U+T+L/\beta$ (83)
uses the same coefficient-separation idea (a parameter separates bands), but the aim is different: we use it for model-honest exact (or commensurate-accuracy) computation under bounded intermediates, rather than to prove a smaller asymptotic exponent in an unbounded-precision algebraic model. Paszynski [17] provides a clear description of how simple uncontrolled coefficient-separation leads to either exponential growth of intermediate values or incorrect results.
10.2 Kronecker substitution and digit packing in exact arithmetic
In exact arithmetic, Kronecker substitution evaluates a polynomial at a sufficiently large base (often a power of two) so that coefficients of the product can be recovered by digit extraction from the resulting integer. This is the canonical “choose $\beta$ large enough so bands do not overlap” principle, used to reduce polynomial multiplication to integer multiplication (see, e.g., Harvey’s treatment and refinements [22]). Our base $\beta$ plays an analogous role, but for block products (quadrants) rather than polynomial coefficients.
10.3 Mixed-domain packing (one multiply does multiple jobs)
A conceptually adjacent line of work packs multiple computations into a single multiplication in a larger algebraic domain (e.g., complex arithmetic so one complex GEMM implements multiple real GEMMs). High-performance “1M” methods for complex matrix multiplication are a concrete example [23]. While this does not use $\beta$ and $\beta^{-1}$ explicitly, it shares the same meta-idea: embed multiple targets into one multiply, then decode by structured linear operations.
10.4 What changes the game: recursive invariant amplification
The difficulty in turning coefficient-separation tricks into a practical, format-robust algorithm is that intermediate magnitudes can grow with dimension and depth, forcing depth-dependent bases $\beta(m)$ and thus growing word sizes. The key design move here is to enforce, at every recursion node, a local invariant (centering plus dyadic $\ell_{2}$ normalization) that gives dimension-free dot-product bounds. This makes the master gap condition hold with a sufficient base $\beta(\sigma)$ (Theorem 7.1 in Part I), bounding intermediate magnitudes and making precision costs structure-dependent; when exactness beyond a fixed format is required, digit-band staging gives a soft-quadratic (polylog overhead) bit-complexity bound.
From a robustness viewpoint, the same recursive normalization gives a strictly larger feasible region under a fixed intermediate budget: global (one-shot) preprocessing does not control the norms of the packed linear combinations that arise deeper in the recursion, whereas the node-local discipline does. In Part IV we abstract this as a general admissibility condition (master condition/BWBM) coupled to an extractor interface.
Part II Float-GPR: baseline comparators and commensurate accuracy (Part II)
11 Scope and goals
What changes from Part I
  • We no longer claim exactness of every intermediate operation (only commensurate accuracy w.r.t. classical)
  • We compare against a baseline comparator executed in a fixed format $\mathsf{F}$ (e.g., IEEE-754).
  • Correctness statements become commensurate accuracy and decision invariance under drift budgets.
What stays the same
  • Recursive blocked structure (schoolbook $2\times 2$ recursion).
  • Graded projection at every node to bound intermediates.
  • A gap/slack parameter $\delta$ controlling discrete extraction/decision stability.
11.1 Proof dependency map (Part II)
The proof spine for float-GPR (commensurate-accuracy) has the following dependencies:
  1. 1.
    Semantics:
    • Definition 12.1 (execution format),
    • Definition 12.4 (baseline comparator),
    • Definition 12.7 (commensurate accuracy).
  2. 2.
    Decision control:
    • Definition 12.8 (decision sites),
    • Definition 12.9 (ledger),
    • Lemma 12.10 (decision invariance),
    • Lemma 13.9 (sieve stability under drift).
  3. 3.
    Guarded normalization:
  4. 4.
    Composition:
    • Theorem 13.10 (float-GPR commensurate theorem),
    • Theorem 15.3 (task-level float reduction template).
  5. 5.
    IEEE instantiation: Lemma 12.3 and its corollaries for $\varepsilon_{\mathrm{round}}$.
12 Execution format model
Definition 12.1 (Execution format $\mathsf{F}$).
An execution format $\mathsf{F}$ is a fully specified model of scalar arithmetic consisting of:
  1. 1.
    Radix and representable set: A radix $b\geq 2$ (here $b=2$) and a set $\mathrm{Val}(\mathsf{F})\subset\mathbb{R}\cup\{\pm\infty,\mathrm{NaN}\}$ of representable values.
  2. 2.
    Precision: A precision parameter $p\in\mathbb{N}$ specifying the number of significand bits (including any hidden leading bit convention). Equivalently, in normal form each nonzero finite representable value has the shape:
    $x~=~\pm mb^{e},$ (84)
    where $m$ is a radix-$b$ significand with $p$ digits/bits.
  3. 3.
    Exponent range and subnormals: Integers $e_{\min}\leq e\leq e_{\max}$ determining the normal exponent range. The format also specifies whether and how subnormals are represented and rounded (including flush-to-zero vs gradual underflow) and what exponent is used for subnormal values.
  4. 4.
    Rounding rule: A deterministic rounding map to the nearest representable value under a specified rule (e.g., round-to-nearest; ties-to-even; or directed roundings). This includes a tie-breaking convention whenever the nearest representable values are not unique.
  5. 5.
    Exception semantics: A specification of how exceptional results are produced and propagated, including: overflow (e.g., to $\pm\infty$ or max finite), underflow/subnormal handling, division by zero, invalid operations, and NaN propagation.
Rounded execution map.
We write $\mathrm{fl}_{\mathsf{F}}(\cdot)$ for the result of executing a scalar real operation in format $\mathsf{F}$, including rounding and exception semantics. Concretely, for a real value $x\in\mathbb{R}$, $\mathrm{fl}_{\mathsf{F}}(x)$ denotes the result of rounding $x$ to $\mathrm{Val}(\mathsf{F})$ according to $\mathsf{F}$’s rounding rule (possibly resulting in $\pm\infty$ or $\mathrm{NaN}$ per the exception rules).
Lift to algorithms.
For an algorithm specified as a sequence of scalar operations on real inputs, its execution in format $\mathsf{F}$ means that every scalar operation is replaced by the corresponding $\mathrm{fl}_{\mathsf{F}}(\cdot)$-wrapped operation, with the same evaluation order.
Definition 12.2 (Baseline-safe (normal-range) regime).
For a given algorithm $\mathcal{A}$ executed in $\mathsf{F}$, we say an input instance lies in the normal-range regime if all primitive operations performed by $\mathcal{A}$ return finite normal numbers (no overflow, no NaN, no subnormals).
Lemma 12.3 (IEEE rounding as additive drift (normal-range)).
Assume $\mathsf{F}$ is an IEEE-754 binary format with precision $p$ and unit roundoff $u=2^{-p}$, and assume the normal-range regime (Definition 12.2) holds, then for any real $x$ with finite normal rounded value:
$\lvert\mathrm{fl}_{\mathsf{F}}(x)-x\rvert~\leq~u\lvert x\rvert$. (85)
Proof.
This is the standard IEEE floating-point model in the normal range under round-to-nearest: the rounded value differs from $x$ by at most half an ulp, which is bounded by $u|x|$ for normal numbers. ∎
12.1 Baseline comparators
Definition 12.4 (Baseline comparator $\mathrm{BaseMM}_{\mathsf{F}}$).
Fix a specific classical matrix multiplication procedure (evaluation order and blocking strategy). Define $\mathrm{BaseMM}_{\mathsf{F}}(A,B)$ to be the result of executing that procedure in format $\mathsf{F}$, i.e., every scalar operation is wrapped by $\mathrm{fl}_{\mathsf{F}}(\cdot)$ with $\mathsf{F}$ exception semantics.
Remark 12.5.
Baseline comparators can be defined similarly for factorization and orthogonalization kernels (LDL, QR, etc.) and will be used in later float-GPR case studies.
12.2 Accuracy notions
Definition 12.6 (Baseline-exactness).
An algorithm $\mathcal{A}$ is baseline-exact w.r.t. $\mathrm{BaseMM}_{\mathsf{F}}$ if for all representable inputs it produces exactly the same output as $\mathrm{BaseMM}_{\mathsf{F}}$ (bitwise in $\mathsf{F}$, including exception behavior).
Definition 12.7 (Commensurate accuracy (feasibility + bounded drift)).
An algorithm $\mathcal{A}$ has commensurate accuracy w.r.t. $\mathrm{BaseMM}_{\mathsf{F}}$ if:
  • (Feasibility) whenever $\mathrm{BaseMM}_{\mathsf{F}}(A,B)$ is finite (no NaNs/Infs), and $\mathcal{A}(A,B)$ is also finite.
  • (Drift bound) the output perturbation of $\mathcal{A}$ relative to exact $AB$ is bounded by a stated function of the baseline perturbation or by an explicit ledger bound that is no worse than baseline under the same format regime.
(This may be instantiated using a norm bound, componentwise bound, or ulp-based bound.)
12.3 Drift ledger and decision sites
Definition 12.8 (Decision sites).
A decision site is any point where a discrete choice depends on rounded values (e.g., sieve extraction boundaries, pivot selection, sign tests, or thresholding). Correctness is reduced to showing decisions match an ideal reference under bounded drift.
Definition 12.9 (Drift budget ledger).
Fix slack $\delta\in(0,1/2]$. Define drift components (all in the units of the rounding argument at the decision site):
  • $\varepsilon_{\mathrm{mean}}$,
  • $\varepsilon_{\mathrm{round}}$,
  • $\varepsilon_{\mathrm{aux}}$.
Here $\varepsilon_{\mathrm{aux}}$ includes auxiliary effects such as guarded $\ell_{2}$ normalization overshoot (Corollary 13.5) and any other bounded truncation/staging effects.
Let:
$\varepsilon_{\mathrm{tot}}~:=~\varepsilon_{\mathrm{mean}}+\varepsilon_{\mathrm{round}}+\varepsilon_{\mathrm{aux}},$ (86)
with required ledger inequality:
$\varepsilon_{\mathrm{tot}}~\leq~\delta$ (87)
to guarantee decision invariance.
12.4 Decision invariance
Lemma 12.10 (Decision invariance from slack).
Let $f:\mathbb{R}\to\mathbb{R}$ be a decision statistic and consider a threshold decision of the form:
$\mathrm{Dec}(x)~=~\mathbf{1}\{f(x)\geq 0\}.$ (88)
Assume an ideal value $s^{\star}=f(x)$ and an implemented value $s=f(x)+\eta=s^{\star}+\eta$ with additive drift $\eta$. If there exists $\delta>0$ such that:
$|s^{\star}|~\geq~\delta~~\text{and}~~|\eta|~\leq~\delta,$ (89)
then $\mathrm{Dec}$ is invariant under the drift: $\mathbf{1}\{s\geq 0\}=\mathbf{1}\{s^{\star}\geq 0\}$.
More generally, if a discrete decision is defined by rounding a real statistic to an integer,
$\mathrm{Dec}(x)~=~g\left(\left\lfloor f(x)\right\rceil\right)$ (90)
for some map $g:\mathbb{Z}\to\mathcal{Y}$, and there exist $m\in\mathbb{Z}$ and $\delta\in(0,1/2]$ such that:
$\lvert f(x)-m\rvert~\leq~\frac{1}{2}-\delta~~\text{and}~~\lvert\eta\rvert~<~\delta,$ (91)
then $\left\lfloor f(x)+\eta\right\rceil=m=\left\lfloor f(x)\right\rceil$, and so the decision is invariant, i.e.:
$g(\left\lfloor f(x)+\eta\right\rceil)=g(\left\lfloor f(x)\right\rceil).$ (92)
Proof.
For the sign-threshold form, if $|s^{\star}|\geq\delta$ and $|\eta|\leq\delta$, then $s=s^{\star}+\eta$ cannot cross zero:
$s^{\star}$ $\geq$ $\delta\Rightarrow s~\geq~\delta-\delta\geq 0,$ (93)
$s^{\star}$ $\leq$ $-\delta\Rightarrow s~\leq~-\delta+\delta\leq 0.$ (94)
Thus $\mathbf{1}\{s\geq 0\}=\mathbf{1}\{s^{\star}\geq 0\}$.
For the rounding-based form, apply Lemma 4.1 (margin-based rounding stability) with $r=f(x)$ and $e=\eta$. The hypotheses imply $\left\lfloor f(x)+\eta\right\rceil=\left\lfloor f(x)\right\rceil=m$, and applying $g$ preserves equality. ∎
13 Float-GPR MM: commensurate accuracy theorem
Definition 13.1 (Compensated guarded squared norm (Kahan)).
Fix an execution format $\mathsf{F}$ with unit roundoff $u$ and assume the normal-range regime holds (Definition 12.2). For $x\in\mathrm{Val}(\mathsf{F})^{m}$, let:
$S~:=~\|x\|_{2}^{2}~=~\sum_{i=1}^{m}x_{i}^{2},$ (95)
and let $S_{\mathrm{Kahan}}$ be the result of Kahan-compensated summation applied to the nonnegative summands $\{x_{i}^{2}\}$. Define the guarded estimator:
$\widehat{S}(x)~:=~(1+\eta_{m})S_{\mathrm{Kahan}},$ (96)
where $\eta_{m}$ is chosen so that $\widehat{S}(x)\geq S$ (one-sided safety) on the baseline-feasible regime. A concrete sufficient choice is given by the bound in Lemma 13.4 (e.g., using a standard Kahan relative-error bound).
Remark 13.2 (Normal-range feasibility of dyadic scaling).
All claims in this subsection are made on the baseline-safe normal-range regime (Definition 12.2). In particular, the exponent shift:
$x~\mapsto~\mathrm{fl}_{\mathsf{F}}(2^{-\widehat{k}(x)}x)$ (97)
is assumed to remain in the normal range (i.e., no subnormals/underflow). Instances for which guarded scaling would produce subnormals are treated as leaving the success set (unless subnormal handling is analyzed separately).
13.1 Guarded dyadic $\ell_{2}$ normalization (float-GPR)
Definition 13.3 (Guarded squared norm and dyadic exponent (float mode)).
Fix an execution format $\mathsf{F}$ with unit roundoff $u$ and assume the normal-range regime holds (Definition 12.2). For a vector $x\in\mathrm{Val}(\mathsf{F})^{m}$, let:
$S(x)~:=~\|x\|_{2}^{2}=\sum_{i=1}^{m}x_{i}^{2}.$ (98)
Let $S_{\mathrm{Kahan}}(x)$ be the Kahan-compensated sum of the nonnegative summands $\{x_{i}^{2}\}_{i=1}^{m}$ computed in format $\mathsf{F}$. Define the guarded estimator:
$\widehat{S}(x)~:=~(1+\eta_{m})S_{\mathrm{Kahan}}(x),$ (99)
where $\eta_{m}\geq 0$ is chosen so that $\widehat{S}(x)\geq S(x)$ on the baseline-feasible regime. Define the guarded dyadic exponent:
$\widehat{k}(x)~:=~\min\{k\in\mathbb{Z}_{\geq 0}:\widehat{S}(x)\leq 2^{2k}\},$ (100)
and the normalized vector:
$\widehat{x}~:=~\mathrm{fl}_{\mathsf{F}}(2^{-\widehat{k}(x)}x).$ (101)
Lemma 13.4 (Kahan guard: safety and bounded overshoot).
Assume Definition 13.3, then on the baseline-feasible regime:
  1. 1.
    (One-sided safety) $\widehat{S}(x)\geq S(x)$, which implies $\|2^{-\widehat{k}(x)}x\|_{2}\leq 1$ in exact real arithmetic.
  2. 2.
    (Bounded overshoot) If Kahan summation satisfies a relative-error bound of the form:
    $|S_{\mathrm{Kahan}}(x)-S(x)|~\leq~\rho_{m}S(x)~\text{with}~\rho_{m}=2u+O(mu^{2}),$ (102)
    then choosing $\eta_{m}\geq\frac{\rho_{m}}{1-\rho_{m}}$ implies:
    $S(x)$ $\leq$ $\widehat{S}(x)~\leq~\Lambda_{m}S(x),$ (103)
    $\Lambda_{m}$ $:=$ $(1+\eta_{m})(1+\rho_{m})~=~1+O(u)+O(mu^{2}).$ (104)
Proof.
  1. 1.
    By construction of $\eta_{m}$ we have $\widehat{S}(x)\geq S(x)$, $\widehat{S}(x)\leq 2^{2\widehat{k}(x)}$ by definition of $\widehat{k}(x)$, so $S(x)\leq 2^{2\widehat{k}(x)}$, and therefore:
    $\|2^{-\widehat{k}(x)}x\|_{2}^{2}~=~2^{-2\widehat{k}(x)}S(x)~\leq~1.$ (105)
  2. 2.
    If $|S_{\mathrm{Kahan}}-S|\leq\rho_{m}S$, then $S_{\mathrm{Kahan}}\geq(1-\rho_{m})S$., so choosing $\eta_{m}\geq\rho_{m}/(1-\rho_{m})$ gives:
    $(1+\eta_{m})S_{\mathrm{Kahan}}~\geq~(1+\eta_{m})(1-\rho_{m})S~\geq~S,$ (106)
    which proves safety. Also, $S_{\mathrm{Kahan}}\leq(1+\rho_{m})S$, so:
    $\widehat{S}$ $=$ $(1+\eta_{m})S_{\mathrm{Kahan}}$ (107)
    $\leq$ $(1+\eta_{m})(1+\rho_{m})S~=~\Lambda_{m}S.$ (108)
Corollary 13.5 (Ledger impact of guarded normalization).
Under the bounded-overshoot conclusion of Lemma 13.4, the guarded exponent satisfies:
$0~\leq~\widehat{k}(x)-k_{\mathrm{ideal}}(x)~\leq~\left\lceil\frac{1}{2}\log_{2}\Lambda_{m}\right\rceil,$ (109)
where $k_{\mathrm{ideal}}(x):=\min\{k:S(x)\leq 2^{2k}\}$. Consequently, the over-normalization contributes a depth-independent auxiliary drift term:
$\varepsilon_{\mathrm{aux}}~=~O(\log_{2}\Lambda_{m})~=~O(u)+O(mu^{2}),$ (110)
and it can only improve the master gap inequality (it shrinks the magnitudes of $T$ and $L$ bands).
Proof.
The bound on $\widehat{k}-k_{\mathrm{ideal}}$ is immediate from $\widehat{S}\in[S,\Lambda_{m}S]$ by taking $\log_{2}$ and dividing by $2$. The ledger statement follows because any effect of over-normalization on decision statistics is bounded by the same constant-factor window. Finally, shrinking operands cannot increase band magnitudes in the three-band identity, so gaps cannot worsen. ∎
Remark 13.6 (Concrete choice of $\eta_{m}$ from a Kahan bound).
A standard bound for Kahan summation of nonnegative terms is $\rho_{m}\leq 2u+cmu^{2}$ for a small constant $c$ in the normal-range regime. A conservative explicit choice is therefore:
$\eta_{m}$ $:=$ $\frac{\rho_{m}}{1-\rho_{m}}$ (111)
$\rho_{m}$ $:=$ $2u+cmu^{2}.$ (112)
For IEEE binary64, $u=2^{-53}$ and $u^{2}=2^{-106}$, so even for very large $m$ the term $mu^{2}$ is negligible in practice.
Remark 13.7 (Scope of the Kahan guard assumption).
We only require one-sided safety $\widehat{S}(x)\geq\|x\|_{2}^{2}$ and bounded overshoot on the baseline-feasible normal-range regime. Any guarded estimator meeting Definition 13.3 is sufficient. Kahan summation is a convenient instantiation because it provides an explicit bound with small constants in standard references (e.g., Higham).
Remark 13.8 (Tightening via error-free transforms).
Instead of plain squaring and Kahan summation, we may compute each square using an error-free transform (e.g., TwoProd) and accumulate using TwoSum-style compensated summation. This reduces the effective $\rho_{m}$ constant in practice and can give bounds of the form $\rho_{m}=O(u)+O(mu^{2})$ with smaller leading constants. The GPR analysis is unchanged, but only $\Lambda_{m}$ improves.
Lemma 13.9 (Sieve stability under additive drift).
Fix $\beta\geq 2$ and consider a packed scalar:
$z^{\star}~=~\beta U+T+L/\beta$ (113)
with $U,T\in D^{-2}\mathbb{Z}$. Let $\delta\in(0,1/2]$, and suppose the ideal master gap holds with slack $\delta$, i.e.:
$\lvert D^{2}L/\beta\rvert+\lvert D^{2}T/\beta\rvert+\lvert D^{2}L/\beta^{2}\rvert~\leq~\frac{1}{2}-\delta.$ (114)
Let the implemented scalar be $z=z^{\star}+\eta$ with additive drift $\eta\in\mathbb{R}$, then if:
$\lvert D^{2}\eta\rvert~\leq~\delta,$ (115)
the sieve extraction is invariant, i.e., $\mathcal{S}_{\beta}(z)=\mathcal{S}_{\beta}(z^{\star})=T$.
Proof.
The sieve uses only the two rounded quantities $\left\lfloor D^{2}z\right\rceil$ and $\left\lfloor D^{2}z/\beta\right\rceil$. Under the ideal slack $\delta$, both rounding arguments are at distance at least $\delta$ from their respective half-integer boundaries. The additive perturbation $\eta$ shifts each rounded argument by at most $\delta$ in the lifted units, so Lemma 12.10 (or Lemma 4.1 applied to the lifted arguments) implies the rounding outcomes do not change. Therefore the sieve output does not change. ∎
Theorem 13.10 (Float-GPR extraction invariance implies commensurate accuracy).
Fix an execution format $\mathsf{F}$ and a baseline comparator $\mathrm{BaseMM}_{\mathsf{F}}$ (Definition 12.4), and fix a slack parameter $\delta\in(0,1/2]$. Consider a float-GPR matrix-multiplication kernel $\mathcal{A}_{\mathrm{GPR}}$ that operates on input $(A,B)$ by:
  1. 1.
    Applying graded projection steps (centering/normalization) to produce node-local operands;
  2. 2.
    Performing packed products; and
  3. 3.
    Recovering middle-band quantities via sieve-style extraction decisions (the same discrete decisions as in Part I), but with all primitive operations executed in format $\mathsf{F}$.
Assume the following hypotheses hold at every recursion node and for every scalar extraction site:
  1. 1.
    Ideal decomposition: In exact real arithmetic, each packed scalar to be decoded admits a three-band form $z^{\star}=\beta U^{\star}+T^{\star}+L^{\star}/\beta$ (as in Part I), and the algorithm’s desired output at that site is a fixed function of the middle band $T^{\star}$.
  2. 2.
    Slack at the ideal decision boundary: The ideal statistics defining the extraction decision have margin at least $\delta$ from their decision boundaries (e.g., the master gap inequality holds with residual slack $\delta$).
  3. 3.
    Ledger-bounded drift: The total drift between the implemented and ideal decision statistics is bounded by $\varepsilon_{\mathrm{tot}}\leq\delta$ (Definition 12.9).
  4. 4.
    Feasible baseline regime. The baseline comparator produces a finite result on $(A,B)$ (no NaNs/Infs), i.e., $(A,B)$ lies in the baseline-feasible regime.
Thus:
  1. 1.
    (Extraction decision invariance) All discrete extraction decisions made by $\mathcal{A}_{\mathrm{GPR}}$ coincide with the ideal real-arithmetic decisions, so the extracted middle-band values equal the ideal middle-band values at every site.
  2. 2.
    (Commensurate feasibility) On the baseline-feasible regime, $\mathcal{A}_{\mathrm{GPR}}(A,B)$ is also finite and defined. More generally, if graded projection keeps intermediate magnitudes bounded more tightly than the baseline, then $\mathcal{A}_{\mathrm{GPR}}$ can be defined on a strictly larger feasible regime (it may avoid spurious overflow/underflow where the baseline fails).
  3. 3.
    (Commensurate accuracy (conservative form)) The output perturbation of $\mathcal{A}_{\mathrm{GPR}}$ is governed solely by the same drift terms budgeted in the ledger. In particular, there is no additional accuracy loss from wrong extraction decisions.
Proof.
We now prove each item:
Extraction decision invariance
Fix any scalar extraction site. By hypothesis (2), the ideal decision statistic has margin at least $\delta$ from its boundary. By hypothesis (3), the implemented statistic differs from the ideal by drift at most $\varepsilon_{\mathrm{tot}}\leq\delta$. Therefore Lemma 12.10 applies and the implemented decision equals the ideal decision. In particular, for sieve-style extraction this means the same integer rounding outcomes (or sign/threshold outcomes) occur as in the ideal analysis, so the extracted value equals the intended middle-band quantity $T^{\star}$ at that site. Because this holds for every extraction site at every recursion node, all extracted middle-band values in the computation coincide with their ideal counterparts. For sieve-style extraction sites, this is exactly Lemma 13.9, i.e., slack $\delta$ plus drift $\varepsilon_{\mathrm{tot}}\leq\delta$ implies the extracted band is unchanged. If graded projection includes guarded dyadic $\ell_{2}$ normalization, then Corollary 13.5 quantifies its contribution to $\varepsilon_{\mathrm{aux}}$ (depth-independent), and Lemma 13.14 guarantees the master gap inequality is preserved (or improved) by over-normalization whenever execution remains in-range.
Commensurate feasibility
On the baseline-feasible regime, all baseline operations remain finite in $\mathsf{F}$. The float-GPR kernel differs from the baseline only by inserting graded projection steps and by changing the evaluation order inside blocked products. It does not introduce new operations with intrinsically larger magnitude than the intermediates it replaces. More specifically, graded projection is designed specifically to bound intermediate magnitudes before multiplication/accumulation, so whenever the baseline remains in range, the float-GPR kernel can be arranged to remain in range as well, giving a finite output. If graded projection bounds intermediates strictly more tightly than the baseline’s pipeline, then there exist instances where the baseline overflows/underflows due to intermediate growth while float-GPR remains in range, giving strict enlargement of the feasible regime.
Commensurate accuracy (conservative form)
By (1), all discrete extraction decisions are identical to the ideal decisions, so the algorithm never incurs a catastrophic error due to decoding the wrong band. Therefore the only sources of numerical error relative to the ideal real-arithmetic computation are those explicitly counted in the ledger (rounding of projection steps, rounding inside the implemented arithmetic, and any auxiliary truncations, if present). This gives the stated conservative commensurate-accuracy guarantee: correctness of the discrete structure plus a bounded continuous drift term. ∎
Remark 13.11.
This lemma converts known magnitude envelopes (from graded projection) into explicit $\varepsilon_{\mathrm{round}}$ bounds for the drift ledger.
Remark 13.12 (What is and is not claimed).
This theorem guarantees feasibility preservation on the baseline-feasible regime and invariance of discrete decisions under a drift ledger. It does not claim a global forward-error inequality unless stated separately.
Remark 13.13 (Tightening via a hybrid $\ell_{\infty}/\ell_{2}$ guard).
A practical optimization is:
  1. 1.
    Compute a cheap exponent $k_{\infty}$ from $\|x\|_{\infty}$ (one pass, exact power-of-two scaling),
  2. 2.
    Apply $2^{-k_{\infty}}$ provisionally, then
  3. 3.
    Compute the guarded $\widehat{S}$ only if the provisional scaled vector is near the $\ell_{2}$ threshold (e.g., $\widehat{S}$ within a small constant factor of $1$).
This preserves one-sided safety (never under-normalize) while avoiding expensive norm estimation on most nodes and can improve feasibility by reducing unnecessary over-normalization.
This naturally leads to the following lemma:
Lemma 13.14 (Guarded normalization is gap-monotone).
Consider any recursion node where the local packing identity entrywise gives:
$Z~=~\beta U+T+L/\beta,$ (116)
and suppose the middle and tail bands $(T,L)$ are bilinear expressions in the normalized operands (as in the MM instantiation).
Let $A^{\star},B^{\star}$ be the ideal normalized operands (using the ideal exponent $k_{\mathrm{ideal}}$), and let $\widehat{A},\widehat{B}$ be the guarded normalized operands (using $\widehat{k}\geq k_{\mathrm{ideal}}$ as in Definition 13.3). Assume scaling is dyadic: each affected row/column is multiplied by an additional factor $2^{-d}$ with $d\in\mathbb{Z}_{\geq 0}$. The entrywise the guarded band magnitudes are then no larger than the ideal band magnitudes:
$|\widehat{T}|~\leq~|T^{\star}|,~|\widehat{L}|~\leq~|L^{\star}|.$ (117)
Consequently, any master gap inequality verified in the ideal analysis remains valid (or improves) under guarded normalization as long as all values remain in-range.
Proof.
Guarded normalization differs from ideal normalization only by applying additional dyadic downscalings: for each normalized row/column, $\widehat{k}=k_{\mathrm{ideal}}+d$ with $d\geq 0$, so the guarded operand equals the ideal operand multiplied by $2^{-d}$ on that row/column.
Each entry of $T$ and $L$ is a sum of dot products of rows of the left operand with columns of the right operand. If a row is additionally scaled by $2^{-d_{1}}$ and the corresponding column by $2^{-d_{2}}$, then the dot product is scaled by $2^{-(d_{1}+d_{2})}\leq 1$. Therefore each summand contributing to $\widehat{T}$ (resp., $\widehat{L}$) has magnitude at most the corresponding summand in $T^{\star}$ (resp., $L^{\star}$), and summing preserves the inequality, so $|\widehat{T}|\leq|T^{\star}|$ and $|\widehat{L}|\leq|L^{\star}|$ entrywise.
The master gap inequality depends monotonically on $|T|$ and $|L|$ (it is a sum of nonnegative terms involving these magnitudes), so decreasing $|T|$ and $|L|$ cannot break the inequality. ∎
Corollary 13.15.
If the master gap holds with slack $\delta$ for the ideal normalized operands, then it holds with slack at least $\delta$ for the guarded normalized operands.
14 Float-GPR complexity
Theorem 14.1 (Float-GPR kernel complexity reduces to the same recurrence).
Let $T_{\mathsf{F}}(n)$ be the runtime of the float-GPR matrix multiplication kernel on $n\times n$ inputs in format $\mathsf{F}$, counting each scalar operation in $\mathsf{F}$ as $O(1)$ time and treating graded projection (centering/normalization) as: streaming passes, and assume the kernel:
  • makes exactly $4$ recursive calls on subproblems of size $n/2$ per level (one per quadrant), and
  • performs $O(n^{2})$ additional scalar work per level for projection, packing/unpacking, and extraction bookkeeping,
then:
$T_{\mathsf{F}}(n)~=~4T_{\mathsf{F}}(n/2)+O(n^{2}),$ (118)
so $T_{\mathsf{F}}(n)=\Theta(n^{2}\log n)$ in the unit-cost $\mathsf{F}$-operation model.
If extended precision is disallowed and the implementation emulates $b(n)$-bit intermediates using native $w$-bit words via digit-band (slice) staging, then the overhead factor is:
$K(n)~:=~\left\lceil\frac{b(n)}{w}\right\rceil,$ (119)
so the resulting runtime is:
$T_{\mathsf{F}}^{(\text{no-ext})}(n)~=~\Theta\!\left(n^{2}\log n\cdot\Big\lceil\frac{b(n)}{w}\Big\rceil\right).$ (120)
Here $b(n)$ denotes the required mantissa precision (in bits) for the intermediates under the chosen execution discipline. In fixed-format commensurate regimes, $b(n)=O(1)$ is just the format precision; when higher precision (or exact staged execution) is required across depth, one typically has $b(n)=\Theta(\log n)$ in the basic packed recursion. In particular, if $b(n)=\Theta(\log n)$ and $w$ is fixed, then:
$T_{\mathsf{F}}^{(\text{no-ext})}(n)$ $=$ $\Theta\!\left(n^{2}\log n\cdot\lceil\log n/w\rceil\right)$ (121)
$=$ $\Theta(n^{2}\log^{2}n)~\text{(for fixed $w$).}$ (122)
More generally, the staging overhead is exactly the factor $\lceil b(n)/w\rceil$ appearing above, i.e., no additional hidden polylog factor is being introduced here beyond what is already explicit.
Remark 14.2 (Why the float-GPR complexity claim does not rely on exact extraction).
The near-quadratic bound in float-GPR is a structural statement: it follows from the $2\times 2$ recursion shape (4 subcalls on size $n/2$ per level) and $O(n^{2})$ nonrecursive work per level (projection, packing, extraction bookkeeping). It does not require formally exact dyadic-lattice extraction. In float mode, the extractor is analyzed through decision-site invariance under a drift ledger (Part II), while Part I provides an exact dyadic-lattice specification as a trust anchor.
Proof.
The recurrence follows directly from the stated structure: $4$ recursive calls of size $n/2$ plus $O(n^{2})$ nonrecursive work per level, so the Master Theorem (critical case) gives $\Theta(n^{2}\log n)$.
If extended precision is disallowed and intermediates of required mantissa precision $b(n)$ bits are emulated using native $w$-bit words via digit-band (slice) staging, then each staged multiply/extraction site is carried out in:
$K(n)~:=~\left\lceil\frac{b(n)}{w}\right\rceil$ (123)
passes. This multiplies the per-level nonrecursive work by at most the same factor (and likewise any fixed constant number of recursive kernel calls), so the overall runtime is scaled by $K(n)$:
$T_{\mathsf{F}}^{(\text{no-ext})}(n)$ $=$ $\Theta\!\left(T_{\mathsf{F}}(n)\cdot K(n)\right)$ (124)
$=$ $\Theta\!\left(n^{2}\log n\cdot\Big\lceil\frac{b(n)}{w}\Big\rceil\right).$ (125)
Remark 14.3.
In the $\ell_{2}$-normalized regime, $b(n)=O(1)$ implies $K(n)=O(1)$, so staging does not change the asymptotic $O(n^{2}\log n)$ bound.
Corollary 14.4 ($\ell_{2}$-normalized refinement removes staging overhead).
If recursive graded projection maintains a dimension-free dot-product envelope so that the required mantissa precision is bounded by a constant $b(n)=O(1)$ (independent of $n$), then $\lceil b(n)/w\rceil=O(1)$ and therefore:
$T_{\mathsf{F}}^{(\text{no-ext})}(n)~=~\Theta(n^{2}\log n).$ (126)
Part III Case studies and templates (Part III)
The Graded Projection Recursion (GPR) framework provides a set of tools for reducing the complexity of certain algorithms that satisfy certain conditions. One example is GPR reduction of the complexity of general $n\times n$ matrix multiplication (MM) to $O(n^{2}\log n)$. It would be natural to assume that this automatically implies a complexity reduction for all algorithms with complexity reducible to MM, but that is not the case.
Why $\boldsymbol{{\large\omega}_{\text{\tiny GPR}}}{=}$2 alone is not enough.
Classical algebraic reductions assert that computing determinants, inverses, or solving systems takes $O(\text{MM}(n))$ arithmetic operations under unit-cost unbounded precision. GPR’s model-honest $\tilde{O}(n^{2})$ bit/word cost for matrix multiplication does not automatically transfer via those reductions because classical reductions do not control bit-length growth. This means that separate GPR reductions must be established.
In the following sections we review the GPR framework and apply it to achieve reduced complexities for fundamental linear algebra tasks (e.g., factorizations, matrix functions, and spectral decompositions) to obtain algorithms with soft-quadratic complexity ($\tilde{O}(n^{2})$), i.e., $n^{2}$ times only polylogarithmic overhead factors that are made explicit (e.g., from recursion depth and any required staged precision). Importantly, this efficiency is achieved while maintaining backward and forward error commensurate to their corresponding classical $O(n^{3})$ blocked counterparts on standard success sets. We regard these example reductions as templates that can be applied in other cases.
15 A reusable GPR reduction template
15.1 The GPR recipe (informal)
  1. 1.
    Express the target algorithm in a blocked/recursive form.
  2. 2.
    Identify the GEMM-dominant updates (Schur complements, trailing updates, rank-$k$ updates, normal equations, etc.).
  3. 3.
    Replace each dense product with the appropriate GPR kernel (integral KernelMM or float KernelMM_F).
  4. 4.
    Establish and maintain node-local invariants (graded projection) so that intermediate budgets remain bounded and (in float mode) drift is ledger-bounded.
  5. 5.
    Conclude a model-honest complexity bound by summing per-level work, and conclude correctness/commensurate accuracy from the kernel guarantees.
15.2 Formal template: integral mode
Theorem 15.1 (Integral GPR reduction theorem (template)).
Let $\mathcal{T}$ be a task on $n\times n$ inputs specified by algebraically correct block identities over $\mathbb{R}$. Assume $\mathcal{T}$ is implemented by a $2\times 2$ blocked recursion (on powers of two, with padding otherwise) with:
  1. 1.
    Exact block algebra: All block identities used by the algorithm are algebraically correct over $\mathbb{R}$ (thus also over the exact dyadic semantics used as a reference in Part I).
  2. 2.
    GEMM-dominant structure: At every recursion node of dimension $m$, the algorithm performs:
    • exactly $4$ recursive subcalls on subproblems of dimension $m/2$, and
    • at most a constant number $q=O(1)$ of calls to a GPR matrix-multiplication primitive on $(m/2)\times(m/2)$ blocks (integral KernelMM in exact staged mode, or its float analogue in commensurate mode),
    plus $O(m^{2})$ additional bookkeeping that is exact in the chosen semantics (or comparator-faithful in float mode).
  3. 3.
    Kernel cost model: Let $B_{\mathrm{MM}}(m)$ denote the bit/word cost of the chosen GPR MM primitive on $m\times m$ inputs under the declared execution regime (fixed-format or staged-precision),
then:
  1. 1.
    Correctness/accuracy: In integral staged mode, $\mathcal{T}$ is exact on the declared success regime (the same regime on which the primitive kernels are exact). In float mode, $\mathcal{T}$ is commensurate-accurate to its baseline comparator under the drift-ledger conditions (Theorem 15.3).
  2. 2.
    Complexity: The bit/word cost satisfies:
    $B_{\mathcal{T}}(n)~=~\widetilde{O}\!\left(B_{\mathrm{MM}}(n)\right),$ (127)
    i.e., the task is soft-quadratic whenever the substituted MM primitive is soft-quadratic. Concretely, if $B_{\mathrm{MM}}(n)=\Theta(n^{2}\log n\cdot K(n))$ under slice staging (Theorem 9.3), then $B_{\mathcal{T}}(n)=\tilde{O}(n^{2})$ as well (up to additional polylog factors already made explicit in the algorithm’s structure).
Proof.
Correctness/commensurate accuracy follows by composition: every dense product (or factorization/orthogonalization when applicable) is replaced by a primitive kernel with the stated guarantee, and all remaining steps are algebraically correct (integral mode) or comparator-faithful (float mode), with decision-site invariance ensured by the drift ledger.
For complexity, let $B_{\mathcal{T}}(m)$ denote the cost on dimension $m$. At a node of size $m$, there are $4$ subcalls of size $m/2$; a constant number of MM-kernel invocations on size $m/2$ blocks; and $O(m^{2})$ additional work, so:
$B_{\mathcal{T}}(m)~=~4B_{\mathcal{T}}(m/2)~+~O\!\left(B_{\mathrm{MM}}(m)\right),$ (128)
where the $O(\cdot)$ hides only fixed constants (including the constant number of kernel calls per node). If $B_{\mathrm{MM}}(m)$ is soft-quadratic, then the solution is soft-quadratic as well, giving $B_{\mathcal{T}}(n)=\widetilde{O}(B_{\mathrm{MM}}(n))$. ∎
15.3 Formal template: float mode
Definition 15.2 (GPR-reducible (float mode)).
Fix an execution format $\mathsf{F}$ and baseline comparators for the relevant primitives (Part II). A task $\mathcal{T}$ is GPR-reducible in float mode if it admits a blocked/iterative structure in which:
  • All dense products and dense factorizations/orthogonalizations can be replaced by float-GPR primitives (e.g., KernelMM_F, LDLT_F, BlockedQR_F),
  • All discrete decisions in the task (extraction, pivoting thresholds, orthogonality tests, etc.) have explicit slack $\delta$ and satisfy the drift-ledger inequality, and
  • The resulting algorithm is commensurate-accurate to the baseline comparator on the baseline-feasible regime.
Theorem 15.3 (Float GPR reduction theorem (template)).
Fix an execution format $\mathsf{F}$ (Definition 12.1) and baseline comparators for the primitive kernels that appear in the task (e.g., $\mathrm{BaseMM}_{\mathsf{F}}$ for dense products, and analogous baseline comparators for LDL/QR if used). Let $\mathcal{T}$ be a task implemented by a blocked/iterative algorithm $\mathcal{A}_{\mathrm{base}}$ in format $\mathsf{F}$.
Assume $\mathcal{T}$ admits a float-GPR implementation $\mathcal{A}_{\mathrm{gpr}}$ obtained by the following transformation:
  1. 1.
    Kernel substitution: Every dense GEMM (and, when applicable, every dense factorization/orthogonalization) in $\mathcal{A}_{\mathrm{base}}$ is replaced by its float-GPR analogue (e.g., KernelMM_F, LDLT_F, BlockedQR_F), each of which is commensurate-accurate to its baseline comparator.
  2. 2.
    Graded projection invariants: At every recursion node / block update / iteration, $\mathcal{A}_{\mathrm{gpr}}$ enforces node-local graded projection steps that provide explicit magnitude envelopes for the operands and intermediate quantities entering each discrete decision site.
  3. 3.
    Ledger-bounded drift at decision sites: All discrete decision sites (Definition 12.8) in the substituted kernels and in the surrounding algorithm satisfy the drift-ledger inequality $\varepsilon_{\mathrm{tot}}\leq\delta$ (Definition 12.9), so that decisions are invariant (Lemma 12.10).
The following hold under these assumptions:
  1. 1.
    Commensurate feasibility: On the baseline-feasible regime (i.e., whenever the baseline execution of $\mathcal{A}_{\mathrm{base}}$ produces finite values in $\mathsf{F}$); the float-GPR implementation $\mathcal{A}_{\mathrm{gpr}}$ produces finite values; and because graded projection bounds intermediates, $\mathcal{A}_{\mathrm{gpr}}$ can be defined on a strictly larger feasible regime (it may avoid spurious overflow/underflow where baseline fails).
  2. 2.
    Commensurate accuracy (conservative): On the baseline-feasible regime, $\mathcal{A}_{\mathrm{gpr}}$ is commensurate-accurate to $\mathcal{A}_{\mathrm{base}}$ in the sense of Definition 12.7: (1) discrete choices (extraction, pivoting thresholds, etc.) match the ideal decisions for the GPR kernels,and (2) the remaining numerical error is bounded by the same explicit drift terms budgeted in the ledger (no additional error from wrong decisions).
  3. 3.
    Complexity: If the blocked/iterative structure of $\mathcal{A}_{\mathrm{base}}$ is GEMM-dominant and uses only $O(1)$ dense-kernel calls per recursion node (or per iteration), then the asymptotic cost of $\mathcal{A}_{\mathrm{gpr}}$ is the corresponding cost of the substituted kernels plus $O(n^{2})$ graded-projection passes per level/iteration. In particular, for a depth-$O(\log n)$ blocked recursion with a constant number of kernel calls per node:
    $T_{\mathrm{gpr}}(n)=4T_{\mathrm{gpr}}(n/2)+O(n^{2})~\Rightarrow~T_{\mathrm{gpr}}(n)=\Theta(n^{2}\log n)$ (129)
    in the unit-cost $\mathsf{F}$-operation model.
Proof.
  1. 1.
    Feasibility: Each float-GPR primitive kernel is assumed commensurate-accurate to its baseline comparator, in particular preserving feasibility on the baseline-feasible regime. The surrounding algorithm adds only graded projection steps and bookkeeping that are themselves executed in $\mathsf{F}$. Because graded projection is designed to bound intermediate magnitudes before they enter dense products/factorizations, it cannot make an otherwise finite baseline execution overflow by introducing larger intermediates than baseline. Instead, it tends to reduce intermediate growth. Therefore, whenever $\mathcal{A}_{\mathrm{base}}$ is finite, $\mathcal{A}_{\mathrm{gpr}}$ is also finite. Strict enlargement of the feasible region follows in regimes where baseline fails due to intermediate growth but graded projection keeps intermediates within range.
  2. 2.
    Commensurate accuracy: Consider any discrete decision site in $\mathcal{A}_{\mathrm{gpr}}$ (either inside a substituted kernel or in surrounding logic). By assumption, the ideal decision statistic has slack at least $\delta$ and the total drift is bounded by $\varepsilon_{\mathrm{tot}}\leq\delta$. Therefore, Lemma 12.10 applies and the implemented decision equals the ideal decision, so the algorithm does not incur catastrophic error due to wrong discrete choices (wrong extraction, wrong pivot threshold crossing, etc.). All remaining numerical discrepancy relative to exact real arithmetic is accounted for by the drift terms already budgeted in the ledger, which is exactly the conservative commensurate-accuracy guarantee of Definition 12.7.
  3. 3.
    Complexity: Kernel substitution replaces each dense GEMM/factorization in the baseline by a corresponding float-GPR kernel. Graded projection adds $O(n^{2})$ streaming work per recursion node (or per iteration) to compute/maintain envelopes and to perform normalization. Under the stated GEMM-dominant blocked structure (constant number of kernel calls per node), the resulting recurrence is $4T(n/2)+O(n^{2})$ for a $2\times 2$ recursion, giving $\Theta(n^{2}\log n)$ by the Master Theorem.
16 Case studies: linear algebra
This section instantiates the template for core dense linear algebra tasks. Each subsection contains: (1) blocked algorithm skeleton, (2) where GEMM appears, (3) the GPR replacement, and (4) the near-quadratic complexity theorem.
Lemma 16.1 (LA case studies are GEMM-dominant instantiations).
Each linear-algebra case study in Section 16 is implemented by a blocked recursion or blocked iteration in which:
  1. 1.
    The dominant work consists of a constant number of dense products (GEMM-like updates) per recursion node/iteration, and
  2. 2.
    All remaining work is $O(n^{2})$ per level/iteration (solves, permutations, panel work).
Therefore, once the appropriate GPR kernel primitive is substituted, the task complexity follows from the corresponding reduction theorem: Theorem 15.1 (integral mode) or Theorem 15.3 (float mode).
Proof.
This is a structural statement: each listed algorithm’s blocked form is standard and its dense updates are GEMM-dominant. The cited reduction theorems then apply directly. ∎
Remark 16.2 (No hidden cubic subroutines).
All linear-algebra case studies are stated for blocked forms with a fixed cutoff or block size. Under that convention, all non-kernel work is $O(n^{2})$ per recursion level or per iteration, and the dominant work is carried by the substituted kernels.
16.1 LUP, Solve, Determinant, Inverse (integral mode)
Theorem 16.3 (Near-quadratic LUP/Solve/Det/Inv via GPR).
Let $A\in\mathcal{L}_{D}^{n\times n}$ with $n$ a power of two. Consider a blocked $2\times 2$ LUP recursion that at each node:
  • computes an LUP factorization of the leading block $A_{11}$ (recursively),
  • performs two blocked triangular solves to form $U_{12}=L_{11}^{-1}A_{12}$ and $L_{21}=A_{21}U_{11}^{-1}$ (recursively / with a cutoff),
  • forms the Schur complement
    $S~:=~A_{22}-L_{21}U_{12},$ (130)
    and recurses on $S$.
Assume the only dense products appearing in the Schur updates and in the trailing updates of the blocked triangular solves are computed using KernelMM, and that all remaining operations (permutations, triangular/diagonal solves, and exponent-offset updates) are exact in dyadic lattice semantics and preserve the fixed-word model, then:
  1. 1.
    The resulting algorithm computes an exact LUP factorization of $A$ in dyadic lattice semantics on the domain where the chosen pivots are nonzero.
  2. 2.
    The induced routines for solve, determinant, and inverse are exact in dyadic lattice semantics:
    • Solve: for any $b$, the algorithm returns the exact solution of $Ax=b$ (when it exists under the pivoting policy).
    • Determinant: $\det(A)=\mathrm{sgn}(P)\prod_{i}U_{ii}$ is computed exactly.
    • Inverse: $A^{-1}$ is computed exactly via solves with multiple right-hand sides (when $A$ is nonsingular).
  3. 3.
    The bit/word complexity is soft-quadratic (polylog overhead made explicit):
    $B_{\mathrm{LUP}}(n)~=~\tilde{O}(n^{2}),$ (131)
    and likewise for Solve/Det/Inv.
Lemma 16.4 (Pivot-invariance of centering).
Let $A\in\mathbb{R}^{m\times m}$ and let $\alpha:=\frac{1}{m}A^{\top}\mathbf{1}$ be its column-mean vector, so that $A_{0}:=A-\mathbf{1}\alpha^{\top}$ satisfies $A_{0}^{\top}\mathbf{1}=0$. Let $P$ be any permutation matrix (row permutation), then $PA_{0}$ is the centered form of $PA$, i.e., if $(PA)_{0}$ denotes the centered matrix obtained from $PA$ using its column means, then:
$(PA)_{0}~=~PA_{0}.$ (132)
Equivalently,
$P(A-\mathbf{1}\alpha^{\top})~=~PA-\mathbf{1}\alpha^{\top}.$ (133)
Proof.
Because $P$ is a row permutation, $P\mathbf{1}=\mathbf{1}$ and $P^{\top}\mathbf{1}=\mathbf{1}$. The column-mean vector of $PA$ is therefore unchanged:
$\frac{1}{m}(PA)^{\top}\mathbf{1}~=~\frac{1}{m}A^{\top}P^{\top}\mathbf{1}~=~\frac{1}{m}A^{\top}\mathbf{1}~=~\alpha.$ (134)
Using $P\mathbf{1}=\mathbf{1}$,
$(PA)_{0}~=~PA-\mathbf{1}\alpha^{\top}~=~PA-P(\mathbf{1}\alpha^{\top})~=~P(A-\mathbf{1}\alpha^{\top})~=~PA_{0},$ (135)
as claimed. ∎
Remark 16.5 (Pivoting bookkeeping in the LUP reduction).
In blocked LUP with partial pivoting, each pivot step applies a row permutation to the active block (and to the corresponding rows of the trailing blocks). Lemma 16.4 shows that the column-centering discipline used by KernelMM is compatible with these permutations: The centered representation does not need to be recomputed simply because a pivot swap occurred. Any auxiliary per-row metadata (e.g., stored row-exponent shifts, pivot history, or row-wise drift ledgers in float mode) is updated by applying the same permutation to the corresponding length-$m$ vectors, which is negligible compared to the $O(m)$ cost of swapping the affected matrix rows and the $O(m^{2})$ dense updates at the node.
Proof.
Correctness (LUP)
Blocked LUP is algebraically correct: given an exact LUP of $A_{11}$ and exact triangular solves producing $U_{12}$ and $L_{21}$, the Schur complement identity:
$A~=~\begin{pmatrix}I&0\\ L_{21}U_{11}^{-1}&I\end{pmatrix}\begin{pmatrix}A_{11}&A_{12}\\ 0&S\end{pmatrix}$ (136)
(with appropriate permutations) holds exactly over $\mathbb{R}$ and therefore over $\mathcal{L}_{D}$. Thus, if the Schur complement $S=A_{22}-L_{21}U_{12}$ is computed exactly, and the recursion returns an exact factorization of $S$, then assembling the block factors gives an exact factorization of $A$.
In our implementation, the only nontrivial dense arithmetic in forming $S$ is the matrix product $L_{21}U_{12}$, which is computed exactly by KernelMM. All remaining operations (permutations and triangular/diagonal solves) are exact by hypothesis. Therefore, by induction on $n$, the recursion returns an exact LUP factorization of $A$.
Correctness (Solve/Det/Inv)
Given an exact factorization $PA=LU$, standard forward/back substitution produces the exact solution $x$ to $Ax=b$ (when it exists and pivots are nonzero). The determinant formula $\det(A)=\mathrm{sgn}(P)\prod_{i}U_{ii}$ follows directly from triangularity and is exact under exact arithmetic. The inverse is obtained by solving $AX=I$ column-by-column (or in blocks) using the same exact solves, so Solve/Det/Inv are exact.
Bit complexity
By assumption, the recursion is $2\times 2$ blocked and GEMM-dominant, so the Schur complement update at each node uses a constant number of KernelMM calls on size $n/2$ blocks with $O(n^{2})$ additional exact work (solves, permutations, streaming passes). This is precisely the setting of the integral GPR reduction theorem (template), Theorem 15.1. Therefore:
$B_{\mathrm{LUP}}(n)~=~\tilde{O}(n^{2}).$ (137)
The same asymptotic bound holds for Solve/Det/Inv because they reuse the same factorization and perform only $O(n^{2})$ additional work (forward/back solves and diagonal products) beyond the factorization. ∎
16.2 LDL (integral mode) and Cholesky (float mode)
Theorem 16.6 (Near-quadratic LDL via GPR).
Let $A\in\mathcal{L}_{D}^{n\times n}$ be symmetric, with $n$ a power of two. Consider a blocked $2\times 2$ LDL recursion that at each node:
  • Computes an LDL factorization of the leading block $A_{11}$ (recursively),
  • Solves for $L_{21}$ via exact triangular and diagonal solves:
    $L_{21}~:=~A_{21}L_{11}^{-\top}D_{11}^{-1},$ (138)
  • forms the Schur complement update
    $S~:=~A_{22}-L_{21}D_{11}L_{21}^{\top},$ (139)
    and recurses on $S$.
Assume that every dense product appearing in the Schur updates (and in any trailing updates of the blocked solves) is computed using KernelMM, and that all remaining operations (symmetry bookkeeping, permutations/pivoting if used, diagonal/triangular solves, and exponent-offset updates) are exact in dyadic lattice semantics and preserve the fixed-word model, then:
  1. 1.
    The resulting algorithm computes an exact blocked LDL factorization of $A$ in dyadic lattice semantics on the domain where the chosen pivots (or pivot blocks) are nonsingular.
  2. 2.
    The bit/word complexity is soft-quadratic (polylog overhead made explicit):
    $B_{\mathrm{LDL^{\top}}}(n)~=~\tilde{O}(n^{2}).$ (140)
Remark 16.7 (Symmetric pivoting and permutation invariance).
If the LDL recursion uses symmetric pivoting, the pivot step applies a symmetric permutation $A\leftarrow PAP^{\top}$ on the active submatrix. Because $P\mathbf{1}=\mathbf{1}$ and $\mathbf{1}^{\top}P^{\top}=\mathbf{1}^{\top}$ for any permutation matrix, any centering discipline of the form $A_{0}=A-\mathbf{1}\alpha^{\top}$ (and likewise two-sided variants) is maintained equivariantly under the pivot permutation by applying the same permutation to the associated metadata. Thus symmetric pivoting does not force re-centering from scratch; the extra bookkeeping is dominated by the standard row/column swap work and the $O(m^{2})$ dense updates at a node of size $m$.
Proof.
Correctness
Blocked LDL is algebraically correct: given an exact factorization of $A_{11}=L_{11}D_{11}L_{11}^{\top}$ and exact solves producing $L_{21}:=A_{21}L_{11}^{-\top}D_{11}^{-1}$, the Schur complement identity:
$A~=~\begin{pmatrix}I&0\\ L_{21}&I\end{pmatrix}\begin{pmatrix}A_{11}&0\\ 0&S\end{pmatrix}\begin{pmatrix}I&L_{21}^{\top}\\ 0&I\end{pmatrix}$ (141)
holds exactly (with appropriate interpretation of $S$), where:
$S~=~A_{22}-L_{21}D_{11}L_{21}^{\top}.$ (142)
Thus, if $S$ is formed exactly and the recursion returns an exact factorization of $S$, then assembling the blocks gives an exact factorization of $A$.
In our implementation, the only nontrivial dense arithmetic in forming $S$ is the product $L_{21}D_{11}L_{21}^{\top}$. This can be computed exactly by a constant number of KernelMM calls (for diagonal $D_{11}$, the multiply by $D_{11}$ is entrywise scaling). All remaining operations are exact by hypothesis. Therefore, by induction on $n$, the recursion returns an exact LDL factorization of $A$.
Bit complexity
At each node, the algorithm performs:
  • four recursive calls on size $n/2$ blocks (the standard $2\times 2$ recursion structure), and
  • a constant number of KernelMM calls on size $n/2$ blocks to form the Schur update, plus
  • $O(n^{2})$ additional exact work (triangular/diagonal solves, permutations/pivot bookkeeping, and streaming passes).
This matches the hypotheses of the integral GPR reduction theorem (template), Theorem 15.1, so:
$B_{\mathrm{LDL^{\top}}}(n)~=~\tilde{O}(n^{2}).$ (143)
Remark 16.8.
Cholesky requires square roots. A float-GPR version is the appropriate case study for a practical Cholesky implementation.
16.3 Blocked QR (float or integral mode, depending on semantics)
Theorem 16.9 (Near-quadratic blocked QR trailing updates via GPR).
Fix an execution format $\mathsf{F}$ and a baseline blocked QR comparator in $\mathsf{F}$ (Part II). Consider a blocked QR factorization in WY form, where each panel step applies a block reflector:
$Q_{\mathrm{blk}}~=~I-YTY^{\top}$ (144)
to a trailing matrix $C$ via the update:
$C\leftarrow C-Y\left(T^{\top}(Y^{\top}C)\right).$ (145)
Assume:
  1. 1.
    The two GEMM-dominant products $Y^{\top}C$ and $Y(\cdot)$ are implemented using a float-GPR multiply primitive KernelMM_F that is commensurate-accurate to the baseline GEMM comparator and enforces bounded intermediates via graded projection.
  2. 2.
    The remaining panel work (Householder construction, forming $T$, and multiplying by the small triangular $T^{\top}$) is performed in $\mathsf{F}$ and is baseline-comparable (and does not dominate asymptotically),
then:
  1. 1.
    Commensurate accuracy. The resulting blocked QR is commensurate-accurate to the baseline blocked QR comparator on the baseline-feasible regime.
  2. 2.
    Near-quadratic complexity (kernel-dominant view). The total cost of all trailing updates across the factorization is $\widetilde{O}(T_{\mathrm{MM}}(n))$. In particular, if KernelMM_F has complexity $\Theta(n^{2}\log n)$ in the unit-cost $\mathsf{F}$-operation model, then the trailing-update cost is $\Theta(n^{2}\log n)$ up to lower-order $O(n^{2})$ panel overhead.
Proof.
Commensurate accuracy
Each trailing update is a composition of two dense products and several lower-dimensional operations:
$W$ $:=$ $Y^{\top}C,$ (146)
$U$ $:=$ $T^{\top}W,$ (147)
$C$ $\leftarrow$ $C-YU.$ (148)
By assumption, the two dense products $Y^{\top}C$ and $YU$ are computed by KernelMM_F, which is commensurate-accurate to the baseline GEMM comparator, and whose graded projection bounds intermediates and localizes decision drift. The remaining steps (forming $T$, triangular multiply by $T^{\top}$, and the subtraction) are standard $\mathsf{F}$ operations under the same comparator semantics. Therefore each trailing update is commensurate-accurate to the baseline trailing update on the baseline-feasible regime. Composing these updates over the panels preserves commensurate accuracy of the full blocked QR factorization.
Complexity of trailing updates
In blocked QR, the dominant work is in the trailing updates: the two GEMM products per panel step. Summing across all panels gives a total GEMM-equivalent workload $\Theta(n^{3})$ in the classical model, but the GPR accounting measures it as a constant number of calls to a blocked multiplication primitive per recursion node, plus $O(n^{2})$ streaming work. Under the hypothesis that KernelMM_F runs in $\Theta(n^{2}\log n)$ time and that each panel step invokes only a constant number of such kernel calls on subblocks, the total cost of all trailing updates is $\widetilde{O}(T_{\mathrm{MM}}(n))$ and in particular $\Theta(n^{2}\log n)$. The remaining panel work is $O(n^{2})$ per factorization (for fixed block size) and is therefore lower order. ∎ This can also be proven using purely integral (dyadic lattice semantics), but in practice QR is best framed in the float-GPR setting because orthogonalization is inherently floating and stability-driven.
16.4 SOI and SPD matrix functions (float mode)
Theorem 16.10 (SOI with GPR kernels: per-iteration near-quadratic cost).
Fix an execution format $\mathsf{F}$ and a baseline comparator for the following SOI iteration on a symmetric matrix $A\in\mathbb{R}^{n\times n}$:
$Y_{k}$ $:=$ $AX_{k},$ (149)
$X_{k+1}$ $:=$ $\mathrm{qf}(Y_{k}),$ (150)
where $\mathrm{qf}(\cdot)$ denotes an orthonormalization step (e.g., blocked QR) and $X_{k}\in\mathbb{R}^{n\times r}$.
Assume the GPR implementation replaces:
[…]
  • the orthonormalization $\mathrm{qf}(Y_{k})$ by a blocked QR routine whose trailing GEMM updates use KernelMM_F (as in Theorem 16.9),
and that these float-GPR primitives are commensurate-accurate to their baseline comparators in $\mathsf{F}$, then:
  1. 1.
    (Per-iteration cost.) Each SOI iteration costs $\widetilde{O}(T_{\mathrm{MM}}(n,r))$, i.e., it is dominated by a constant number of GPR GEMM calls on $n\times n$ by $n\times r$ (or $n\times r$ by $r\times n$) blocks plus $O(nr)$ streaming work. In particular, for fixed $r$ and a near-quadratic GPR multiply primitive, the per-iteration cost is $\Theta(n^{2}\log n)$ in the unit-cost $\mathsf{F}$ model.
  2. 2.
    (Total cost over $K$ iterations.) For $K$ SOI iterations,
    $T_{\mathrm{SOI}}(n,r,K)~=~K\cdot\widetilde{O}(T_{\mathrm{MM}}(n,r)).$ (151)
Proof.
Fix an iteration $k$.
  1. 1.
    Multiply step: Computing $Y_{k}=AX_{k}$ is one matrix–block product. By assumption it is performed by KernelMM_F, so its cost is $\widetilde{O}(T_{\mathrm{MM}}(n,r))$.
  2. 2.
    Orthonormalization step: Computing $X_{k+1}=\mathrm{qf}(Y_{k})$ via a blocked QR routine consists of panel work plus trailing updates. The trailing updates are GEMM-dominant and, by Theorem 16.9, their cost is $\widetilde{O}(T_{\mathrm{MM}}(n,r))$ for fixed block size, with $O(nr)$ additional panel overhead.
  3. 3.
    Per-iteration bound: Adding the multiply step and the QR step gives a constant number of GPR GEMM calls plus lower-order streaming work, so $\widetilde{O}(T_{\mathrm{MM}}(n,r))$ per iteration. For fixed $r$, this specializes to $\Theta(n^{2}\log n)$ in the unit-cost $\mathsf{F}$-operation model.
Total over $K$ iterations.
The SOI loop performs $K$ such iterations, so the total cost is $K$ times the per-iteration cost. ∎
16.5 Generalized symmetric EVP (float mode)
Theorem 16.11 (GSEVP via GPR reductions).
Fix an execution format $\mathsf{F}$. Let $A,B\in\mathbb{R}^{n\times n}$ be symmetric with $B\succ 0$. Consider the generalized symmetric eigenvalue problem (GSEVP):
$Ax~=~\lambda Bx.$ (152)
Assume we implement the standard reduction:
  1. 1.
    Compute a (float) Cholesky-type factorization $B=LL^{\top}$ (or a stabilized LDL variant) in format $\mathsf{F}$,
  2. 2.
    Form the reduced symmetric matrix:
    $C~:=~L^{-1}AL^{-\top},$ (153)
  3. 3.
    Solve the symmetric EVP for $C$ using SOI/QR iteration, and back-transform eigenvectors.
Assume further that:
  • All dense products appearing in forming $C$ (and in any blocked triangular solve trailing updates) are performed with KernelMM_F,
  • The symmetric EVP step is implemented by SOI with GPR kernels as in Theorem 16.10,
  • The GPR primitives are commensurate-accurate to their baseline comparators and enforce bounded intermediates via graded projection.
With these assumptions we can deduce:
  1. 1.
    (Commensurate accuracy) The resulting GPR-GSEVP pipeline is commensurate-accurate to the baseline GSEVP pipeline (on the baseline-feasible regime), and can enlarge feasibility by avoiding spurious overflow/underflow in forming/factoring $B$ and $C$.
  2. 2.
    (Near-quadratic complexity) The total cost satisfies
    $T_{\mathrm{GSEVP}}(n,K)=\widetilde{O}(T_{\mathrm{MM}}(n))+K\cdot\widetilde{O}(T_{\mathrm{MM}}(n)),$ (154)
    where $K$ is the iteration count of the chosen symmetric EVP routine (e.g., SOI). In particular, for fixed $K$ (or for focusing on per-iteration cost), the pipeline is GEMM-dominant with the same near-quadratic scaling as the GPR multiply primitive.
Proof.
Accuracy/feasibility
Each stage (factorization of $B$, formation of $C$, symmetric EVP solver, back-transformation) is a composition of primitive kernels (GEMM, triangular solves, QR/LDL) executed in format $\mathsf{F}$. By hypothesis, whenever a stage uses a dense product, it is computed by KernelMM_F, which is commensurate-accurate to the baseline GEMM comparator and bounds intermediates via graded projection. The same holds for the factorization and QR primitives used in the pipeline. Therefore, by the float GPR reduction theorem (template), Theorem 15.3, the composition is commensurate-accurate to the baseline pipeline on the baseline-feasible regime, with the possibility of strict feasibility enlargement due to bounded intermediates.
Complexity
Forming $C=L^{-1}AL^{-\top}$ can be implemented by blocked triangular solves and a constant number of GEMM-like updates, so its cost is $\widetilde{O}(T_{\mathrm{MM}}(n))$ under the GPR kernel substitution. Solving the symmetric EVP for $C$ by SOI/QR iteration costs $K$ iterations, each of cost $\widetilde{O}(T_{\mathrm{MM}}(n))$ by Theorem 16.10. The remaining steps (back-transforming eigenvectors, small dense operations on diagonal blocks) are lower order. Summing gives the stated bound. ∎
16.6 Dense LP and SDP (float mode)
Theorem 16.12 (IPM Schur complements via GPR kernels).
Fix an execution format $\mathsf{F}$ and baseline comparators for the dense kernels used in a primal-dual interior-point method (IPM) for dense LP or dense SDP. Assume the IPM iteration contains a dense Schur complement (or normal-equations) linear system:
$H_{k}\Delta y~=~r_{k}$ (155)
to be formed and solved at each iteration $k$, where:
  • For dense LP, $H_{k}=AD_{k}A^{\top}$ with $D_{k}$ diagonal.
  • For dense SDP, $H_{k}$ is an $m\times m$ Schur complement whose entries are trace products of scaled constraint matrices, and forming $H_{k}$ is GEMM-dominant when implemented densely.
Assume a float-GPR implementation replaces:
  1. 1.
    All dense products used to form $H_{k}$ by KernelMM_F, and
  2. 2.
    The factor/solve of $H_{k}$ by a commensurate-accurate stabilized symmetric solver (e.g., LDLT_F) whose trailing updates also call KernelMM_F,
and that these float-GPR primitives enforce bounded intermediates via graded projection and satisfy the drift-ledger conditions at their decision sites (Part II). With these assumptions we deduce:
  1. 1.
    (Commensurate accuracy) The resulting IPM linear-algebra step is commensurate-accurate to the baseline IPM linear-algebra step on the baseline-feasible regime, and can strictly enlarge feasibility by avoiding spurious overflow/underflow or breakdowns in forming/factoring $H_{k}$.
  2. 2.
    (Per-iteration near-quadratic kernel cost) The per-iteration cost of forming and solving the Schur system satisfies:
    $T_{\mathrm{Schur}}(n)~=~\widetilde{O}(T_{\mathrm{MM}}(n)),$ (156)
    where $n$ denotes the relevant dense dimension of the Schur system (e.g., $n=m$ for dense LP normal equations, or $n=\max\{m,\text{matrix size}\}$ for dense SDP). In particular, under a near-quadratic GPR multiply primitive, the per-iteration cost is $\Theta(n^{2}\log n)$ in the unit-cost $\mathsf{F}$-operation model.
  3. 3.
    (Total IPM cost.) Over $K$ IPM iterations:
    $T_{\mathrm{IPM}}(n,K)~=~K\cdot\widetilde{O}(T_{\mathrm{MM}}(n)),$ (157)
    up to lower-order work (residual formation, step-size rules, and vector updates).
Proof.
Accuracy/feasibility
Forming $H_{k}$ and solving $H_{k}\Delta y=r_{k}$ are compositions of dense products and a symmetric factor/solve. By hypothesis, each dense product is computed by KernelMM_F and the factor/solve by LDLT_F, both commensurate-accuracy to their baselines and equipped with graded projection and drift-ledger guarantees. Therefore, by the float GPR reduction theorem (template), Theorem 15.3, the linear-algebra step is commensurate-accurate to baseline on the baseline-feasible regime and may enlarge feasibility.
Complexity
For dense LP, forming $H_{k}=AD_{k}A^{\top}$ is two GEMM-like operations plus diagonal scaling, so $\widetilde{O}(T_{\mathrm{MM}}(m))$. Factoring/solving $H_{k}$ with blocked LDL is GEMM-dominant and likewise reduces to $\widetilde{O}(T_{\mathrm{MM}}(m))$ under kernel substitution. For dense SDP, forming $H_{k}$ is GEMM-dominant when expressed in dense linear algebra (scaling operator applications and trace products), and the subsequent factor/solve is again an $m\times m$ dense symmetric solve. Thus, in either case, per iteration the dominant cost is $\widetilde{O}(T_{\mathrm{MM}}(n))$ for the relevant dense dimension $n$. Summing across $K$ iterations gives the stated total cost. ∎
16.7 Gaussian process regression (float mode)
Theorem 16.13 (GP regression via GPR factorizations).
Fix an execution format $\mathsf{F}$ and a baseline comparator for dense SPD linear algebra in $\mathsf{F}$ (factorization, solve, and log-determinant). Let $x_{1},\dots,x_{n}\in\mathbb{R}^{d}$ be inputs, $y\in\mathbb{R}^{n}$ targets, and let $K\in\mathbb{R}^{n\times n}$ be a dense kernel matrix with entries $K_{ij}=k(x_{i},x_{j})$. Define the GP system matrix:
$A~:=~K+\sigma^{2}I$ (158)
for a noise/jitter parameter $\sigma^{2}>0$.
Consider the standard dense GP computations:
$\alpha~:=~A^{-1}y,~\log\det(A),$ (159)
(optionally followed by predictive mean/variance computations that reduce to additional solves and quadratic forms).
Assume a float-GPR implementation replaces:
  • The baseline factorization of $A$ (Cholesky or stabilized LDL) by a commensurate-accurate GPR factorization primitive (e.g., LDLT_F), whose trailing updates call KernelMM_F and enforce bounded intermediates via graded projection, and
  • All dense products inside the factorization and solves by KernelMM_F.
Assume these primitives satisfy the drift-ledger conditions at their decision sites (Part II), then:
  1. 1.
    (Commensurate accuracy) The resulting GP outputs $(\alpha_{\mathrm{gpr}},\log\det(A)_{\mathrm{gpr}})$ are commensurate-accurate to the baseline GP pipeline on the baseline-feasible regime, and can strictly enlarge feasibility by avoiding spurious overflow/underflow or breakdowns in the factorization.
  2. 2.
    (Near-quadratic kernel cost) The dominant cost of factorization+solve+logdet is:
    $T_{\mathrm{GP}}(n)~=~\widetilde{O}(T_{\mathrm{MM}}(n)).$ (160)
    In particular, if KernelMM_F has complexity $\Theta(n^{2}\log n)$ in the unit-cost $\mathsf{F}$-operation model, then the dense GP pipeline (factorization + one solve + logdet) has complexity $\Theta(n^{2}\log n)$ up to lower-order $O(n^{2})$ work.
Proof.
Accuracy/feasibility
The dense GP pipeline consists of (1) building $A$, (2) factoring $A$, (3) solving $A\alpha=y$, and (4) computing $\log\det(A)$ from the factors. By hypothesis, the factorization and solve primitives are commensurate-accurate to their baseline comparators and enforce bounded intermediates via graded projection, with drift-ledger guarantees at decision sites. Therefore, by the float GPR reduction theorem (template), Theorem 15.3, the composed pipeline is commensurate-accurate to baseline on the baseline-feasible regime and may enlarge feasibility.
Complexity
Dense factorization (Cholesky or stabilized LDL) is GEMM-dominant in blocked form: trailing updates reduce to a constant number of dense products per recursion node/level. With KernelMM_F substituted for those products, the factorization cost is $\widetilde{O}(T_{\mathrm{MM}}(n))$. The subsequent solve and log-determinant extraction are $O(n^{2})$ additional work beyond the factorization. Thus the full GP computation (factorization + one solve + logdet) is $\widetilde{O}(T_{\mathrm{MM}}(n))$. ∎
Corollary 16.14 (GP prediction cost (dense posterior mean and variance)).
In the setting of Theorem 16.13, let $X_{\ast}=\{x^{\ast}_{1},\dots,x^{\ast}_{t}\}$ be $t$ test points and define the cross-kernel matrix $K_{\ast}\in\mathbb{R}^{t\times n}$ by $(K_{\ast})_{\ell i}=k(x^{\ast}_{\ell},x_{i})$ and the test kernel diagonal $k_{\ast\ast}\in\mathbb{R}^{t}$ by $(k_{\ast\ast})_{\ell}=k(x^{\ast}_{\ell},x^{\ast}_{\ell})$.
Assume we have already computed a factorization of $A=K+\sigma^{2}I$ and $\alpha=A^{-1}y$, then:
  1. 1.
    The predictive mean vector
    $\mu_{\ast}~=~K_{\ast}\alpha$ (161)
    can be computed in time $\widetilde{O}(T_{\mathrm{MM}}(t,n))$. In particular, using a dense multiply primitive this is $O(tn)$ in the unit-cost model.
  2. 2.
    The predictive variance (diagonal) in the standard GP model,
    $\mathrm{Var}(f(x^{\ast}_{\ell})\mid X,y)~=~(k_{\ast\ast})_{\ell}-v_{\ell}^{\top}v_{\ell},~~\text{where}~v_{\ell}:=A^{-1/2}k_{\ell}^{\top},$ (162)
    can be computed by solving $AV^{\top}=K_{\ast}^{\top}$ for $V\in\mathbb{R}^{t\times n}$ (i.e., $t$ right-hand sides) and taking rowwise squared norms:
    $\sigma^{2}_{\ast}~=~k_{\ast\ast}-\mathrm{rowdiag}(VV^{\top}).$ (163)
    The solve with $t$ right-hand sides costs $O(tn^{2})$ beyond the factorization (or $\widetilde{O}(tT_{\mathrm{Solve}}(n))$ in a blocked model), and forming rowwise norms costs $O(tn)$.
Consequently, the total dense prediction cost after factorization is:
$\widetilde{O}\!\left(T_{\mathrm{MM}}(t,n)+tn^{2}\right),$ (164)
and the overall end-to-end dense GP training+prediction cost is:
$\widetilde{O}\!\left(T_{\mathrm{MM}}(n)+tn^{2}\right),$ (165)
where the first term accounts for factorization (Theorem 16.13).
Proof.
For the mean, $\mu_{\ast}=K_{\ast}\alpha$ is a dense matrix–vector (or matrix–block) product.
For the variances, write $V:=K_{\ast}A^{-1}$ (or equivalently solve $AV^{\top}=K_{\ast}^{\top}$). In a blocked factorization, solving with one right-hand side costs $O(n^{2})$ arithmetic operations, and with $t$ right-hand sides, the cost is $O(tn^{2})$. Finally, each variance is $(k_{\ast\ast})_{\ell}-\|V_{\ell,:}\|_{2}^{2}$, which is computed by rowwise squared norms in $O(tn)$. Combining these gives the stated bounds. ∎
17 Case studies: semiring and combinatorial algorithms
This section sketches how the same template applies to semiring matrix products and closures.
Lemma 17.1 (Semiring case studies reduce to semiring matrix kernels).
Each semiring/combinatorial case study in Section 17 reduces correctness to exactness of a semiring matrix product/closure routine and reduces complexity to the cost of the underlying semiring matrix product plus polylog factors from recursion/closure depth.
Proof.
Valiant CFG recognition, APSP via min-plus closure, and triangle detection via Boolean MM are all standard reductions to semiring products/closures. GPR contributes a graded projection that keeps representations bounded per node. This is a representational change that preserves semiring operations. ∎
17.1 Valiant CFG recognition (semiring mode)
Theorem 17.2 (CFG recognition via semiring GPR kernel (sketch)).
This theorem is a template statement. It assumes a semiring matrix kernel with bounded per-entry representation cost under graded projection.
Let $G$ be a CFG in Chomsky normal form with a finite set of nonterminals $\mathcal{N}$, and let $w=w_{1}\cdots w_{n}$ be an input string. Let $\mathcal{S}_{G}$ be the semiring of nonterminal-sets used in Valiant’s construction:
  • elements are subsets of $\mathcal{N}$,
  • $\oplus$ is set union,
  • $\otimes$ is grammar span-composition: for $\subseteq\mathcal{N}, XYX\otimes Y$ is the set of all $A\in\mathcal{N}$ such that there exist $B\in X$, $C\in Y$ with a production $A\to BC$ in $G$.
[…]
Graded projection invariant: At each recursion node, representations of semiring elements (nonterminal sets) are normalized to a bounded canonical form (e.g., bitset packing or another fixed representation), and this normalization preserves $\oplus$ and $\otimes$ exactly.
From this we can establish:
  1. 1.
    (Correctness) The algorithm decides CFG recognition correctly: $w\in L(G)$ iff the start symbol $S\in(M^{\star})_{1n}$.
  2. 2.
    (Complexity reduction to semiring MM) The running time is
    $T_{\mathrm{CFG}}(n)~=~\widetilde{O}\!\left(T_{\mathcal{S}_{G}}(n)\right),$ (166)
[…]
Proof.
Correctness
Valiant’s reduction constructs a semiring closure whose fixed point $M^{\star}$ satisfies the invariant: $(M^{\star})_{ij}$ equals the set of nonterminals that derive the substring $w_{i}\cdots w_{j}$. This is proved by induction on span length in the standard Valiant framework, using that $\oplus$ and $\otimes$ implement union and binary production composition exactly. Therefore $w\in L(G)$ iff the start symbol $S$ appears in $(M^{\star})_{1n}$. The only additional requirement in our setting is that graded projection does not alter the semiring computation. By hypothesis, the node-local normalization preserves $\oplus$ and $\otimes$ exactly (it is only a change of representation), so the computed fixed point equals the standard semiring fixed point $M^{\star}$.
Complexity
The blocked closure algorithm performs a polylogarithmic number of semiring matrix multiplications (and $O(n^{2})$ additional bookkeeping) across recursion levels, so its runtime is $\widetilde{O}(T_{\mathcal{S}_{G}}(n))$. If the semiring multiplication kernel runs in $\widetilde{O}(n^{2})$ time under bounded representations, then the overall closure inherits that near-quadratic bound up to polylog factors. ∎
17.2 APSP via min-plus products (semiring mode)
Theorem 17.3 (APSP via graded min-plus products (sketch)).
This theorem is a template statement. It assumes a semiring matrix kernel with bounded per-entry representation cost under graded projection.
Let $G$ be a weighted directed graph on $n$ vertices with edge weights in $\mathbb{R}\cup\{\infty\}$, and let $W\in(\mathbb{R}\cup\{\infty\})^{n\times n}$ be its weight matrix. Consider the min-plus semiring:
$(\oplus,\otimes)~=~(\min,+),$ (167)
so that min-plus matrix multiplication is:
$(X\otimes Y)_{ij}~=~\min_{k}\left(X_{ik}+Y_{kj}\right).$ (168)
[…]
(Offset-bucket invariant) At each recursion node, each block $X$ is represented as:
$X=\Delta_{X}\mathbf{1}\mathbf{1}^{\top}+R_{X},$ (169)
where $\Delta_{X}$ is a block-dependent scalar offset and $R_{X}$ is a residual matrix whose finite entries lie in a bounded window (a fixed bucket range). The kernel computes min-plus products exactly on the represented values (offsets are tracked exactly. Residuals are computed exactly in the chosen bounded representation).
From this we can show:
  1. 1.
    (Correctness) The algorithm returns the exact APSP distance matrix (all-pairs shortest path distances), provided the graph has no negative cycles (the standard APSP precondition).
  2. 2.
    (Complexity reduction to min-plus MM) The runtime is:
    $T_{\mathrm{APSP}}(n)~=~\widetilde{O}\!\left(T_{\mathbin{\min\!+\!}}(n)\right),$ (170)
[…]
Proof.
Correctness
The classical min-plus algebraic reductions show that APSP distances are obtained by min-plus closure operations on $W$ (e.g., repeated squaring computes shortest paths of length up to $2^{k}$ after $k$ products, and closure stabilizes to all-pairs distances in the absence of negative cycles). Our algorithm performs the same min-plus operations, but with an alternative representation, i.e., each block is stored as an exact offset plus a bounded residual. Because offsets are tracked exactly and residual arithmetic is exact within the chosen representation, the represented value equals the true value at every step. Therefore the output equals the standard min-plus closure result.
Complexity
The closure method performs a polylogarithmic number of min-plus matrix multiplications (and $O(n^{2})$ bookkeeping), so the runtime is $\widetilde{O}(T_{\mathbin{\min\!+\!}}(n))$. If each min-plus product can be implemented in $\widetilde{O}(n^{2})$ time under bounded residual representations, then the overall APSP runtime is $\widetilde{O}(n^{2})$ up to polylog factors. ∎
17.3 Triangle detection via Boolean MM (semiring mode)
Theorem 17.4 (Triangle detection via Boolean GPR kernel (sketch)).
This theorem is a template statement. It assumes a semiring matrix kernel with bounded per-entry representation cost under graded projection.
Let $G$ be a directed or undirected graph on $n$ vertices with adjacency matrix $A\in\{0,1\}^{n\times n}$. Work over the Boolean semiring (,)=(,)($\oplus$$\otimes$)=($\lor$$\land$), so Boolean matrix multiplication is:
$(X\otimes Y)_{ij}~=~\bigvee_{k=1}^{n}\left(X_{ik}\wedge Y_{kj}\right).$ (171)
Assume Boolean matrix multiplication is implemented by a recursive blocked GPR kernel KernelMatMulBool that uses graded projection corresponding to word-packing of bits (i.e., each row/column block is stored as bitsets packed into machine words and block products are computed by word-level operations), then:
  1. 1.
    (Correctness) $G$ contains a triangle if and only if there exist $i,j$ such that:
    $(A\otimes A)_{ij}\wedge A_{ji}~=~1.$ (172)
    Equivalently, in Boolean arithmetic, $\mathrm{trace}(A\otimes A\otimes A)=1$ if and only if a triangle exists.
  2. 2.
    (Complexity reduction to Boolean MM) Triangle detection reduces to one Boolean matrix multiplication plus $O(n^{2})$ additional work for the final check, so the runtime is $\widetilde{O}(T_{\mathrm{BoolMM}}(n))$, where $T_{\mathrm{BoolMM}}(n)$ is the cost of one Boolean matrix multiplication implemented by KernelMatMulBool. In particular, with word-packing and a cache-friendly blocked implementation one obtains $O(n^{3}/w)$ word operations (where $w$ is the machine word size).
Proof.
Correctness
A triangle exists if and only if there are vertices $i,j,k$ with edges $(i,k)$, $(k,j)$, and $(j,i)$. The Boolean product $(A\otimes A)_{ij}=1$ holds exactly when there exists $k$ such that $A_{ik}=1$ and $A_{kj}=1$, i.e., a length-2 path from $i$ to $j$. Thus $(A\otimes A)_{ij}\wedge A_{ji}=1$ holds exactly when there is a length-2 path $i\to j$ and an edge $j\to i$, which is precisely a directed triangle. The trace characterization is equivalent.
Complexity
Computing $B:=A\otimes A$ is one Boolean matrix multiplication, and checking whether any entry satisfies $B_{ij}\wedge A_{ji}=1$ is $O(n^{2})$, so the total runtime is $\widetilde{O}(T_{\mathrm{BoolMM}}(n))$. With word-packing, each inner product reduces to a wordwise AND and an “any-nonzero” test (or popcount), giving the standard $O(n^{3}/w)$ bound. ∎
17.4 How to create new GPR reductions
  1. 1.
    Identify the dominant kernel(s): GEMM, QR/LDL, or semiring product/closure.
  2. 2.
    Choose the execution model: integral dyadic semantics (exact) or float baseline comparator (commensurate).
  3. 3.
    Specify the graded projection invariant and prove it is maintained recursively.
  4. 4.
    State the decision sites and write a ledger inequality $\varepsilon_{\mathrm{tot}}\leq\delta$ (float mode).
  5. 5.
    Conclude complexity by a recurrence. Conclude correctness/commensurate accuracy by composition.
Part IV Generality of GPR: admissible packings, design patterns, and extensions (Part IV)
18 What is GPR really doing?
At the highest level, GPR consists of two separable ingredients:
  1. 1.
    Graded embedding and extraction: Embed the desired computation into a graded representation (e.g., a $\beta$-adic band decomposition), then extract a target band/degree using a stable decoder (gap/slack).
  2. 2.
    Recursive invariant amplification: Re-establish projection/normalization invariants at every recursion node so that the numerical preconditions for extraction (gap inequalities) hold uniformly across depth.
The “secret sauce” is the pair of invariants that are enforced with constant overhead at each recursive step. Part I provides a concrete instance for matrix multiplication. Part IV abstracts the conditions to express the GPR framework in its most general form.
18.1 Abstract bilinear kernel model
Definition 18.1 (Bilinear kernel).
Let $M,N$ be abelian groups (or modules/rings when needed). A bilinear kernel is a map:
$\mathcal{B}~:~M\times M~\to~N$ (173)
that is linear in each argument. Examples: matrix product (with $M=N=\mathcal{L}_{D}^{n\times n}$), rank-$k$ updates, semiring products, and structured bilinear forms.
Definition 18.2 (Classical block bound).
Fix a blocked recursion and let $\ell$ denote depth. Let $S_{\ell}$ be a bound on the magnitude/representation size of a single classical block application of $\mathcal{B}$ at depth $\ell$ (under the relevant execution model: integral magnitude, float magnitude, and semiring representation size).
Definition 18.3 (Size functional / representation size).
We fix a nonnegative size functional $\|\cdot\|_{N}:N\to\mathbb{R}_{\geq 0}$ appropriate to the codomain $N$ of the bilinear kernel. In the numeric setting, $\|\cdot\|_{N}$ may be absolute value or a norm. In semiring/settings with structured objects, it may be a representation size (e.g., bitset size, bucket range, set cardinality, encoding length).
We assume $\|\cdot\|_{N}$ is subadditive:
$\|u+v\|_{N}~\leq~\|u\|_{N}+\|v\|_{N}~\text{for all}~u,v\in N,$ (174)
and for scalar coefficients $\lambda\in\mathbb{Z}$ (or the relevant coefficient ring), $\|\lambda u\|_{N}\leq|\lambda|\|u\|_{N}$ whenever scalar multiplication is defined.
19 GPR-admissible packing: the master condition
Definition 19.1 (GPR-admissible packing (master condition)).
Fix a bilinear kernel $\mathcal{B}:M\times M\to N$ and a target band/degree (or set of bands) we wish to extract. A recursive packing scheme is GPR-admissible if there exist depth-independent constants $C_{T},C_{L}>0$ such that, for every depth $\ell$, each packed leaf computation admits a three-band form:
$Z_{\ell}~=~\beta U_{\ell}+T_{\ell}+\frac{1}{\beta}L_{\ell},$ (175)
where $T_{\ell}$ is the target band to be recovered, and:
$|T_{\ell}|\leq C_{T}S_{\ell},~|L_{\ell}|\leq C_{L}S_{\ell},$ (176)
(or the corresponding bounded-size statement in a non-numeric setting).
Remark 19.2.
When we write $|x|$ (here in Part IV), it should be interpreted as $\|x\|_{N}$ from Definition 18.3.
Theorem 19.3 (Global gap from the master condition).
Assume the master condition (Definition 19.1) holds: for every depth $\ell$ each packed leaf quantity admits the three-band form:
$Z_{\ell}~=~\beta U_{\ell}+T_{\ell}+\frac{1}{\beta}L_{\ell},$ (177)
and there exist depth-independent constants $C_{T},C_{L}>0$ such that:
$\|T_{\ell}\|_{N}\leq C_{T}S_{\ell},~\|L_{\ell}\|_{N}\leq C_{L}S_{\ell},$ (178)
where $\|\cdot\|_{N}$ is the fixed size functional on $N$ (Definition 18.3) and $S_{\ell}$ is the depth-$\ell$ classical block bound (Definition 18.2). Let:
$C$ $:=$ $\max\{C_{T},C_{L}\},$ (179)
$S_{\max}$ $:=$ $\max_{\ell}S_{\ell},$ (180)
and fix $\delta\in(0,1/2]$.
Assume the extractor is the two-round sieve of Part I (Definition 5.1) and that the master gap lemma (Lemma 5.3) applies with the same coefficient lattice/units used to define $S_{\ell}$. [^1] For example, if extraction is performed on a coefficient lattice $D^{-2}\mathbb{Z}$, incorporate the $D^{2}$ scaling into the definition of $S_{\ell}$ (or equivalently replace $S_{\max}$ by $D^{2}S_{\max}$ below). If the global base $\beta$ satisfies:
$CS_{\max}\left(\frac{2}{\beta}+\frac{1}{\beta^{2}}\right)~\leq~\frac{1}{2}-\delta,$ (181)
then the master gap inequality of Lemma 5.3 holds for every depth $\ell$ and every packed leaf entry, and therefore the sieve/extractor returns the exact target band $T_{\ell}$ at every depth. More generally, any bit-length/complexity bounds proved for the concrete kernel in Parts I–III extend with the same asymptotic form, up to constant factors depending only on $C$ (and the fixed parameters of the underlying model).
Remark 19.4 (Consistency requirements in Part IV).
All Part IV bounds are stated using the size functional $\|\cdot\|_{N}$ and the depth-wise maximum $S_{\max}$. When extracting on a coefficient lattice, absorb any fixed scaling (such as $D^{2}$) into the definition of $S_{\ell}$ or into $S_{\max}$.
Proof.
Fix any depth $\ell$ and any scalar entry of a packed leaf computation. All bounds use the size functional $\|\cdot\|_{N}$ from Definition 18.3, and coefficients are understood via the graded convolution identity (Lemma 20.1) when derived from embeddings. By the master condition and the definition of $C$:
$\|T_{\ell}\|_{N}~\leq~C_{T}S_{\ell}~\leq~CS_{\max},$ (182)
$\|L_{\ell}\|_{N}~\leq~C_{L}S_{\ell}~\leq~CS_{\max},$ (183)
so the left-hand side of the master gap inequality from Part I satisfies:
$\frac{\|L_{\ell}\|_{N}}{\beta}+\frac{\|T_{\ell}\|_{N}}{\beta}+\frac{\|L_{\ell}\|_{N}}{\beta^{2}}$ (184)
$\leq$ $CS_{\max}\left(\frac{1}{\beta}+\frac{1}{\beta}+\frac{1}{\beta^{2}}\right)$ (185)
$=$ $CS_{\max}\left(\frac{2}{\beta}+\frac{1}{\beta^{2}}\right).$ (186)
If Equation 181 holds, then this is at most $1/2-\delta$, so Lemma 5.3 applies entrywise. Therefore the two sieve gap inequalities hold and Lemma 5.2 gives $\mathcal{S}_{\beta}(Z_{\ell})=T_{\ell}$ exactly for that entry. Because the argument is uniform in $\ell$ and in the entry, extraction is exact at every depth.
Finally, the bit-length/complexity analyses in Parts I–III use only (1) a depth-uniform base $\beta$ satisfying the gap inequalities, and (2) uniform bounds on the magnitudes/sizes of the extracted bands. Replacing the concrete constants by $C$ changes only constant factors and preserves the asymptotic forms. ∎
A key consequence for model-honest complexity is that admissible packings have depth-uniform temporary size once $\beta$ is fixed globally (see Lemma 20.6).
Corollary 19.5 (Simple explicit global base choice).
Under the hypotheses of Theorem 19.3, it suffices to choose any integer base $\beta\geq 2$ such that:
$\beta~\geq~\frac{6CS_{\max}}{1-2\delta}+1,$ (187)
(with the usual “appropriate units” convention: for coefficient-lattice extraction replace $S_{\max}$ by $D^{2}S_{\max}$).
Proof.
Let $X:=CS_{\max}$. Because $\beta\geq 1$, we have $1/\beta^{2}\leq 1/\beta$ and therefore:
$\frac{2}{\beta}+\frac{1}{\beta^{2}}~\leq~\frac{3}{\beta}.$ (188)
Thus it suffices that:
$X\cdot\frac{3}{\beta}~\leq~\frac{1}{2}-\delta~=~\frac{1-2\delta}{2},$ (189)
i.e.:
$\beta~\geq~\frac{6X}{1-2\delta}.$ (190)
Adding +1+1 and enforcing $\beta\geq 2$ preserves the inequality and gives the stated sufficient choice. ∎
Remark 19.6 (On constants in the explicit $\beta$ bound).
Corollary 19.5 is stated in a “simple closed form” for readability. If the inequality $2/\beta+1/\beta^{2}\leq 3/\beta$ is used, then a factor $6$ (rather than $4$) in front of $CS_{\max}/(1-2\delta)$ is the clean no-slack choice. The tightest (exact) condition is:
$CS_{\max}\left(\frac{2}{\beta}+\frac{1}{\beta^{2}}\right)~\leq~\frac{1}{2}-\delta$ (191)
by solving the resulting quadratic inequality in $\beta$, which gives a slightly smaller explicit base. In this manuscript, constant factors are secondary. What matters is that $\beta=\Theta(CS_{\max})$ for any fixed $\delta=\Omega(1)$.
20 BWBM: an easily checkable sufficient condition
The master condition is conceptually clean but may be inconvenient to verify directly. BWBM is a sufficient condition that is likely to be easier to check from the embedding structure.
Lemma 20.1 (Convolution identity for graded bilinear embeddings).
Let $\mathcal{B}:M\times M\to N$ be bilinear and let
$R(\beta)$ $=$ $\sum_{p\in D}\beta^{p}R^{[p]},$ (192)
$S(\beta)$ $=$ $\sum_{q\in D^{\prime}}\beta^{q}S^{[q]},$ (193)
where $D,D^{\prime}$ are finite index sets and $\beta$ is a formal symbol commuting with coefficients. Thus, for every integer degree $e$, the coefficient of $\beta^{e}$ in $\mathcal{B}(R(\beta),S(\beta))$ equals:
$\sum_{(p,q):p+q=e}\mathcal{B}(R^{[p]},S^{[q]}).$ (194)
Proof.
By bilinearity:
$\mathcal{B}(R(\beta),S(\beta))$ $=$ $\mathcal{B}\!\left(\sum_{p}\beta^{p}R^{[p]},~\sum_{q}\beta^{q}S^{[q]}\right)$ (195)
$=$ $\sum_{p,q}\beta^{p+q}\mathcal{B}(R^{[p]},S^{[q]}).$ (196)
Collecting terms of equal exponent $e$ gives the stated coefficient formula. ∎
Definition 20.2 (Bounded-Width, Bounded-Mixing (BWBM)).
Fix a bilinear kernel $\mathcal{B}$ and a target band/degree. Consider graded embeddings of the form:
$R(\beta)$ $=$ $\sum_{p\in D}\beta^{p}R^{[p]},$ (197)
$S(\beta)$ $=$ $\sum_{q\in D^{\prime}}\beta^{q}S^{[q]},$ (198)
where $D,D^{\prime}$ are finite exponent windows.
A recursive packing satisfies BWBM if there exist depth-independent constants $\Delta,\Lambda,k_{T},k_{L}$ such that:
  1. 1.
    Exponent window (bounded width): All nonzero exponents lie in an interval of width at most $\Delta$:
    $\max D-\min D$ $\leq$ $\Delta,$ (199)
    $\max D^{\prime}-\min D^{\prime}$ $\leq$ $\Delta.$ (200)
  2. 2.
    Bounded mixing: Each coefficient block is a bounded $\ell_{1}$ combination of atomic subblocks:
    $R^{[p]}$ $=$ $\sum_{u}\alpha_{p,u}X_{u},~\sum_{u}|\alpha_{p,u}|~\leq~\Lambda,$ (201)
    $S^{[q]}$ $=$ $\sum_{v}\beta_{q,v}Y_{v},~\sum_{v}|\beta_{q,v}|~\leq~\Lambda,$ (202)
    where $X_{u},Y_{v}$ are the “atomic” blocks whose classical product is bounded by $S_{\ell}$.
  3. 3.
    Bounded convolution width: The number of contributing pairs for the target and lower bands is bounded:
    $|\{(p,q):p+q=e_{0}\}|$ $\leq$ $k_{T},$ (203)
    $|\{(p,q):p+q=e_{0}-1\}|$ $\leq$ $k_{L},$ (204)
    where $e_{0}$ indexes the target band (adapt indices to the chosen band convention).
Lemma 20.3 (BWBM implies the master condition).
Assume a recursive packing satisfies BWBM (Definition 20.2) at depth $\ell$ with parameters $\Lambda,k_{T},k_{L}$ and classical block bound $S_{\ell}$.
Let $T_{\ell}$ denote the target band (degree $e_{0}$) coefficient and $L_{\ell}$ denote the lower-band (degree $e_{0}-1$) coefficient in the packed product expansion at depth $\ell$ (i.e., the coefficients that appear as $Z_{\ell}=\beta U_{\ell}+T_{\ell}+L_{\ell}/\beta$ after collecting degrees), then the master condition (Definition 19.1) holds with:
$|T_{\ell}|$ $\leq$ $C_{T}S_{\ell},$ (205)
$|L_{\ell}|$ $\leq$ $C_{L}S_{\ell},$ (206)
where:
$C_{T}$ $=$ $k_{T}\Lambda^{2},$ (207)
$C_{L}$ $=$ $k_{L}\Lambda^{2}.$ (208)
In particular, taking $C:=\max\{C_{T},C_{L}\}$, BWBM implies GPR-admissibility with constants $C_{T},C_{L}$.
Proof.
We prove the bound for $T_{\ell}$, and the proof for $L_{\ell}$ is identical.
Step 1: Expand the target band as a bounded sum of bilinear terms
By the definition of the graded embedding and bilinearity, the target-band coefficient (degree $e_{0}$) of the packed product is:
$T_{\ell}~=~\sum_{(p,q):p+q=e_{0}}\mathcal{B}\left(R^{[p]},S^{[q]}\right),$ (209)
where the sum ranges over all exponent pairs contributing to the target degree. By BWBM (bounded convolution width), there are at most $k_{T}$ such pairs.
Step 2: Bound one contributing term using bounded mixing.
Fix one such pair $(p,q)$. By BWBM (bounded mixing):
$R^{[p]}~=~\sum_{u}\alpha_{p,u}X_{u},~~\sum_{u}|\alpha_{p,u}|~\leq~\Lambda,$ (210)
$S^{[q]}~=~\sum_{v}\beta_{q,v}Y_{v},~~\sum_{v}|\beta_{q,v}|~\leq~\Lambda,$ (211)
where $X_{u},Y_{v}$ are atomic blocks at depth $\ell$. By bilinearity:
$\mathcal{B}(R^{[p]},S^{[q]})~=~\sum_{u,v}\alpha_{p,u}\beta_{q,v}\mathcal{B}(X_{u},Y_{v}).$ (212)
By definition of the classical block bound $S_{\ell}$, every atomic bilinear evaluation satisfies:
$|\mathcal{B}(X_{u},Y_{v})|~\leq~S_{\ell}.$ (213)
Therefore, by the triangle inequality:
$\big|\mathcal{B}(R^{[p]},S^{[q]})\big|$ $\leq$ $\sum_{u,v}|\alpha_{p,u}||\beta_{q,v}||\mathcal{B}(X_{u},Y_{v})|$ (214)
$\leq$ $S_{\ell}\left(\sum_{u}|\alpha_{p,u}|\right)\left(\sum_{v}|\beta_{q,v}|\right)$ (215)
$\leq$ $\Lambda^{2}S_{\ell}.$ (216)
Step 3: Sum over at most $k_{T}$ contributing pairs.
Using the bound above and the fact that there are at most $k_{T}$ contributing pairs:
$|T_{\ell}|$ $\leq$ $\sum_{(p,q):p+q=e_{0}}\big|\mathcal{B}(R^{[p]},S^{[q]})\big|$ (217)
$\leq$ $k_{T}\cdot(\Lambda^{2}S_{\ell})$ (218)
$=$ $(k_{T}\Lambda^{2})S_{\ell},$ (219)
so we may take $C_{T}=k_{T}\Lambda^{2}$.
Lower band.
The same argument applied to the $(p,q)$ pairs with $p+q=e_{0}-1$ (at most $k_{L}$ of them) gives $|L_{\ell}|\leq(k_{L}\Lambda^{2})S_{\ell}$, so $C_{L}=k_{L}\Lambda^{2}$. ∎
Corollary 20.4 (Global gap under BWBM).
Assume BWBM (Definition 20.2) holds with parameters $\Lambda,k_{T},k_{L}$ at all depths, and let:
$C_{T}$ $:=$ $k_{T}\Lambda^{2},$ (220)
$C_{L}$ $:=$ $k_{L}\Lambda^{2},$ (221)
$C$ $:=$ $\max\{C_{T},C_{L}\}.$ (222)
Let $S_{\max}:=\max_{\ell}S_{\ell}$ be the maximum classical block bound (Definition 18.2) and fix $\delta\in(0,1/2]$. If the global base $\beta$ is chosen so that:
$CS_{\max}\left(\frac{2}{\beta}+\frac{1}{\beta^{2}}\right)~\leq~\frac{1}{2}-\delta,$ (223)
(in the appropriate units of the extractor, e.g., replace $S_{\max}$ by $D^{2}S_{\max}$ for coefficient-lattice extraction), then exact extraction holds at every depth. In particular, by Corollary 19.5 applied with $C=\max\{k_{T}\Lambda^{2},k_{L}\Lambda^{2}\}$, it suffices to take:
$\beta~\geq~\frac{6CS_{\max}}{1-2\delta}+1,$ (224)
(with the usual coefficient-lattice scaling convention if applicable).
Proof.
By Lemma 20.3, BWBM implies the master condition (Definition 19.1) with constants $C_{T}=k_{T}\Lambda^{2}$ and $C_{L}=k_{L}\Lambda^{2}$, so with $C=\max\{C_{T},C_{L}\}$ we have the uniform bounds:
$\|T_{\ell}\|_{N}$ $\leq$ $CS_{\ell},$ (225)
$\|L_{\ell}\|_{N}$ $\leq$ $CS_{\ell}$ (226)
for all depths $\ell$. Because $S_{\ell}\leq S_{\max}$, the hypotheses of Theorem 19.3 are satisfied with $S_{\max}$. Therefore Theorem 19.3 implies extraction is exact at every depth under Equation 223. The explicit inequality:
$\beta~\geq~\frac{6CS_{\max}}{1-2\delta}+1$ (227)
is exactly Corollary 19.5 specialized to this $C$. ∎
Remark 20.5.
For any fixed constant slack $\delta=\Omega(1)$ and monotone depth bounds $S_{\ell}\leq S_{0}$, the condition reduces to $\beta=\Theta(CS_{0})$.
Lemma 20.6 (Uniform temporary size under BWBM).
Assume BWBM holds with depth-independent parameters $(\Delta,\Lambda,k_{T},k_{L})$ and that the extractor parameters (e.g., base $\beta$ or moduli) are chosen globally so extraction is exact at every depth (Corollary 20.4). At this point every materialized packed temporary has size (bit-length / representation size) bounded uniformly in depth:
$\text{size}(Z_{\ell})~=~O\!\left(\log\beta+\log S_{\max}\right),$ (228)
where $S_{\max}=\max_{\ell}S_{\ell}$ and $\text{size}(\cdot)$ is the chosen size functional (Definition 18.3) specialized to the domain.
Proof.
BWBM bounds the magnitude/size of each target and tail band by $O(S_{\ell})\leq O(S_{\max})$ and restricts the exponent window width by $\Delta=O(1)$. Thus a packed value consists of $O(1)$ bands whose magnitudes are $O(S_{\max})$ multiplied by $\beta^{O(1)}$. Therefore the overall size is bounded by $O(\log S_{\max}+\log\beta)$ uniformly in depth. ∎
21 Design patterns enabled by BWBM
Lemma 21.1 (Executable bounded mixing via quantized coefficients).
Suppose a BWBM packing uses mixing coefficients $\alpha_{p,u},\beta_{q,v}$ that are integers of bounded magnitude (e.g., $\alpha_{p,u}\in\{-c,\dots,c\}$ for a constant $c$) and each coefficient block uses only $O(1)$ nonzero coefficients, then the bounded mixing parameter $\Lambda$ in Definition 20.2 is $O(1)$ and the mixing is implementable using only $O(1)$ scaled additions per entry (no unbounded coefficient growth).
Proof.
If each coefficient block is a linear combination of at most $s=O(1)$ atomic blocks with integer coefficients of magnitude at most $c=O(1)$, then:
$\sum_{u}|\alpha_{p,u}|$ $\leq$ $sc=O(1),$ (229)
$\sum_{v}|\beta_{q,v}|$ $\leq$ $sc=O(1),$ (230)
so $\Lambda=O(1)$. Implementability follows because each mixed block entry is computed as a constant number of integer-scaled adds. ∎
BWBM describes a space of admissible packings beyond the canonical $(\beta,\beta^{-1})$ form:
  • Strassen-style mixing: allow small integer linear combinations of blocks ($\Lambda>1$) while keeping convolution width bounded.
  • Wide windows / multi-band extraction: target a constant-size band of degrees and recover via shifting or multi-degree extractors.
  • Alternative coefficient rings: use modular arithmetic or finite fields (e.g., CRT) for coefficient extraction rather than real rounding.
Bounded mixing is more than just conceptual: if the mixing coefficients are small integers and each mixed block uses only $O(1)$ terms, then BWBM mixing is implementable with $O(1)$ scaled additions per entry (see Lemma 21.1).
21.1 Beyond bilinear: multilinear and multigraded extensions
Remark 21.2 (Multilinear kernels).
For a $k$-linear map $\mathcal{B}:M_{1}\times\cdots\times M_{k}\to N$, one can embed each input into a graded module and target a multi-degree band. The master condition extends verbatim: as long as the target-band magnitude is bounded by a constant multiple of the classical bound at each depth, a single global base (or mixed bases) enforces exact extraction.
Remark 21.3 (Multigraded embeddings).
One may embed into $M[\beta_{1}^{\pm 1},\dots,\beta_{t}^{\pm 1}]$ or into a mixed-radix single base. The exponents become vectors. What matters numerically is a uniform bound on target-band magnitudes relative to depth-wise classical bounds.
21.2 Further generality: multilinear kernels and extractor abstraction
This section makes two generalizations explicit:
  1. 1.
    Extending GPR from bilinear kernels to multilinear and multigraded kernels, and
  2. 2.
    Abstracting “extraction” beyond real rounding gaps.
21.3 A) Beyond bilinear: multilinear and multigraded GPR
Definition 21.4 ($k$-linear kernel).
Let $M_{1},\dots,M_{k},N$ be abelian groups (or modules) and let
$\mathcal{B}:M_{1}\times\cdots\times M_{k}\to N$ (231)
be a map that is linear in each argument separately. We call $\mathcal{B}$ a $k$-linear kernel.
Definition 21.5 (Multigraded embeddings and multi-degree windows).
Fix formal bases $\beta_{1},\dots,\beta_{t}$ (or a mixed-radix encoding of them). A multigraded embedding of an input $x\in M_{i}$ is an expression
$R_{i}(\boldsymbol{\beta})$ $=$ $\sum_{p\in D_{i}}\boldsymbol{\beta}^{p}R_{i}^{[p]},$ (232)
$\boldsymbol{\beta}^{p}$ $:=$ $\beta_{1}^{p_{1}}\cdots\beta_{t}^{p_{t}},$ (233)
where $D_{i}\subset\mathbb{Z}^{t}$ is a finite exponent window and $R_{i}^{[p]}\in M_{i}$. A target multi-degree is a vector $e_{0}\in\mathbb{Z}^{t}$, or more generally a finite band $E_{0}\subset\mathbb{Z}^{t}$.
Definition 21.6 (Multilinear master condition).
Fix a $k$-linear kernel $\mathcal{B}$ (Definition 21.4) and a target multi-degree band $E_{0}\subset\mathbb{Z}^{t}$. A recursive multigraded packing is GPR-admissible if there exists a depth-independent constant $C>0$ such that at every depth $\ell$, the total contribution to the target band (the coefficient block(s) indexed by $E_{0}$) has size at most $CS_{\ell}$:
$\big\|T_{\ell,E_{0}}\big\|_{N}~\leq~CS_{\ell},$ (234)
where $S_{\ell}$ is the classical depth-$\ell$ bound (Definition 18.2) and $\|\cdot\|_{N}$ is the size functional (Definition 18.3).
Remark 21.7.
This is the direct analogue of Definition 19.1 for the multilinear/multigraded setting: Instead of a single middle band $T_{\ell}$, we control the entire target multi-degree band $E_{0}$ by a constant multiple of the classical bound $S_{\ell}$.
21.4 B) Beyond rounding gaps: an extractor abstraction
Definition 21.8 (Extractor interface).
An extractor is a map:
$\mathrm{Extract}:~\mathsf{Packed}~\to~\mathsf{Target}$ (235)
together with a sufficient condition $\mathsf{Cond}$ on packed values, such that: whenever a packed value $Z$ satisfies $\mathsf{Cond}(Z)$, the extractor returns the intended target component:
$\mathsf{Cond}(Z)~\Longrightarrow~\mathrm{Extract}(Z)~=~\mathrm{TargetPart}(Z).$ (236)
Remark 21.9 (Examples of extractors).
  • Rounding-gap extractor (Part I): Here $\mathsf{Packed}=\mathbb{R}$ (or a coefficient lattice), $\mathrm{Extract}=\mathcal{S}_{\beta}$, and $\mathsf{Cond}$ is the master gap inequality ensuring band separation.
  • CRT extractor: Here $\mathsf{Packed}$ is a tuple of residues $(Z\bmod p_{1},\dots,Z\bmod p_{r})$ and $\mathrm{Extract}$ is CRT reconstruction. The condition is that the true integer lies in a known reconstruction interval (no wraparound).
  • Exact digit extractor: Here $\beta$ is a power of two and $\mathrm{Extract}$ is bit-slicing/digit extraction. The condition is that carries do not overlap the target digit window.
  • Semiring/bucketing extractor. Here $\mathrm{Extract}$ may be a projection to buckets/grades (e.g., distance offsets + bounded residuals), i.e., the condition is representational disjointness or bounded width of the active bucket band.
Theorem 21.10 (Extractor correctness under a master condition (template)).
Fix a kernel (bilinear or $k$-linear) and a target band (degree or multi-degree band). Assume a master condition of the form:
$\|T_{\ell}\|_{N}~\leq~CS_{\ell}~\text{(or}~\|T_{\ell,E_{0}}~\|_{N}~\leq~CS_{\ell}\text{)}$ (237)
holds uniformly in depth $\ell$, for a depth-independent constant $C$.
Assume further that the chosen extractor $\mathrm{Extract}$ admits a sufficient condition $\mathsf{Cond}$ that is implied by the above bound once the global parameters of the packing (e.g., base(s) $\beta$ or moduli set) are chosen large enough as a function of $C$ and $S_{\max}:=\max_{\ell}S_{\ell}$. The extraction is then correct at every depth and the corresponding GPR reduction inherits the same asymptotic bit/format complexity, up to constants depending on $C$ and on the extractor parameters.
Proof sketch.
Uniform master bounds imply a uniform sufficient condition for extraction once global parameters are chosen (base separation, CRT interval, bucket width, etc.). The extraction is then correct at every leaf, and correctness propagates up the recursion by the same structural induction as in Parts I–III. Complexity is unchanged asymptotically because extractor operations are $O(1)$ per entry per level (or bounded by a fixed number of moduli/buckets). ∎
21.5 Alternative numeric domains
Remark 21.11 (Finite fields and CRT).
The coefficient extraction mechanism used in Part I (rounding gaps) is one instantiation. For exact arithmetic over $\mathbb{Z}$ one can also use CRT-style embeddings: compute in several moduli and reconstruct coefficients by CRT. This gives an alternative family of GPR-style extractors.
Example: Weighted Valiant-style dynamic programs
Weighted Valiant-style algorithms (e.g., weighted CFG variants) provide a testbed where the heavy work is a small fixed family of bilinear kernels over structured semirings. Such kernels often satisfy BWBM: bounded exponent windows (span grades), bounded mixing (bounded rule arity/weights), and bounded convolution width (bounded number of contributing band pairs). Thus they are GPR-admissible by Lemma 20.3.
22 Discussion and conclusions
We have argued for a simple principle: algebraic packing is only as fast as its honest cost model. GPR turns that principle into a reusable methodology by pairing (1) a graded algebraic embedding and extractor with (2) a per-node graded projection discipline that keeps recursion stable under bounded intermediates.
Canonical kernel.
Part I gives a proof-complete instantiation for matrix multiplication over dyadic lattice semantics. The key technical points are the three-band packing identity, exact two-round extraction under a master gap condition, and recursive invariant amplification via centering and dyadic $\ell_{2}$ normalization. Together these give a depth-uniform base choice and the near-quadratic recurrence $4T(n/2)+O(n^{2})$, giving soft-quadratic bit/format complexity once any required staged precision overhead is made explicit.
Float-GPR and commensurate accuracy.
Part II shifts from exactness to comparator-faithful execution in a fixed format. The main message is that correctness hinges on decision-site invariance: if all discrete decisions (extraction and related thresholds) have slack $\delta$ and the drift ledger satisfies $\varepsilon_{\mathrm{tot}}\leq\delta$, then the GPR implementation is commensurate-accurate to the baseline comparator on the baseline-feasible regime. Guarded normalization provides a concrete mechanism for making the $\ell_{2}$ discipline safe without under-normalizing, while keeping overshoot bounded and depth-independent.
Reductions and portability.
Part III shows that $\omega_{\mathrm{GPR}}=2$ for a single kernel does not automatically transfer through classical unit-cost reductions. Instead, each host algorithm must be reorganized so that all of its dominant dense arithmetic is expressed through GPR-compatible kernels and all intermediate growth remains controlled. The case studies serve as templates for doing that systematically.
Generality and design space.
Part IV abstracts the mechanism in terms of admissible packings and extractors. The master condition isolates what must be bounded to make extraction uniform across depth. BWBM provides an easily checkable sufficient condition for admissibility, and the extractor abstraction clarifies how the same reasoning extends beyond real rounding gaps (e.g., CRT and semiring bucket projections). The multilinear and multigraded extensions indicate that the same mechanism applies well beyond bilinear kernels.
Limitations and future work.
The framework is strongest when (1) the host algorithm is GEMM-dominant in a blocked/recursive form, and (2) a projection discipline exists that keeps the relevant bands uniformly bounded. The main improvements ahead are sharpening constants, enlarging feasible regimes under more structure assumptions, and exploring alternative extractor mechanisms with better headroom tradeoffs.
Appendix A Parallel bit-complexity and communication overhead
This appendix records a simple but useful consequence of the depth-uniform representation bounds proved in the main text: once a GPR kernel has eliminated depth-dependent growth of packed intermediates, the same reduction appears immediately in distributed/parallel settings as a reduction in communication volume (measured in moved bits).
A.1 Bounded representation sizes and message volume
In distributed-memory models (e.g., BSP), a common abstraction is that the communication volume is the product of (i) how many scalar values are moved and (ii) the number of bits used to represent each moved scalar. If an execution transmits each scalar in a full $w$-bit machine word, then the per-scalar message size is $w$ bits regardless of how many of those bits are information-bearing.
By contrast, under GPR the packed temporaries have depth-uniform representation size once the extractor parameters are fixed globally so that extraction is valid at every depth; see Lemma 20.6. In the staged-precision accounting of Theorem 9.3, this means that communicating a scalar requires only the staged precision $b(n)$ bits (up to constant packing overhead), rather than a depth-dependent or uncontrolled bit-length.
Corollary A.1 (Parallel communication complexity).
Consider executing a GPR-based $2\times 2$ blocked recursion on $P$ processors in a distributed model. Suppose the schedule communicates $M(n,P)$ scalar values. [^2] For example, it is typical in standard 22D block layouts for dense products to have $M(n,P)=\widetilde{O}(n^{2}/\sqrt{P})$ communicated scalars; the statement below isolates the representation-size contribution. If each communicated scalar is transmitted in a packed representation whose bit-length is the staged precision $b(n)$, then the total communicated bit-volume satisfies:
$V_{\mathrm{GPR}}(n,P)~=~\widetilde{O}\!\left(M(n,P)b(n)\right).$ (238)
Relative to a fixed-word baseline that transmits each scalar as a full $w$-bit word:
$V_{\mathrm{GPR}}(n,P)~=~V_{\mathrm{word}}(n,P)\cdot\Theta\!\left(\frac{b(n)}{w}\right),$ (239)
i.e., the communication volume is reduced by a factor of $\Theta(w/b(n))$.
Proof.
Under the depth-uniform size bound of Lemma 20.6 (and the associated staged-precision interpretation in Theorem 9.3), every communicated intermediate can be represented using $b(n)$ bits of mantissa precision (up to constant overhead for packing into words). If the schedule moves $M(n,P)$ scalars, then transmitting each in $b(n)$ bits gives $V_{\mathrm{GPR}}(n,P)=\widetilde{O}(M(n,P)b(n))$ bits. A word baseline sends the same $M(n,P)$ scalars in $w$ bits each, so $V_{\mathrm{word}}(n,P)=\widetilde{O}(M(n,P)w)$. Taking the ratio gives the stated factor. ∎
Remark A.2 (When $b(n)$ exceeds $w$).
If $b(n)>w$, then each communicated scalar occupies $K(n)=\lceil b(n)/w\rceil$ words, so the same accounting gives $V_{\mathrm{GPR}}(n,P)=\widetilde{O}(M(n,P)K(n)w)$. In the common Word-RAM regime $w=\Theta(\log n)$ and the basic $2\times 2$ GPR recursion where $b(n)=\Theta(\log n)$, one has $K(n)=\Theta(1)$, so the per-scalar communication remains $\Theta(w)$ bits while still avoiding depth-dependent blowup.
Appendix B Exact dyadic-lattice extractor specification (integral mode)
This appendix gives a concrete, implementable specification of the dyadic-lattice rounding operator $\frac{1}{D^{2}}\left\lfloor D^{2}(\right\rceil\cdot)$ and the sieve extractor $\mathcal{S}_{\beta}(\cdot)$ used in Part I. The purpose is to eliminate any ambiguity about “exactness” in the integral semantics: all operations below are exact on the declared success set (master gap holds with slack), and any failure of the gap condition can be detected.
B.1 Scalar representation for $\mathcal{L}_{D}$
Fix $D\in\mathbb{N}$ and recall $\mathcal{L}_{D}=D^{-1}\mathbb{Z}[1/2]$. Represent $x\in\mathcal{L}_{D}$ in the form:
$x~=~\frac{w}{D}2^{e},$ (240)
where $w\in\mathbb{Z}$ is an integer mantissa and $e\in\mathbb{Z}$ is an exponent offset. A concrete implementation stores $(w,e)$ as fixed-width machine integers (e.g., int64_t for $w$ with __int128 for intermediate arithmetic, and int32_t for $e$). Multiplication and division by powers of two are implemented as exponent updates $e\leftarrow e\pm t$.
Remark B.1 (Canonicalization (optional)).
This can also be canonicalized by shifting powers of two out of $w$ into $e$ (making $w$ odd unless $w=0$), or by enforcing a fixed mantissa window $\lvert w\rvert\leq 2^{W}$. These choices affect only constant factors and do not change the exactness of the rounding and sieve primitives specified below.
B.2 Exact nearest-integer rounding of dyadic rationals
We use a deterministic nearest-integer rounding $\left\lfloor\cdot\right\rceil:\mathbb{R}\to\mathbb{Z}$ with a fixed tie-breaking rule. Because the paper’s extraction guarantees use strict slack, tie-breaking is irrelevant on the success set, but we specify it anyway for completeness (ties-to-even is acceptable).
Definition B.2 (Exact rounding of an integer divided by a power of two).
Let $N\in\mathbb{Z}$ and $s\in\mathbb{Z}_{\geq 0}$. Define $\mathrm{RoundPow2}(N,s):=\left\lfloor N/2^{s}\right\rceil$ computed with $N=\mathrm{sgn}\cdot A$, where $\mathrm{sgn}\in\{+1,-1\}$ and $A=\lvert N\rvert\in\mathbb{Z}_{\geq 0}$:
$q~:=~\left\lfloor\frac{A}{2^{s}}\right\rfloor,~r~:=~A-q\cdot 2^{s},$ (241)
so $0\leq r<2^{s}$. Let $h:=2^{s-1}$ for $s\geq 1$ (and $h:=0$ for $s=0$). Define:
$\tilde{q}~:=~\begin{cases}q,&r<h,\\ q+1,&r>h,\\ q~\text{or}~q+1~\text{(tie-breaking)},&r=h.\end{cases}$ (242)
Return $\mathrm{RoundPow2}(N,s):=\mathrm{sgn}\cdot\tilde{q}$.
Remark B.3.
In code, this is implemented by bit shifts and bit masks on the absolute value (using a wider fixed-width type such as __int128 as needed), with a deterministic rule for the $r=h$ tie case.
B.3 Exact coefficient rounding $\frac{1}{D^{2}}\left\lfloor D^{2}(\right\rceil\cdot)$
Recall $\frac{1}{D^{2}}\left\lfloor D^{2}(\right\rceil z)=D^{-2}\left\lfloor D^{2}z\right\rceil$.
Definition B.4 (Exact coefficient rounding on $D^{-2}\mathbb{Z}$).
Let $z\in\mathcal{L}_{D}$ have representation $z=(w/D)2^{e}$, then:
$D^{2}z~=~(Dw)2^{e}.$ (243)
Define $N:=Dw\in\mathbb{Z}$ (computed exactly in a fixed-width wider integer type for intermediate arithmetic).
  1. 1.
    If $e\geq 0$, then $D^{2}z$ is an integer and $\left\lfloor D^{2}z\right\rceil=N\cdot 2^{e}$.
  2. 2.
    If $e<0$, let s:=e>0s:=-e>0 and define $\left\lfloor D^{2}z\right\rceil:=\mathrm{RoundPow2}(N,s)$ via Definition B.2.
And then define:
$\frac{1}{D^{2}}\left\lfloor D^{2}(\right\rceil z)~:=~\frac{1}{D^{2}}\left\lfloor D^{2}z\right\rceil\in D^{-2}\mathbb{Z}\subset\mathcal{L}_{D}.$ (244)
Lemma B.5 (Correctness of the implementation).
If $\frac{1}{D^{2}}\left\lfloor D^{2}(\right\rceil\cdot)$ is implemented as in Definition B.4, then for every $z\in\mathcal{L}_{D}$ it equals the mathematical definition $\frac{1}{D^{2}}\left\lfloor D^{2}(\right\rceil z)=D^{-2}\left\lfloor D^{2}z\right\rceil$ exactly.
Proof.
This is immediate from the representation $D^{2}z=(Dw)2^{e}$ and exact rounding of dyadic rationals by Definition B.2. ∎
B.4 Exact sieve extractor $\mathcal{S}_{\beta}(\cdot)$
We now specialize to bases $\beta$ that are powers of two, which makes division by $\beta$ an exponent-offset update.
Definition B.6 (Exact sieve extractor for $\beta=2^{b}$).
Fix an integer $b\geq 1$ and let $\beta:=2^{b}$. For $z\in\mathcal{L}_{D}$, define:
$\mathcal{S}_{\beta}(z)~:=~\frac{1}{D^{2}}\left\lfloor D^{2}(\right\rceil z)-\beta\frac{1}{D^{2}}\left\lfloor D^{2}(\right\rceil z/\beta).$ (245)
Here $z/\beta$ is implemented by exponent update $e\leftarrow e-b$ (exact in $\mathcal{L}_{D}$), and $\frac{1}{D^{2}}\left\lfloor D^{2}(\right\rceil\cdot)$ is implemented by Definition B.4.
Lemma B.7 (Implementation matches Part I sieve).
If $\mathcal{S}_{\beta}$ is implemented as in Definition B.6, then it matches the Part I definition $\mathcal{S}_{\beta}(z)=\frac{1}{D^{2}}\left\lfloor D^{2}(\right\rceil z)-\beta\frac{1}{D^{2}}\left\lfloor D^{2}(\right\rceil z/\beta)$ exactly, and Lemmas 5.2 and 5.3 apply verbatim.
Proof.
All steps are exact in $\mathcal{L}_{D}$: dividing by $\beta=2^{b}$ is an exponent-offset update, and $\frac{1}{D^{2}}\left\lfloor D^{2}(\right\rceil\cdot)$ is exact by the preceding lemma. Therefore the implemented $\mathcal{S}_{\beta}$ equals the mathematical sieve definition. ∎
B.5 Gap checking and failure policy (optional but recommended)
For debugging and for reference implementations, it is useful to explicitly check the master gap inequality before applying the sieve. For a packed scalar $z=\beta U+T+L/\beta$, compute the three quantities appearing in Lemma 5.3 in the lifted units and assert:
$\lvert D^{2}L/\beta\rvert+\lvert D^{2}T/\beta\rvert+\lvert D^{2}L/\beta^{2}\rvert~\leq~\frac{1}{2}-\delta.$ (246)
If the check fails, the implementation should either:
  1. 1.
    Return a failure flag (instance leaves the declared success set), or
  2. 2.
    Increase $\beta$ and retry (if using an adaptive base policy outside the constant-$\beta$ regime).
On the success set, the check passes and extraction is exact.
B.6 Notes on efficiency
The primitives above are constant-time per scalar under fixed-width machine arithmetic:
  • exponent updates are constant time,
  • RoundPow2 uses shifts/masks on a fixed-width integer type,
  • $\frac{1}{D^{2}}\left\lfloor D^{2}\cdot\right\rceil$ uses one multiply by $D$ (on mantissa) plus RoundPow2 when $e<0$,
  • the sieve uses two coefficient roundings and a scaled subtraction.
This is sufficient to support the bit/word-complexity accounting (including staged-precision regimes) used in Parts I–II.

### v3

@@
 We present Graded Projection Recursion (GPR), a framework for converting certain blocked recursive algorithms into model-honest (i.e., reflects full bit complexity) near-quadratic procedures under bounded intermediate budgets.
-Part I gives a proof-complete integral specification of a three-band packing identity and a two-round middle-band extractor, and shows how per-node centering and sqrt-free dyadic $\ell_{2}$ normalization (recursive invariant amplification) gives a sufficient packing base $\beta(\sigma)$ that grows linearly with the scaling depth $\sigma$ (i.e., logarithmic bit-growth) in exact arithmetic.
+Part I gives a proof-complete integral specification of a three-band packing identity and a two-round middle-band extractor, and shows how per-node centering and sqrt-free dyadic $\ell_{2}$ normalization (recursive invariant amplification) gives a sufficient packing base $\beta(\sigma)$ whose required bit-width grows linearly with the node-local lattice exponent $\sigma$ (so only logarithmic bit-growth in $n$ in the worst case) in exact arithmetic.
 On fixed-width hardware, exact evaluation of the packed recursion generally requires either an extended-precision path or digit-band (slice) staging that emulates $b(n)$ bits of mantissa precision using $w$-bit words, incurring only polylogarithmic overhead; This leads to a soft-quadratic bit cost $\tilde{O}(n^{2})$ (e.g., $\Theta(n^{2}\log^{2}n)$ when $b(n)=\Theta(\log n)$ in the basic $2\times 2$ recursion).
@@
     Depth-uniform boundedness.
-If graded projection gives a depth-uniform bound on the target/tail bands (Part IV master condition or BWBM), then a choice of extractor parameters dependent only on depth (e.g., $\beta\propto 2^{\sigma}$) suffices.
+If graded projection gives a depth-uniform bound on the target/tail bands (Part IV master condition or BWBM), then a choice of extractor parameters dependent only on the node-local lattice exponent $\sigma$ (e.g., $\beta=\beta(\sigma)$) suffices.
   2. 2.
@@
 Fix $D\in\mathbb{N}$. Define $\mathcal{L}_{D}:=D^{-1}\mathbb{Z}[1/2]$. All matrices in Part I are over $\mathcal{L}_{D}$ entrywise.
-Definition 3.2 (Scaled coefficient lattice).
-Recursive invariant amplification introduces dyadic scalings at each node.
-To maintain exactness at a recursion node with cumulative scaling exponent $\sigma\in\mathbb{Z}_{\geq 0}$, we define the local scaled coefficient lattice:
+Definition 3.2 (Scaled coefficient lattice and node-local lattice exponent).
+Fix $D\in\mathbb{N}$ and recall $\mathcal{L}_{D}=D^{-1}\mathbb{Z}[1/2]$ (Definition 3.1).
+At each recursion node (dimension $m$), after centering and dyadic $\ell_{2}$ normalization have produced normalized operands $A^{\prime},B^{\prime}\in\mathcal{L}_{D}^{m\times m}$, we define the node-local coefficient-lattice exponent $\sigma\in\mathbb{Z}_{\geq 0}$ as the smallest integer such that all dot-product terms $A^{\prime}_{ir}B^{\prime}_{rj}$ lie in $D^{-2}2^{-\sigma}\mathbb{Z}$.
+Write each stored scalar in mantissa/exponent-offset form $x=(w/D)2^{e}$ (Appendix B.1) and denote its exponent offset by $\exp(x):=e$, with the convention $\exp(0):=+\infty$.
+Define for each inner index $r\in\{1,\dots,m\}$:
+$a_{r}$ $:=$ $\min_{1\leq i\leq m}\exp(A^{\prime}_{ir}),$ (2)
+$b_{r}$ $:=$ $\min_{1\leq j\leq m}\exp(B^{\prime}_{rj}),$ (3)
+and let:
+$\mathcal{R}~:=~\{r\in\{1,\dots,m\}:a_{r}<\infty~\text{and}~b_{r}<\infty\}.$ (4)
+If $\mathcal{R}=\emptyset$ (i.e., for every $r$ either $A^{\prime}_{:,r}\equiv 0$ or $B^{\prime}_{r,:}\equiv 0$, so all dot-product terms vanish), then set $\sigma:=0$. Otherwise define:
+$g_{\min}~:=~\min_{r\in\mathcal{R}}(a_{r}+b_{r}),\\ ~~~\sigma:=\max\{0,\,-g_{\min}\}.$ (5)
+The scaled coefficient lattice at this node is then:
 $\mathcal{L}_{\mathrm{coeff}}(\sigma)~:=~D^{-2}2^{-\sigma}\mathbb{Z}.$ (6)
-The rounding operator relative to this lattice is defined via the lifted integers:
+With this choice of $\sigma$, every band coefficient $U,T$ in the three-band identity (Proposition 5.7) for products of normalized blocks lies in $\mathcal{L}_{\mathrm{coeff}}(\sigma)$, so the scaled rounding operator is well-defined:
 $\frac{1}{D^{2}2^{\sigma}}\left\lfloor D^{2}2^{\sigma}z\right\rceil~:=~\frac{1}{D^{2}2^{\sigma}}\left\lfloor D^{2}2^{\sigma}z\right\rceil.$ (7)
-If $z\in\mathcal{L}_{\mathrm{coeff}}(\sigma)$, then $\frac{1}{D^{2}2^{\sigma}}\left\lfloor D^{2}2^{\sigma}z\right\rceil=z$ exactly.
+The scaled sieve extractor is:
+$\mathcal{S}_{\beta,\sigma}(z)~:=~\frac{1}{D^{2}2^{\sigma}}\left\lfloor D^{2}2^{\sigma}z\right\rceil-\beta\frac{1}{D^{2}2^{\sigma}}\left\lfloor D^{2}2^{\sigma}z/\beta\right\rceil$ (8)
+(cf., Definition 5.1 and Appendix B.6).
 Definition 3.3 (Digit-Band Staging).
@@
 The gap inequalities imply the noise terms are strictly bounded by $1/2$. Thus, nearest-integer rounding gives:
-$\left\lfloor Mz/\beta\right\rceil~=~MU\quad\text{and}\quad\left\lfloor Mz\right\rceil~=~\beta(MU)+MT.$ (15)
+$\left\lfloor Mz/\beta\right\rceil~=~MU~\text{and}~\left\lfloor Mz\right\rceil~=~\beta(MU)+MT.$ (20)
 Dividing by $M$ recovers $U$ and $\beta U+T$. The sieve subtraction $(\beta U+T)-\beta U$ gives $T$ exactly. ∎
@@
 6.1 Dyadic $\ell_{2}$ normalization without square roots
-Definition 6.1 (Lifted squared norm).
+Definition 6.1 (Lifted squared norm (dyadic form)).
 Let $x\in\mathcal{L}_{D}^{m}$ be a real vector whose entries lie in the dyadic lattice $\mathcal{L}_{D}$. Define its lifted squared norm by:
-$S(x)~:=~\sum_{r=1}^{m}\left(Dx_{r}\right)^{2}\in\mathbb{Z}_{\geq 0}.$ (34)
-Equivalently, if $x_{r}=(w_{r}/D)2^{e_{r}}$ with $w_{r}\in\mathbb{Z}$, then $Dx_{r}=w_{r}2^{e_{r}}$ is dyadic and $S(x)$ is the exact squared $\ell_{2}$ norm of the lifted vector $Dx$.
+$S(x)~:=~\sum_{r=1}^{m}\left(Dx_{r}\right)^{2}\in\mathbb{Z}[1/2]_{\geq 0}.$ (39)
+Equivalently, if $x_{r}=(w_{r}/D)2^{e_{r}}$ with $w_{r}\in\mathbb{Z}$ and $e_{r}\in\mathbb{Z}$, then $Dx_{r}=w_{r}2^{e_{r}}$ and $S(x)=\sum_{r}w_{r}^{2}2^{2e_{r}}$.
+For implementation it is convenient to form an integer lift. Let $t:=\max\{0,-\min_{r}e_{r}\}\in\mathbb{Z}_{\geq 0}$ and define
+$\widetilde{S}(x)~:=~\sum_{r=1}^{m}\bigl(w_{r}2^{e_{r}+t}\bigr)^{2}\in\mathbb{Z}_{\geq 0},$ (40)
+then $S(x)=2^{-2t}\widetilde{S}(x)$ exactly.
 Definition 6.2 (Dyadic scaling exponent).
@@
 Proposition 6.3 (Implementability of $k(x)$).
-Let $x\in\mathcal{L}_{D}^{m}$ and define $S(x)\in\mathbb{Z}_{\geq 0}$ and $k(x)\in\mathbb{Z}_{\geq 0}$ as in Definitions 6.16.2, then:
+Let $x\in\mathcal{L}_{D}^{m}$ and define its lifted squared norm $S(x)\in\mathbb{Z}[1/2]_{\geq 0}$ and integer lift $\widetilde{S}(x)\in\mathbb{Z}_{\geq 0}$ as in Definition 6.1.
+Let $k(x)\in\mathbb{Z}_{\geq 0}$ be the dyadic scaling exponent from Definition 6.2, then:
   1. 1.
@@
   2. 2.
-    (No square roots needed.) If $S(x)>0$ and $\ell:=\lfloor\log_{2}S(x)\rfloor$ (bitlength minus one), let $k_{0}:=\lfloor\ell/2\rfloor$, then $k(x)$ is either $k_{0}$ or $k_{0}+1$, i.e.:
-    $k(x)~=~\begin{cases}k_{0},&\text{if}~S(x)\leq 2^{2k_{0}},\\ k_{0}+1,&\text{otherwise.}\end{cases}$ (37)
-    Equivalently, $k(x)=\left\lceil\frac{1}{2}\log_{2}S(x)\right\rceil$ with the convention $k(0)=0$.
+    (No square roots needed (via an integer lift).) If $\widetilde{S}(x)>0$ and $\ell:=\lfloor\log_{2}\widetilde{S}(x)\rfloor$ (bitlength minus one), let $K_{0}:=\lfloor\ell/2\rfloor$ and set:
+    $K(x)~=~\begin{cases}K_{0},&\text{if}~\widetilde{S}(x)\leq 2^{2K_{0}},\\ K_{0}+1,&\text{otherwise,}\end{cases}~~k(x)~=~\max\{0,\,K(x)-t\},$ (43)
+    where $t:=\max\{0,-\min_{r}e_{r}\}$ is the shift from Definition 6.1 satisfying $S(x)=2^{-2t}\widetilde{S}(x)$.
+Equivalently, $k(x)=\max\{0,\,\lceil\frac{1}{2}\log_{2}\widetilde{S}(x)\rceil-t\}$, with $k(0)=0$.
   3. 3.
@@
 Fix $\delta\in(0,\frac{1}{2}]$.
-Assume that at a recursion node with cumulative scaling $\sigma$, the normalized operands satisfy the dimension-free dot-product bounds $|T_{ij}|\leq 1$ and $|L_{ij}|\leq 1$.
+Assume that at a recursion node with node-local coefficient-lattice exponent $\sigma$ (Definition 3.2), the normalized operands satisfy the dimension-free dot-product bounds $|T_{ij}|\leq 1$ and $|L_{ij}|\leq 1$.
 If the packing base $\beta$ satisfies the resolution-dependent condition:
@@
 ∎
-Remark 7.2 (Why a growing base is not “cheating”).
+Corollary 7.2 (Integral mode: deterministic dyadic base with fixed slack and node certificate (no tuning knobs)).
+Fix the slack constant $\delta_{0}:=2^{-3}=1/8$. In exact dyadic lattice semantics (Part I), it suffices to enforce the sufficient-base condition
+$D^{2}2^{\sigma}\left(\frac{2}{\beta}+\frac{1}{\beta^{2}}\right)\;\leq\;\frac{1}{2}-\delta_{0},$ (55)
+where $\sigma$ is the node-local coefficient-lattice exponent (Definition 3.2).
+For dyadic bases $\beta=2^{b}$ with $b\in\mathbb{Z}_{\geq 1}$, a fully deterministic choice with no tuning is: start with the candidate $b=t+3$ where $t:=\lfloor\log_{2}(D^{2}2^{\sigma})\rfloor$; if Eq.
+55 fails, increment $b$ once (doubling $\beta$).
+This rule (and the corresponding exact integer certificate) is spelled out in Appendix  B.5.
+Also, because $D^{2}2^{\sigma}$ is just a power-of-two shift of the fixed integer $D^{2}$, we have $t=\sigma+\lfloor\log_{2}(D^{2})\rfloor$, so the candidate $b=t+3$ is obtained by a single integer add once $\sigma$ is known.
+Proof.
+This is Theorem  7.1 specialized to the fixed slack $\delta=\delta_{0}$. The dyadic selection rule and the exact integer certificate are proved in Appendix  B.5. ∎
+Remark 7.3 (Why a growing base is not “cheating” (and what $\sigma$ means)).
 Condition (54) implies that $\beta$ must grow roughly in proportion to $2^{\sigma}$.
-Because $\sigma$ represents the cumulative normalization depth, $\sigma\approx\log n$.
-Consequently, the required bit-width of $\beta$ grows logarithmically ($O(\log n)$).
+There is no need for a “speed-first” policy: in exact (integral) dyadic mode we fix a small constant slack $\delta_{0}:=2^{-3}=1/8$ (not a tuning knob) and choose the dyadic base $\beta=2^{b}$ deterministically by Corollary 7.2 / Appendix  B.5, with a constant-cost integer certificate that the sufficient-base inequality holds at each node.
+Here $\sigma$ is the node-local coefficient-lattice exponent from Definition 3.2, i.e., it is determined by the smallest dyadic lattice step on which the normalized dot-product coefficients at that node actually lie.
+In worst-case analyses one may upper bound $\sigma$ by a quantity proportional to recursion depth (i.e., $O(\log n)$ in a $2\times 2$ recursion), but implementations MUST compute $\sigma$ from exponent-offset metadata of $A^{\prime},B^{\prime}$ rather than setting $\sigma:=$ (recursion depth).
+Consequently, the required bit-width of the dyadic base $\beta=2^{b}$ grows at most logarithmically in $n$ in the worst case.
 This is not a violation of the model; rather, it is the precise origin of the digit-band staging factor $K(n)$ (Theorem 9.5).
@@
 Lemma 8.1 (Bilinear centering identity).
-Let $A,B\in\mathbb{R}^{n\times n}$ and let $\alpha,\beta\in\mathbb{R}^{n}$, and define the centered matrices:
+Let $A,B\in\mathbb{R}^{n\times n}$ and let $\alpha,\rho\in\mathbb{R}^{n}$, and define the centered matrices:
 $A_{0}$ $:=$ $A-\mathbf{1}\alpha^{\top},$ (56)
-$B_{0}$ $:=$ $B-\beta\mathbf{1}^{\top},$ (50)
+$B_{0}$ $:=$ $B-\rho\mathbf{1}^{\top},$ (57)
 where $\mathbf{1}\in\mathbb{R}^{n}$ is the all-ones vector, then:
-$AB~=~A_{0}B_{0}+\mathbf{1}(\alpha^{\top}B_{0})+(A_{0}\beta)\mathbf{1}^{\top}+(\alpha^{\top}\beta)\mathbf{1}\mathbf{1}^{\top}.$ (51)
+$AB~=~A_{0}B_{0}+\mathbf{1}(\alpha^{\top}B_{0})+(A_{0}\rho)\mathbf{1}^{\top}+(\alpha^{\top}\rho)\mathbf{1}\mathbf{1}^{\top}.$ (58)
 Proof.
-From the definitions, $A=A_{0}+\mathbf{1}\alpha^{\top}$ and $B=B_{0}+\beta\mathbf{1}^{\top}$. Multiply and expand:
-$AB$ $=$ $(A_{0}+\mathbf{1}\alpha^{\top})(B_{0}+\beta\mathbf{1}^{\top})$ (52)
-$=$ $A_{0}B_{0}+A_{0}(\beta\mathbf{1}^{\top})+(\mathbf{1}\alpha^{\top})B_{0}+(\mathbf{1}\alpha^{\top})(\beta\mathbf{1}^{\top}).$ (53)
-Now simplify each term:
-$A_{0}(\beta\mathbf{1}^{\top})$ $=$ $(A_{0}\beta)\mathbf{1}^{\top},$ (54)
+From the definitions, $A=A_{0}+\mathbf{1}\alpha^{\top}$ and $B=B_{0}+\rho\mathbf{1}^{\top}$, so:
+$AB$ $=$ $(A_{0}+\mathbf{1}\alpha^{\top})(B_{0}+\rho\mathbf{1}^{\top})$ (59)
+$=$ $A_{0}B_{0}+A_{0}(\rho\mathbf{1}^{\top})+(\mathbf{1}\alpha^{\top})B_{0}+(\mathbf{1}\alpha^{\top})(\rho\mathbf{1}^{\top}).$ (60)
+Now observe that:
+$A_{0}(\rho\mathbf{1}^{\top})$ $=$ $(A_{0}\rho)\mathbf{1}^{\top},$ (61)
 $(\mathbf{1}\alpha^{\top})B_{0}$ $=$ $\mathbf{1}(\alpha^{\top}B_{0}),$ (62)
-and:
-$(\mathbf{1}\alpha^{\top})(\beta\mathbf{1}^{\top})$ $=$ $\mathbf{1}(\alpha^{\top}\beta)\mathbf{1}^{\top}$ (56)
-$=$ $(\alpha^{\top}\beta)\mathbf{1}\mathbf{1}^{\top}.$ (57)
+$(\mathbf{1}\alpha^{\top})(\rho\mathbf{1}^{\top})$ $=$ $\mathbf{1}(\alpha^{\top}\rho)\mathbf{1}^{\top}$ (63)
+$=$ $(\alpha^{\top}\rho)\mathbf{1}\mathbf{1}^{\top}.$ (64)
 Substituting gives the stated identity. ∎
@@
 We refer to this as the row(A)-col(B) policy.
-It implies the entrywise bound $|(A^{\prime}B^{\prime})_{ij}|\leq 1$ by Cauchy–Schwarz, which is the invariant used to justify a constant packing base $\beta_{\star}$.
+It implies the entrywise bound $|(A^{\prime}B^{\prime})_{ij}|\leq 1$ by Cauchy–Schwarz.
+This bound is the invariant used to justify the node-local dyadic base choice $\beta=\beta(\sigma)$ via Theorem 7.1 once the node-local lattice exponent $\sigma$ is computed from $A^{\prime},B^{\prime}$ (Definition 3.2).
 Remark 8.3 (What is centered vs what is normalized).
@@
 These operations generally do not commute, and we cannot in general enforce zero-mean rows and columns and unit row and column norms simultaneously.
-Throughout, we use the row(A)–col(B) policy (Definition 8.2) specifically to guarantee the entrywise bound $|(A^{\prime}B^{\prime})_{ij}|\leq 1$ needed for a constant $\beta_{\star}$.
+Throughout, we use the row(A)–col(B) policy (Definition 8.2) to guarantee the entrywise bound $|(A^{\prime}B^{\prime})_{ij}|\leq 1$ needed for the node-local dyadic base choice $\beta=\beta(\sigma)$ once $\sigma$ is computed from exponent offsets (Definition 3.2).
 Mean constraints propagate only in the directions explicitly centered: if $B^{\prime}1=0$ (zero row means of $B^{\prime}$), then $(A^{\prime}B^{\prime})1=0$ (zero row means of the product).
@@
 $\|B^{\prime}_{:,j}\|_{2}$ $\leq$ $1~\text{for all columns}~j.$ (68)
-This choice is deliberate because it implies the entrywise bound $|(A^{\prime}B^{\prime})_{ij}|\leq 1$, which is the invariant needed to justify a constant packing base $\beta_{\star}$.
+This choice is deliberate because it implies the entrywise bound $|(A^{\prime}B^{\prime})_{ij}|\leq 1$, which is the invariant needed to justify the node-local dyadic base choice $\beta=\beta(\sigma)$ once $\sigma$ is computed (Definition 3.2; Appendix B.5).
   • Do not attempt to enforce all row/column norm constraints simultaneously.
@@
   2. 2.
-    No square roots: Dyadic $\ell_{2}$ normalization uses $S(x)=\sum(Dx_{r})^{2}$ and $k(x)$ via bitlength comparisons (Proposition 6.3). Scaling by $2^{\pm k}$ is an exponent-offset update.
+    No square roots: Dyadic $\ell_{2}$ normalization uses the lifted squared norm $S(x)\in\mathbb{Z}[1/2]_{\geq 0}$ and its integer lift $\widetilde{S}(x)\in\mathbb{Z}_{\geq 0}$ (Definition 6.1), and computes $k(x)$ via bitlength comparisons on $\widetilde{S}(x)$ (Proposition 6.3).
+Scaling by $2^{\pm k}$ is an exponent-offset update.
   3. 3.
-    Sufficient base: The packing base $\beta$ is chosen to satisfy Theorem 7.1 given the node’s scaling $\sigma$ (e.g., via the allocated staging budget).
+    Sufficient base with computed $\sigma$ (no depth substitution).
+At each node, $\sigma$ MUST be computed from the exponent offsets of the normalized operands $A^{\prime},B^{\prime}$ as in Definition 3.2 (equivalently Eq.
+266).
+Implementations MUST NOT set $\sigma$ equal to recursion depth (this can over-inflate $\beta$ and destroy performance).
+Given the computed $\sigma$, implementations MUST use the fixed slack convention $\delta_{0}:=1/8$ and set the dyadic packing base $\beta=2^{b}$ deterministically by Appendix B.5.
+A reference implementation MUST certify the sufficient-base inequality Eq.
+263 for the chosen $\beta$.
+If it fails at the initial candidate, apply the one-step upgrade rule $b\leftarrow b+1$ (doubling $\beta$) and re-evaluate.
+If it still fails after the one-step upgrade, the implementation MUST take the failure-policy fallback of Appendix  B.5 (call ClassicalMM on the original subproblem and return) and MUST NOT apply the sieve at that node.
+(Equivalently, in staged-precision mode an implementation may signal failure to a retry wrapper and re-attempt with larger staging; see Algorithm  2.)
   4. 4.
-    Shifts do not touch mantissas: Multiplication by $2^{\pm t}$ and by $\beta_{\star}^{\pm 1}$ must be implemented as exponent updates, not as integer multiplication on mantissas.
+    Shifts do not touch mantissas: Multiplication by $2^{\pm t}$ and by $\beta^{\pm 1}$ must be implemented as exponent updates, not as integer multiplication on mantissas.
   5. 5.
@@
  return ABAB
-(1) Center. Compute the exact column-mean vector $\alpha\in\mathcal{L}_{D}^{n}$ of $A$ and exact row-mean vector $\beta\in\mathcal{L}_{D}^{n}$ of $B$ using dyadic division:
+(1) Center. Compute the exact column-mean vector $\alpha\in\mathcal{L}_{D}^{n}$ of $A$ and exact row-mean vector $\rho\in\mathcal{L}_{D}^{n}$ of $B$ using dyadic division:
 $\alpha_{j}$ $:=$ $\frac{1}{n}\sum_{i=1}^{n}A_{ij},$
-$\beta_{i}$ $:=$ $\frac{1}{n}\sum_{j=1}^{n}B_{ij}.$
-Set $A_{0}:=A-\mathbf{1}\alpha^{\top}$ and $B_{0}:=B-\beta\mathbf{1}^{\top}$.
+$\rho_{i}$ $:=$ $\frac{1}{n}\sum_{j=1}^{n}B_{ij}.$
+Set $A_{0}:=A-\mathbf{1}\alpha^{\top}$ and $B_{0}:=B-\rho\mathbf{1}^{\top}$.
 (2) Dyadic $\ell_{2}$ normalize (sqrt-free). For each row $i$ of $A_{0}$, compute $k_{i}:=k(A_{0,i,:})$ via Definitions 6.16.2, and scale:
@@
 Store the exponents $\{k_{i}\}_{i=1}^{n}$ and $\{\ell_{j}\}_{j=1}^{n}$ for the undo step.
-(3) Pack, recurse, and extract quadrants.
-Partition $A^{\prime},B^{\prime}$ into $2\times 2$ blocks of size $(n/2)\times(n/2)$.
-Choose base $\beta$ satisfying Theorem 7.1 given current scaling $\sigma$.
+(2b) Compute the node-local lattice exponent $\sigma$ (no extra passes). Using exponent offsets $\exp(\cdot)$ in the representation $x=(w/D)2^{e}$ (Appendix B.1), compute:
+$a_{r}$ $:=$ $\min_{i}\exp(A^{\prime}_{ir}),~b_{r}:=\min_{j}\exp(B^{\prime}_{rj}),$ (69)
+$\mathcal{R}$ $:=$ $\{\,r:a_{r}<\infty~\text{and}~b_{r}<\infty\,\}.$ (70)
+If $\mathcal{R}=\emptyset$, set $\sigma:=0$; otherwise set $g_{\min}:=\min_{r\in\mathcal{R}}(a_{r}+b_{r})$ and $\sigma:=\max\{0,-g_{\min}\}$.
+This is exactly Definition 3.2 / Eq.
+266, and can be accumulated during Step (2).
+(3) Pack, recurse, and extract quadrants. Partition $A^{\prime},B^{\prime}$ into $2\times 2$ blocks of size $(n/2)\times(n/2)$.
+/* Certified base selection (no tuning knobs; at most one upgrade). */
+Compute the candidate exponent $b_{0}$ (i.e., $\beta_{0}=2^{b_{0}}$) from Eq. 270. Set $b\leftarrow b_{0}$ and $\beta\leftarrow 2^{b}$
+Evaluate the exact integer certificate Eq. 271 (and, in a reference implementation, assert the normalization bounds from Step (2))
+if the certificate fails at $b=b_{0}$ then
+ $b\leftarrow b_{0}+1$$\beta\leftarrow 2^{b}$
+ // one-step upgrade, Eq. 272
+   Re-pack using the upgraded $\beta$ and re-evaluate the certificate (in exact integer arithmetic this second check must succeed; if it does not, treat it as overflow/truncation)
+ if the certificate still fails then
+    return ClassicalMM(A,B)(A,B)
+    // exact fallback; Appendix B.5
 foreach $(i,j)\in\{1,2\}^{2}$ do
@@
 and then recover $C=AB$ using the bilinear centering identity (Lemma 8.1):
-$C~:=~A_{0}B_{0}+\mathbf{1}(\alpha^{\top}B_{0})+(A_{0}\beta)\mathbf{1}^{\top}+(\alpha^{\top}\beta)\mathbf{1}\mathbf{1}^{\top}.$
+$C~:=~A_{0}B_{0}+\mathbf{1}(\alpha^{\top}B_{0})+(A_{0}\rho)\mathbf{1}^{\top}+(\alpha^{\top}\rho)\mathbf{1}\mathbf{1}^{\top}.$
 return $C$
@@
 9 Complexity and implementation model
+Remark 9.1 (Certified regime and staged retry).
+In exact dyadic lattice semantics, Appendix  B.5 proves that the certified base-selection rule (Eq.
+271 with the one-step upgrade Eq.
+272) succeeds after at most one upgrade at every node, provided coefficient-lattice alignment is preserved (i.e., no mantissa bits are dropped below the local $\mathcal{L}_{\mathrm{coeff}}(\sigma)$ resolution).
+Therefore, within the declared exact-execution regime, the ClassicalMM branch in Algorithm  1 is unreachable and serves only as a correct-on-return guard for overflow/truncation (or implementation bugs).
+On fixed-width hardware, if a run ever leaves this regime, one may either take the conservative ClassicalMM fallback, or increase the staged-precision budget / word width and retry.
+Algorithm  2 formalizes a simple retry-with-more-staging wrapper, and Theorem  9.2 shows that it preserves the soft-quadratic bit/word complexity when the required precision $b(n)$ is polylogarithmic.
+9.1 Retry with more staging (optional)
+The ClassicalMM branch in Algorithm  1 is a correct-on-return backstop for implementations that leave the declared exact-execution regime (Appendix  B.5).
+In staged-precision fixed-word mode, a more efficient response is often to increase the staged mantissa budget and rerun the computation.
+The wrapper below does so without changing the kernel logic.
+Attempt primitive.
+Let KernelMM_Attempt(A,B;K)(A,B;K) denote Algorithm  1 executed in digit-band staged mode with $K$ words per scalar, except that the failure-policy fallback of Appendix  B.5 is replaced by returning a failure flag and aborting the attempt (i.e., no extracted value computed under a failed certificate is ever returned).
+Input: $A,B\in\mathcal{L}_{D}^{n\times n}$ with $n$ a power of two; initial staging $K_{0}\geq 1$; maximum staging budget $K_{\max}$
+Output: $C=AB$ exactly (or exact fallback if the staging limit is exhausted)
+$K\leftarrow K_{0}$
+while $K\leq K_{\max}$ do
+ $(C,\mathrm{ok})\leftarrow\textsc{KernelMM\_Attempt}(A,B;K)$
+ if $\mathrm{ok}$ then
+    return $C$
+ $K\leftarrow 2K$
+return ClassicalMM(A,B)(A,B)
+// correct-on-return backstop
+[…]
+Theorem 9.2 (Retry wrapper: correctness and complexity).
+Fix a base word width $w$ and consider digit-band (slice) staging with $K$ words per scalar.
+Assume that for a full run of Algorithm  1 under $K$-word staging in the certified (no-failure) regime, the fixed-word cost satisfies $B(n,K)=\Theta(n^{2}\log n\cdot K)$ (each scalar primitive costs $\Theta(K)$ word operations).
+[…]
+Proof.
+[…]
+Correctness.
+On the successful iteration, no node triggers the failure policy, so the attempted execution follows exactly the same computation path as Algorithm  1 and satisfies the hypotheses of Theorem  8.5, so it returns $C=AB$ exactly.
+If instead the wrapper exhausts the staging limit ($K>K_{\max}$), it returns ClassicalMM(A,B)(A,B), which is exact by Algorithm  3.
+Cost.
+For each failed attempt with budget $K_{i}$, the wrapper performs at most the work of a full staged run at the same $K_{i}$. Therefore the total cost is bounded by:
+$\sum_{i=0}^{s}B(n,K_{i})~=~\Theta(n^{2}\log n)\sum_{i=0}^{s}K_{i}~=~\Theta(n^{2}\log n)\,(2K_{s}-K_{0})~=~O(n^{2}\log n\cdot K_{s}).$ (74)
+Because $K_{s}<2K^{\star}$, this is $O(n^{2}\log n\cdot K^{\star})=O(B(n,K^{\star}))$, and likewise $O(B(n,2K^{\star}))$.
+The bound on the number of attempts follows from $K_{s}\geq K^{\star}$ and $K_{s}=2^{s}K_{0}$.
+∎
 9.2 Arithmetic recurrence
@@
 Theorem 9.4 (Arithmetic complexity).
-Let $T(n)$ denote the arithmetic time of KernelMM on $n\times n$ inputs, measured as the number of scalar ring operations (additions, multiplications by fixed constants such as powers of two, and nearest-integer rounding calls), assuming $n$ is a power of two, then:
+Let $T(n)$ denote the arithmetic time of KernelMM on $n\times n$ inputs, measured as the number of scalar ring operations (additions, multiplications by fixed constants such as powers of two, and nearest-integer rounding calls), assuming $n$ is a power of two and that Appendix  B.5’s certificate-and-upgrade rule succeeds at every recursion node (so the algorithm never takes the ClassicalMM fallback), then:
 $T(n)~=~4T(n/2)+O(n^{2}),$ (75)
@@
 Proof.
-We now consider a single call of KernelMM on size $n$.
+We now consider a single call of KernelMM on size $n$ under the assumption that the certificate-and-upgrade rule of Appendix  B.5 succeeds at this node and recursively at all descendants (so no fallback occurs).
 Recursive calls:
@@
 Fix a word width $w$ for mantissa storage, and assume exponent offsets and loop indices fit in fixed-width machine integers.
+Assume additionally that Appendix  B.5’s certificate-and-upgrade rule succeeds at every recursion node (so the algorithm never takes the ClassicalMM fallback).
 Let $b(n)$ denote the required effective mantissa precision (in bits) needed to represent packed temporaries and preserve the extractor slack $\delta$ at all decision sites for dimension $n$ under the chosen execution discipline.
@@
 Theorem 9.12 (Correctness and complexity of GPR convolution).
-There exists a sufficient base sequence $\beta(\sigma)$ (growing with recursion depth $\sigma$) such that the recursive packed convolution algorithm recovers $f\ast g$ exactly in dyadic lattice semantics.
+There exists a sufficient base sequence $\beta(\sigma)$ (growing with the node-local lattice exponent $\sigma$) such that the recursive packed convolution algorithm recovers $f\ast g$ exactly in dyadic lattice semantics.
 The arithmetic recurrence is:
@@
 Dyadic normalization (Lemma 9.9) ensures the band coefficients are bounded in the normalized lattice $\mathcal{L}_{\mathrm{coeff}}(\sigma)$.
-By choosing $\beta$ to satisfy the sufficient-base condition (Theorem 7.1) for the current depth $\sigma$, the master gap inequality holds, and the scaled sieve $\mathcal{S}_{\beta,\sigma}$ recovers the middle band $T$ exactly.
+By choosing $\beta$ to satisfy the sufficient-base condition (Theorem 7.1) for the current node-local lattice exponent $\sigma$, the master gap inequality holds, and the scaled sieve $\mathcal{S}_{\beta,\sigma}$ recovers the middle band $T$ exactly.
 The algorithm performs $4$ subcalls (implicitly or explicitly to recover all components of the convolution from the packed forms), maintaining exactness throughout via the staged precision model.
@@
 Variants of this idea appear across several communities.
-What is new in this note is not the existence of coefficient/band extraction, but the way recursive invariant amplification (centering plus dyadic $\ell_{2}$ normalization at every recursion node) makes coefficient extraction compatible with a model-honest intermediate-budget analysis: it gives a constant base $\beta_{\star}$, eliminates bit-growth of mantissas, and improves feasibility/robustness by bounding intermediates at every depth rather than only at the root.
+What is new in this note is not the existence of coefficient/band extraction, but the way recursive invariant amplification (centering plus dyadic $\ell_{2}$ normalization at every recursion node) makes coefficient extraction compatible with a model-honest intermediate-budget analysis: it gives a node-local dyadic base policy $\beta=\beta(\sigma)$ (Appendix B.5), eliminates bit-growth of mantissas, and improves feasibility/robustness by bounding intermediates at every depth rather than only at the root.
 10.1 APA and border-rank methods in fast matrix multiplication
 In fast matrix multiplication theory, approximate (APA) bilinear algorithms introduce a formal parameter (often written $\lambda$ or $\varepsilon$) to separate contributions across different powers of that parameter and then recover a desired coefficient via an exactification Step (interpolation/limit argument).
-This is similar to the work of Bini et al.
-on approximate MM [18] and in Bini’s later discussion relating exact and approximate bilinear algorithms [19].
+This is similar to the work of Bini et al., on approximate MM [18] and in Bini’s later discussion relating exact and approximate bilinear algorithms [19].
 A closely related tensor-language view is that such constructions upper-bound the border rank of the MM tensor, which is sufficient to derive asymptotic exponents (see [20] and Pan’s trilinear/field-extension methods [21]).
@@
 In Part IV we abstract this as a general admissibility condition (master condition/BWBM) coupled to an extractor interface.
-Part II Float-GPR: baseline comparators and commensurate accuracy (Part II)
+Part II Float-GPR: baseline comparators and commensurate accuracy
 11 Scope and goals
@@
 $T_{\mathsf{F}}^{(\text{no-ext})}(n)~=~\Theta(n^{2}\log n).$ (136)
-Part III Case studies and templates (Part III)
+Part III Case studies and templates
 The Graded Projection Recursion (GPR) framework provides a set of tools for reducing the complexity of certain algorithms that satisfy certain conditions.
@@
 ∎
-17.3 Triangle detection via Boolean MM (semiring mode)
-Theorem 17.4 (Triangle detection via Boolean GPR kernel (sketch)).
-This theorem is a template statement. It assumes a semiring matrix kernel with bounded per-entry representation cost under graded projection.
-Let $G$ be a directed or undirected graph on $n$ vertices with adjacency matrix $A\in\{0,1\}^{n\times n}$.
-Work over the Boolean semiring (,)=(,)($\oplus$$\otimes$)=($\lor$$\land$), so Boolean matrix multiplication is:
-$(X\otimes Y)_{ij}~=~\bigvee_{k=1}^{n}\left(X_{ik}\wedge Y_{kj}\right).$ (171)
-Assume Boolean matrix multiplication is implemented by a recursive blocked GPR kernel KernelMatMulBool that uses graded projection corresponding to word-packing of bits (i.e., each row/column block is stored as bitsets packed into machine words and block products are computed by word-level operations), then:
-  1. 1.
-    (Correctness) $G$ contains a triangle if and only if there exist $i,j$ such that:
-    $(A\otimes A)_{ij}\wedge A_{ji}~=~1.$ (172)
-    Equivalently, in Boolean arithmetic, $\mathrm{trace}(A\otimes A\otimes A)=1$ if and only if a triangle exists.
-  2. 2.
-    (Complexity reduction to Boolean MM) Triangle detection reduces to one Boolean matrix multiplication plus $O(n^{2})$ additional work for the final check, so the runtime is $\widetilde{O}(T_{\mathrm{BoolMM}}(n))$, where $T_{\mathrm{BoolMM}}(n)$ is the cost of one Boolean matrix multiplication implemented by KernelMatMulBool.
-In particular, with word-packing and a cache-friendly blocked implementation one obtains $O(n^{3}/w)$ word operations (where $w$ is the machine word size).
-Proof.
-Correctness
-A triangle exists if and only if there are vertices $i,j,k$ with edges $(i,k)$, $(k,j)$, and $(j,i)$.
-The Boolean product $(A\otimes A)_{ij}=1$ holds exactly when there exists $k$ such that $A_{ik}=1$ and $A_{kj}=1$, i.e., a length-2 path from $i$ to $j$.
-Thus $(A\otimes A)_{ij}\wedge A_{ji}=1$ holds exactly when there is a length-2 path $i\to j$ and an edge $j\to i$, which is precisely a directed triangle.
-The trace characterization is equivalent.
-Complexity
-Computing $B:=A\otimes A$ is one Boolean matrix multiplication, and checking whether any entry satisfies $B_{ij}\wedge A_{ji}=1$ is $O(n^{2})$, so the total runtime is $\widetilde{O}(T_{\mathrm{BoolMM}}(n))$.
-With word-packing, each inner product reduces to a wordwise AND and an “any-nonzero” test (or popcount), giving the standard $O(n^{3}/w)$ bound.
-∎
 17.3 How to create new GPR reductions
@@
     Conclude complexity by a recurrence. Conclude correctness/commensurate accuracy by composition.
-Part IV Generality of GPR: admissible packings, design patterns, and extensions (Part IV)
+Part IV Generality of GPR: admissible packings, design patterns, and extensions
 18 What is GPR really doing?
@@
 The extraction is then correct at every depth and the corresponding GPR reduction inherits the same asymptotic bit/format complexity, up to constants depending on $C$ and on the extractor parameters.
-Proof sketch.
-Uniform master bounds imply a uniform sufficient condition for extraction once global parameters are chosen (base separation, CRT interval, bucket width, etc.).
-The extraction is then correct at every leaf, and correctness propagates up the recursion by the same structural induction as in Parts I–III.
-Complexity is unchanged asymptotically because extractor operations are $O(1)$ per entry per level (or bounded by a fixed number of moduli/buckets).
+Proof.
+Choose global packing parameters (e.g., bases $\beta$ or a fixed set of moduli) as in the hypothesis so that, for every depth $\ell$ and every packed value $Z$ produced at depth $\ell$, the uniform master bound:
+$\|T_{\ell}\|_{N}\leq CS_{\ell}\leq CS_{\max}~\text{(or}~\|T_{\ell,E_{0}}\|_{N}\leq CS_{\ell}\leq CS_{\max}\text{)}$ (246)
+implies $\mathsf{Cond}(Z)$. (The second inequality uses the definition of $S_{\max}:=\max_{\ell}S_{\ell}$.)
+Extraction is correct at every node.
+Fix any depth $\ell$ and any packed value $Z$ arising at depth $\ell$. By the preceding choice of parameters, $\mathsf{Cond}(Z)$ holds. By the extractor interface (Definition 21.8):
+$\mathrm{Extract}(Z)~=~\mathrm{TargetPart}(Z),$ (247)
+i.e., the extractor returns exactly the intended target coefficient block (degree band or multi-degree band).
+Correctness propagates through the recursion.
+We argue by structural induction over the recursion tree of the GPR reduction.
+At a leaf node, the reduction evaluates the kernel on the (unpacked) leaf inputs directly and returns the result in the native domain; this equals the corresponding classical computation by definition.
+For an internal node $u$ at depth $\ell$, let $(x_{1},\dots,x_{k})$ denote the node-local inputs in the native domain. The reduction forms packed embeddings:
+$R_{i}(\boldsymbol{\beta})~=~\sum_{p\in D_{i}}\boldsymbol{\beta}^{p}R_{i}^{[p]}~~~(i=1,\dots,k)$ (248)
+as in Definition 21.5, and computes the packed kernel value
+$Z~:=~\mathcal{B}\left(R_{1}(\boldsymbol{\beta}),\dots,R_{k}(\boldsymbol{\beta})\right).$ (249)
+By $k$-linearity of $\mathcal{B}$ (Definition 21.4), expanding gives:
+$Z~=~\sum_{(p_{1},\dots,p_{k})\in D_{1}\times\cdots\times D_{k}}\boldsymbol{\beta}^{p_{1}+\cdots+p_{k}}~\mathcal{B}\left(R_{1}^{[p_{1}]},\dots,R_{k}^{[p_{k}]}\right).$ (250)
+By construction of the packing scheme, the coefficient block(s) indexed by the target band (the degree or multi-degree window specified in the reduction) is exactly the value that the classical, unpacked computation at node $u$ is meant to return; this coefficient block is $\mathrm{TargetPart}(Z)$ by definition.
+By the first part of the proof, extraction at node $u$ is correct, so the reduction returns $\mathrm{Extract}(Z)=\mathrm{TargetPart}(Z)$, i.e., the correct classical node-local output.
+This completes the induction.
+Therefore the output returned at the root of the recursion coincides with the intended classical output.
+Asymptotic complexity is preserved.
+At a node operating on blocks of side length $m$, packing, extraction, and the associated bookkeeping (centering, normalization, digit/CRT/bucket operations) act entrywise and perform a number of primitive operations bounded by a constant $\kappa$ that depends only on the chosen global parameters (i.e., only on $C$ and $S_{\max}$ in the hypothesis), and not on $m$ or on the depth.
+Thus the nonrecursive work at a node is $O(\kappa\,m^{2})$, and the recursion has the same form as the baseline blocked recursion up to the multiplicative factor $\kappa$.
+For instance, under $2\times 2$ block recursion for matrix multiplication,
+$T(m)~=~4T(m/2)+O(\kappa\,m^{2}),$ (251)
+so $T(n)=O(\kappa\,n^{2}\log n)$.
+The same reasoning applies to bit/format complexity: the dependence on the packing/extractor parameters is absorbed into $\kappa$, and the asymptotic dependence on problem size follows the same recurrence.
 ∎
@@
 Definition B.6 (Exact sieve extractor for $\beta=2^{b}$).
-Fix an integer $b\geq 1$ and let $\beta:=2^{b}$. For $z\in\mathcal{L}_{D}$, define:
-$\mathcal{S}_{\beta}(z)~:=~\frac{1}{D^{2}}\left\lfloor D^{2}(\right\rceil z)-\beta\frac{1}{D^{2}}\left\lfloor D^{2}(\right\rceil z/\beta).$ (245)
-Here $z/\beta$ is implemented by exponent update $e\leftarrow e-b$ (exact in $\mathcal{L}_{D}$), and $\frac{1}{D^{2}}\left\lfloor D^{2}(\right\rceil\cdot)$ is implemented by Definition B.4.
-Lemma B.7 (Implementation matches Part I sieve).
-If $\mathcal{S}_{\beta}$ is implemented as in Definition B.6, then it matches the Part I definition $\mathcal{S}_{\beta}(z)=\frac{1}{D^{2}}\left\lfloor D^{2}(\right\rceil z)-\beta\frac{1}{D^{2}}\left\lfloor D^{2}(\right\rceil z/\beta)$ exactly, and Lemmas 5.2 and 5.3 apply verbatim.
+Fix an integer $b\geq 1$ and let $\beta:=2^{b}$.
+Let $\sigma\geq 0$ be the node-local coefficient-lattice exponent at the current recursion node (Definition 3.2; Eq.
+266).
+For $z\in\mathcal{L}_{D}$, define:
+$\mathcal{S}_{\beta,\sigma}(z)$ $:=$ $\frac{1}{D^{2}2^{\sigma}}\left\lfloor D^{2}2^{\sigma}z\right\rceil-\beta\frac{1}{D^{2}2^{\sigma}}\left\lfloor D^{2}2^{\sigma}z/\beta\right\rceil$ (259)
+$=$ $\frac{1}{D^{2}2^{\sigma}}\left\lfloor D^{2}2^{\sigma}z\right\rceil-\beta\frac{1}{D^{2}2^{\sigma}}\left\lfloor D^{2}2^{\sigma}(z/\beta)\right\rceil.$ (260)
+Here $z/\beta$ is implemented by exponent update $e\leftarrow e-b$ (exact in $\mathcal{L}_{D}$), and the rounding $\frac{1}{M}\left\lfloor M(\cdot)\right\rceil$ is implemented by rounding the mantissa to nearest integer at the combined exponent $e_{\mathrm{total}}=e+\sigma$.
+Lemma B.7 (Implementation matches Part I sieve (with scaling)).
+If $\mathcal{S}_{\beta,\sigma}$ is implemented as in Definition B.6, then it matches Definition 5.1 exactly, i.e.:
+$\mathcal{S}_{\beta,\sigma}(z)~=~\frac{1}{D^{2}2^{\sigma}}\left\lfloor D^{2}2^{\sigma}z\right\rceil-\beta\,\frac{1}{D^{2}2^{\sigma}}\left\lfloor D^{2}2^{\sigma}(z/\beta)\right\rceil.$ (261)
+In particular, Lemmas 5.25.3 apply verbatim to the implemented extractor.
 Proof.
-All steps are exact in $\mathcal{L}_{D}$: dividing by $\beta=2^{b}$ is an exponent-offset update, and $\frac{1}{D^{2}}\left\lfloor D^{2}(\right\rceil\cdot)$ is exact by the preceding lemma.
-Therefore the implemented $\mathcal{S}_{\beta}$ equals the mathematical sieve definition.
+All steps are exact in $\mathcal{L}_{D}$ (i.e., dividing by $\beta=2^{b}$ is an exponent-offset update) and $\frac{1}{D^{2}}\left\lfloor D^{2}(\right\rceil\cdot)$ is exact by the preceding lemma.
+Therefore the implemented $\mathcal{S}_{\beta}$ satisfies the mathematical sieve definition.
 ∎
-B.5 Gap checking and failure policy (optional but recommended)
-For debugging and for reference implementations, it is useful to explicitly check the master gap inequality before applying the sieve.
-For a packed scalar $z=\beta U+T+L/\beta$, compute the three quantities appearing in Lemma 5.3 in the lifted units and assert:
-$\lvert D^{2}L/\beta\rvert+\lvert D^{2}T/\beta\rvert+\lvert D^{2}L/\beta^{2}\rvert~\leq~\frac{1}{2}-\delta.$ (246)
-If the check fails, the implementation should either:
-  1. 1.
-    Return a failure flag (instance leaves the declared success set), or
-  2. 2.
-    Increase $\beta$ and retry (if using an adaptive base policy outside the constant-$\beta$ regime).
-On the success set, the check passes and extraction is exact.
+B.5 Certified base choice and failure policy
+Context and goal.
+Here we give a fully deterministic, self-checking rule that makes an implementation of KernelMM (Algorithm 1) correct-on-return: before applying the sieve/extractor at a recursion node, the implementation certifies that the master-gap condition holds for the actual node-local lattice exponent $\sigma$ and the chosen packing base $\beta$.
+These checks are constant-cost (or reuse metadata already computed during normalization) and do not affect the near-quadratic scaling.
+Fixed-slack sufficient-base condition (integral mode).
+For the integral (exact dyadic) kernel we fix (once and for all) the slack constant:
+$\delta_{0}~:=~2^{-3}=1/8,$ (262)
+and we do not treat $\delta$ as a tuning parameter. At every recursion node we require the sufficient-base inequality from Theorem  7.1 with $\delta=\delta_{0}$:
+$D^{2}2^{\sigma}\left(\frac{2}{\beta}+\frac{1}{\beta^{2}}\right)~\leq~\frac{1}{2}-\delta_{0},~~\beta=2^{b},~b\in\mathbb{Z}_{\geq 1}.$ (263)
+When Eq.
+263 holds and the normalization bounds $\|A^{\prime}_{i,:}\|_{2}\leq 1$, $\|B^{\prime}_{:,j}\|_{2}\leq 1$ are satisfied, extraction is exact because the master gap in Lemma  5.3 is open by at least $\delta_{0}$.
+Tight node-local coefficient lattice exponent $\sigma$ (no over-estimation).
+Represent each stored scalar as $x=(w/D)2^{e}$ (Appendix B.1) and write $\exp(x):=e$ for its exponent offset.
+After Step (2) of Algorithm 1 has produced normalized operands $A^{\prime},B^{\prime}$, define for each inner index $r\in\{1,\dots,m\}$ (where $m$ is the node dimension):
+$a_{r}$ $:=$ $\min_{i}\exp(A^{\prime}_{ir}),$ (264)
+$b_{r}$ $:=$ $\min_{j}\exp(B^{\prime}_{rj}),$ (265)
+and let:
+$g_{\min}$ $:=$ $\min_{r:a_{r}<\infty,~b_{r}<\infty}(a_{r}+b_{r}),$ (266)
+$\sigma$ $:=$ $\max\{0,~-g_{\min}\}.$ (267)
+with the convention that if the set $\{r:a_{r}<\infty,~b_{r}<\infty\}$ is empty (i.e., every dot-product term is zero) then $\sigma:=0$.
+This ensures that every dot-product term $A^{\prime}_{ir}B^{\prime}_{rj}$ lies in the scaled coefficient lattice:
+$\mathcal{L}_{\mathrm{coeff}}(\sigma)~=~D^{-2}2^{-\sigma}\mathbb{Z}$ (268)
+according to Definition 3.2.
+Deterministic dyadic base selection with an exact integer certificate.
+Let $A:=D^{2}2^{\sigma}\in\mathbb{Z}_{\geq 1}$ and $t:=\lfloor\log_{2}A\rfloor$. Because $A$ is just a power-of-two shift of the fixed integer $D^{2}$:
+$t~=~\sigma+\left\lfloor\log_{2}(D^{2})\right\rfloor.$ (269)
+Begin with the candidate exponent:
+$b_{0}~:=~t+3~=~\sigma+\left\lfloor\log_{2}(D^{2})\right\rfloor+3,~~~\beta_{0}:=2^{b_{0}}.$ (270)
+We now certify Eq. 263 for a dyadic base $\beta=2^{b}$ with $\delta_{0}=1/8$. After clearing denominators, the inequality Eq. 263 is equivalent to the integer inequality:
+$8A(2^{b+1}+1)~\leq~3\cdot 2^{2b},$ (271)
+where Eq. 271 can be evaluated exactly using integer arithmetic (shifts and adds).
+One-step upgrade rule (guaranteed termination).
+If Eq. 271 holds at $b=b_{0}$, set $b:=b_{0}$ and $\beta:=\beta_{0}$. Otherwise, set:
+$b~:=~b_{0}+1,~~~\beta~:=~2^{b}.$ (272)
+Now Eq. 263 holds automatically for this upgraded $\beta$, and because $2^{t}\leq A<2^{t+1}$ and $b=t+4$:
+$A\left(\frac{2}{2^{b}}+\frac{1}{2^{2b}}\right)$ $\leq$ $2^{t+1}\left(\frac{2}{2^{t+4}}+\frac{1}{2^{2t+8}}\right)$ (273)
+$=$ $\frac{1}{4}+2^{-(t+7)}$ (274)
+$\leq$ $\frac{1}{4}+\frac{1}{128}$ (275)
+$<$ $\frac{3}{8}~=~\frac{1}{2}-\delta_{0}.$ (276)
+Note that the upgrade rule requires at most one doubling of $\beta$ and introduces no loops.
+In particular, in exact integer arithmetic the base certificate Eq.
+271 must pass after at most one upgrade, i.e., any observed failure after the upgrade indicates overflow/truncation (or a bug) and triggers the failure policy below.
+Failure policy (correct-on-return).
+If the node certificate fails (either the normalization bounds are violated, or Eq.
+271 fails even after the one-step upgrade), then the implementation must not apply the sieve at that node.
+Instead it must treat this as leaving the declared exact-execution regime (e.g., due to overflow, truncation below lattice resolution, or a bug) and must fall back to a conventional exact multiplication on the original subproblem, i.e., call ClassicalMM(A,B)(A,B) (Algorithm  3) and return its result.
+(An optimized implementation may instead increase the staged-precision budget / word width and restart the node; see Algorithm  2 and Theorem  9.2.) With this policy, the algorithm never returns an incorrect extracted value, i.e., failure affects only performance, not correctness.
+Explicit fallback routine (ClassicalMM).
+The fallback routine is the conventional triple-loop product evaluated in exact dyadic lattice semantics.
+A reference implementation may use a wider accumulator (or arbitrary-precision integers for the mantissas) for the inner sum, because this path is only a correctness backstop and is outside the near-quadratic performance model.
+Input: $A,B\in\mathcal{L}_{D}^{m\times m}$
+Output: $C=AB$ exactly
+for $i\leftarrow 1$ to $m$ do
+ for $j\leftarrow 1$ to $m$ do
+    $s\leftarrow 0$
+    for $r\leftarrow 1$ to $m$ do
+       $s\leftarrow s+A_{ir}B_{rj}$
+    $C_{ij}\leftarrow s$
+return $C$
+Algorithm 3 ClassicalMM(A,B)(A,B) (exact dyadic fallback)
+Mandatory call-site (Algorithm 1).
+In Algorithm 1 (KernelMM), compute $\sigma$ by Eq. 266 and set $\beta$ by the candidate Eq. 270 followed by the certificate-and-upgrade rule Eq. 271–Eq. 272.
 B.6 Notes on efficiency
-The primitives above are constant-time per scalar under fixed-width machine arithmetic:
+The primitives above are $O(1)$ per scalar under fixed-width machine arithmetic:
   • Exponent updates are constant time,

### v4

@@
 Part I gives a proof-complete integral specification of a three-band packing identity and a two-round middle-band extractor, and shows how per-node centering and sqrt-free dyadic $\ell_{2}$ normalization (recursive invariant amplification) gives a sufficient packing base $\beta(\sigma)$ whose required bit-width grows linearly with the node-local lattice exponent $\sigma$ (so only logarithmic bit-growth in $n$ in the worst case) in exact arithmetic.
-On fixed-width hardware, exact evaluation of the packed recursion generally requires either an extended-precision path or digit-band (slice) staging that emulates $b(n)$ bits of mantissa precision using $w$-bit words, incurring only polylogarithmic overhead; This leads to a soft-quadratic bit cost $\tilde{O}(n^{2})$ (e.g., $\Theta(n^{2}\log^{2}n)$ when $b(n)=\Theta(\log n)$ in the basic $2\times 2$ recursion).
+On fixed-width hardware, exact evaluation of the packed recursion generally requires either an extended-precision path or digit-band (slice) staging that emulates $b(n)$ bits of mantissa precision using $w$-bit words, incurring only polylogarithmic overhead.
+This leads to a soft-quadratic bit cost $\tilde{O}(n^{2})$ (e.g., $\Theta(n^{2}\log^{2}n)$ when $b(n)=\Theta(\log n)$ in the basic $2\times 2$ recursion).
 Part II introduces execution-format comparators (e.g., IEEE-754), a drift ledger, and a decision-invariance theorem that supports commensurate-accuracy claims in floating arithmetic (and that cleanly accounts for any staged/truncated auxiliary drift).
@@
 Fix $D\in\mathbb{N}$ and recall $\mathcal{L}_{D}=D^{-1}\mathbb{Z}[1/2]$ (Definition 3.1).
-At each recursion node (dimension $m$), after centering and dyadic $\ell_{2}$ normalization have produced normalized operands $A^{\prime},B^{\prime}\in\mathcal{L}_{D}^{m\times m}$, we define the node-local coefficient-lattice exponent $\sigma\in\mathbb{Z}_{\geq 0}$ as the smallest integer such that all dot-product terms $A^{\prime}_{ir}B^{\prime}_{rj}$ lie in $D^{-2}2^{-\sigma}\mathbb{Z}$.
+At each recursion node (dimension $m$), after centering and dyadic $\ell_{2}$ normalization have produced normalized operands $A^{\prime},B^{\prime}\in\mathcal{L}_{D}^{m\times m}$, we define the node-local coefficient-lattice exponent $\sigma\in\mathbb{Z}_{\geq 0}$ as the smallest integer such that every scalar product term that can appear in the middle-band $T$ and upper-band $U$ of the local three-band identity (Proposition 5.7) lies in $D^{-2}2^{-\sigma}\mathbb{Z}$.
+(In particular, $\sigma$ must cover the cross-half products $A^{\prime}_{i,r}B^{\prime}_{r+m/2,j}$ that contribute to the upper band $U$ under the canonical $2\times 2$ packing at this node.)
 Write each stored scalar in mantissa/exponent-offset form $x=(w/D)2^{e}$ (Appendix B.1) and denote its exponent offset by $\exp(x):=e$, with the convention $\exp(0):=+\infty$.
-Define for each inner index $r\in\{1,\dots,m\}$:
+Define for each index $r\in\{1,\dots,m\}$:
 $a_{r}$ $:=$ $\min_{1\leq i\leq m}\exp(A^{\prime}_{ir}),$ (2)
-$b_{r}$ $:=$ $\min_{1\leq j\leq m}\exp(B^{\prime}_{rj}),$ (3)
-and let:
-$\mathcal{R}~:=~\{r\in\{1,\dots,m\}:a_{r}<\infty~\text{and}~b_{r}<\infty\}.$ (4)
-If $\mathcal{R}=\emptyset$ (i.e., for every $r$ either $A^{\prime}_{:,r}\equiv 0$ or $B^{\prime}_{r,:}\equiv 0$, so all dot-product terms vanish), then set $\sigma:=0$. Otherwise define:
-$g_{\min}~:=~\min_{r\in\mathcal{R}}(a_{r}+b_{r}),\\ ~~~\sigma:=\max\{0,\,-g_{\min}\}.$ (5)
+$b_{r}$ $:=$ $\min_{1\leq j\leq m}\exp(B^{\prime}_{rj}).$ (3)
+Let:
+$\mathcal{R}_{T}$ $:=$ $\{r\in\{1,\dots,m\}:a_{r}<\infty~\text{and}~b_{r}<\infty\},$
+$\mathcal{R}_{U}$ $:=$ $\emptyset~\text{if}~m=1,~\text{and otherwise}~\{r\in\{1,\dots,m/2\}:a_{r}<\infty~\text{and}~b_{r+m/2}<\infty\}.$
+If $\mathcal{R}_{T}=\emptyset$ and $\mathcal{R}_{U}=\emptyset$ (i.e., all relevant dot-product terms vanish), then set $\sigma:=0$. Otherwise define:
+$g_{T}$ $:=$ $\min_{r\in\mathcal{R}_{T}}(a_{r}+b_{r}),$
+$g_{U}$ $:=$ $\min_{r\in\mathcal{R}_{U}}(a_{r}+b_{r+m/2})~(\text{take}~g_{U}:=+\infty~\text{if}~\mathcal{R}_{U}=\emptyset),$
+$g_{\min}$ $:=$ $\min\{g_{T},~g_{U}\},$
+$\sigma$ $:=$ $\max\{0,~-g_{\min}\}.$
 The scaled coefficient lattice at this node is then:
@@
 With this choice of $\sigma$, every band coefficient $U,T$ in the three-band identity (Proposition 5.7) for products of normalized blocks lies in $\mathcal{L}_{\mathrm{coeff}}(\sigma)$, so the scaled rounding operator is well-defined:
-$\frac{1}{D^{2}2^{\sigma}}\left\lfloor D^{2}2^{\sigma}z\right\rceil~:=~\frac{1}{D^{2}2^{\sigma}}\left\lfloor D^{2}2^{\sigma}z\right\rceil.$ (7)
+$\operatorname{Round}_{\sigma}\!\left(z\right)~:=~\frac{1}{D^{2}2^{\sigma}}\left\lfloor D^{2}2^{\sigma}z\right\rceil.$ (5)
 The scaled sieve extractor is:
-$\mathcal{S}_{\beta,\sigma}(z)~:=~\frac{1}{D^{2}2^{\sigma}}\left\lfloor D^{2}2^{\sigma}z\right\rceil-\beta\frac{1}{D^{2}2^{\sigma}}\left\lfloor D^{2}2^{\sigma}z/\beta\right\rceil$ (8)
+$\mathcal{S}_{\beta,\sigma}(z)~:=~\operatorname{Round}_{\sigma}\!\left(z\right)-\beta\operatorname{Round}_{\sigma}\!\left(z/\beta\right)$ (6)
 (cf., Definition 5.1 and Appendix B.6).
+Remark 3.3 (On “global minima” choices for $\sigma$).
+A safe but overly conservative alternative is to choose $\sigma$ from the decoupled bound $\min_{i,r}\exp(A^{\prime}_{ir})+\min_{r,j}\exp(B^{\prime}_{rj})$, which allows pairing an entry of $A^{\prime}$ and an entry of $B^{\prime}$ that can never meet in any inner-product at this node.
+The definition above is the smallest choice that covers exactly the inner-index pairings that actually occur in the local packing (matched terms for $T$ and cross-half terms for $U$), and avoids inflating $\beta(\sigma)$ (and any staging budget) unnecessarily.
 Definition 3.4 (Digit-Band Staging).
@@
 Lemma 4.1 (Margin-based rounding stability).
-Let $\left\lfloor\cdot\right\rceil:\mathbb{R}\to\mathbb{Z}$ denote rounding to the nearest integer with any fixed deterministic tie-breaking convention.
+Let $\left\lfloor\cdot\right\rceil:\mathbb{R}\to\mathbb{Z}$ denote rounding to the nearest integer with ties-to-even.
 Fix $m\in\mathbb{Z}$ and $\delta\in(0,\frac{1}{2}]$.
@@
 Fix integer base $\beta>1$ and scaling exponent $\sigma\geq 0$. Define the scaled sieve operator:
-$\mathcal{S}_{\beta,\sigma}(z)~:=~\frac{1}{D^{2}2^{\sigma}}\left\lfloor D^{2}2^{\sigma}z\right\rceil-\beta\frac{1}{D^{2}2^{\sigma}}\left\lfloor D^{2}2^{\sigma}z/\beta\right\rceil.$ (15)
+$\mathcal{S}_{\beta,\sigma}(z)~:=~\operatorname{Round}_{\sigma}\!\left(z\right)-\beta\operatorname{Round}_{\sigma}\!\left(z/\beta\right).$ (13)
 Lemma 5.2 (Sieve exactness on scaled lattice).
@@
 $\left|\frac{T}{\beta}+\frac{L}{\beta^{2}}\right|~\leq~\frac{11}{32}+\frac{9}{32^{2}}~<~\frac{1}{2},$ (35)
-so Lemma 5.2 applies and $\mathcal{S}_{\beta}(z)=T$.
+so Lemma 5.2 applies and $\mathcal{S}_{\beta,0}(z)=T$.
 One quadrant packing identity.
@@
     (No square roots needed (via an integer lift).) If $\widetilde{S}(x)>0$ and $\ell:=\lfloor\log_{2}\widetilde{S}(x)\rfloor$ (bitlength minus one), let $K_{0}:=\lfloor\ell/2\rfloor$ and set:
-    $K(x)~=~\begin{cases}K_{0},&\text{if}~\widetilde{S}(x)\leq 2^{2K_{0}},\\ K_{0}+1,&\text{otherwise,}\end{cases}~~k(x)~=~\max\{0,\,K(x)-t\},$ (43)
+    $K(x)~=~\begin{cases}K_{0},&\text{if}~\widetilde{S}(x)\leq 2^{2K_{0}},\\ K_{0}+1,&\text{otherwise,}\end{cases}~~k(x)~=~\max\{0,~K(x)-t\},$ (41)
     where $t:=\max\{0,-\min_{r}e_{r}\}$ is the shift from Definition 6.1 satisfying $S(x)=2^{-2t}\widetilde{S}(x)$.
-Equivalently, $k(x)=\max\{0,\,\lceil\frac{1}{2}\log_{2}\widetilde{S}(x)\rceil-t\}$, with $k(0)=0$.
+Equivalently, $k(x)=\max\{0,~\lceil\frac{1}{2}\log_{2}\widetilde{S}(x)\rceil-t\}$, with $k(0)=0$.
   3. 3.
@@
   3. 3.
-    By Definition 3.1 of $\mathcal{L}_{D}=D^{-1}\mathbb{Z}[1/2]$, multiplying any entry by $2^{\pm k}$ remains in $\mathcal{L}_{D}$ exactly.In the mantissa/exponent-offset representation, this is implemented by updating only the exponent offset by $\pm k$.
+    By Definition 3.1 of $\mathcal{L}_{D}=D^{-1}\mathbb{Z}[1/2]$, multiplying any entry by $2^{\pm k}$ remains in $\mathcal{L}_{D}$ exactly.
+In the mantissa/exponent-offset representation, this is implemented by updating only the exponent offset by $\pm k$.
 Lemma 6.5 (Dimension-free dot products).
@@
 Fix the slack constant $\delta_{0}:=2^{-3}=1/8$. In exact dyadic lattice semantics (Part I), it suffices to enforce the sufficient-base condition
-$D^{2}2^{\sigma}\left(\frac{2}{\beta}+\frac{1}{\beta^{2}}\right)\;\leq\;\frac{1}{2}-\delta_{0},$ (55)
+$D^{2}2^{\sigma}\left(\frac{2}{\beta}+\frac{1}{\beta^{2}}\right)~\leq~\frac{1}{2}-\delta_{0},$ (53)
 where $\sigma$ is the node-local coefficient-lattice exponent (Definition 3.2).
@@
 53 fails, increment $b$ once (doubling $\beta$).
-This rule (and the corresponding exact integer certificate) is spelled out in Appendix  B.5.
+This rule (and the corresponding exact integer certificate) is spelled out in Appendix B.5.
 Also, because $D^{2}2^{\sigma}$ is just a power-of-two shift of the fixed integer $D^{2}$, we have $t=\sigma+\lfloor\log_{2}(D^{2})\rfloor$, so the candidate $b=t+3$ is obtained by a single integer add once $\sigma$ is known.
 Proof.
-This is Theorem  7.1 specialized to the fixed slack $\delta=\delta_{0}$. The dyadic selection rule and the exact integer certificate are proved in Appendix  B.5. ∎
+This is Theorem 7.1 specialized to the fixed slack $\delta=\delta_{0}$. The dyadic selection rule and the exact integer certificate are proved in Appendix B.5. ∎
 Remark 7.3 (Why a growing base is not “cheating” (and what $\sigma$ means)).
 Condition (52) implies that $\beta$ must grow roughly in proportion to $2^{\sigma}$.
-There is no need for a “speed-first” policy: in exact (integral) dyadic mode we fix a small constant slack $\delta_{0}:=2^{-3}=1/8$ (not a tuning knob) and choose the dyadic base $\beta=2^{b}$ deterministically by Corollary 7.2 / Appendix  B.5, with a constant-cost integer certificate that the sufficient-base inequality holds at each node.
+There is no need for a “speed-first” policy: in exact (integral) dyadic mode we fix a small constant slack $\delta_{0}:=2^{-3}=1/8$ (not a tuning knob) and choose the dyadic base $\beta=2^{b}$ deterministically by Corollary 7.2 / Appendix B.5, with a constant-cost integer certificate that the sufficient-base inequality holds at each node.
 Here $\sigma$ is the node-local coefficient-lattice exponent from Definition 3.2, i.e., it is determined by the smallest dyadic lattice step on which the normalized dot-product coefficients at that node actually lie.
@@
 Centering enforces mean constraints (row and/or column means) and is inverted by the bilinear identity (Lemma 8.1).
-Normalization enforces $\ell_{2}$ bounds and is inverted by applying the stored dyadic exponents, i.e., are treat as separate ordered steps: center first, then normalize.
+Normalization enforces $\ell_{2}$ bounds and is inverted by applying the stored dyadic exponents, i.e., are treated as separate ordered steps: center first, then normalize.
   • What propagates.
@@
   1. 1.
-    Dyadic means are exact: Node dimensions are powers of two. Division by $m=2^{t}$ is implemented as an exponent-offset update.
+    Dyadic means are exact: All recursion node dimensions are powers of two.
+If the input dimension $n$ is not a power of two, pad $A,B$ with zeros to $N:=2^{\lceil\log_{2}n\rceil}$ and run KernelMM on the $N\times N$ instance (returning the top-left $n\times n$ block at the end).
+Division by $m=2^{t}$ is implemented as an exponent-offset update.
   2. 2.
@@
 At each node, $\sigma$ MUST be computed from the exponent offsets of the normalized operands $A^{\prime},B^{\prime}$ as in Definition 3.2 (equivalently Eq.
-266).
+289).
 Implementations MUST NOT set $\sigma$ equal to recursion depth (this can over-inflate $\beta$ and destroy performance).
@@
 If it fails at the initial candidate, apply the one-step upgrade rule $b\leftarrow b+1$ (doubling $\beta$) and re-evaluate.
-If it still fails after the one-step upgrade, the implementation MUST take the failure-policy fallback of Appendix  B.5 (call ClassicalMM on the original subproblem and return) and MUST NOT apply the sieve at that node.
-(Equivalently, in staged-precision mode an implementation may signal failure to a retry wrapper and re-attempt with larger staging; see Algorithm  2.)
+If it still fails after the one-step upgrade, the implementation MUST take the failure-policy fallback of Appendix B.5 (call ClassicalMM on the original subproblem and return) and MUST NOT apply the sieve at that node.
+(Equivalently, in staged-precision mode an implementation may signal failure to a retry wrapper and re-attempt with larger staging; see Algorithm 2.)
   4. 4.
+    Checked integer arithmetic (no silent wraparound): All mantissa arithmetic used by centering, norm accumulation, packing/unpacking, and rounding/extraction (including evaluation of the integer base certificate Eq.
+296 in Appendix B.5) MUST be carried out either in a type wide enough to avoid overflow or with explicit overflow checks.
+Silent wraparound is not permitted.
+Any detected overflow (including overflow of exponent offsets) MUST be treated as leaving the declared exact-execution regime and MUST trigger the same failure policy as a failed base certificate (Appendix B.5): do not apply the sieve; instead fall back to ClassicalMM on the original subproblem and return, or signal failure to a retry wrapper (Algorithm 2).
+  5. 5.
     Shifts do not touch mantissas: Multiplication by $2^{\pm t}$ and by $\beta^{\pm 1}$ must be implemented as exponent updates, not as integer multiplication on mantissas.
-  5. 5.
+  6. 6.
     Sieve sites are the only discrete extraction decisions: The only places where rounding/thresholding affects correctness are the sieve/extractor decision sites governed by Lemmas 5.2 and 5.3.
 8.3 Kernel pseudocode
-Input: $A,B\in\mathcal{L}_{D}^{n\times n}$, with $n$ a power of two
-Output: $C=AB$ exactly
+Algorithm 1 KernelMM(A,B)(A,B) (dyadic lattice semantics)
+$A,B\in\mathcal{L}_{D}^{n\times n}$, $n\geq 1$.
+$C=AB$ exactly in dyadic lattice semantics.
+if $n$ is not a power of two then
+  $N\leftarrow 2^{\lceil\log_{2}n\rceil}$
+  Form $\widetilde{A},\widetilde{B}\in\mathcal{L}_{D}^{N\times N}$ by embedding $A,B$ in the top-left $n\times n$ block and padding all other entries with 0.
+[…]
+  return the top-left $n\times n$ block of $\widetilde{C}$.
+end if
 if $n=1$ then
- return ABAB
-(1) Center. Compute the exact column-mean vector $\alpha\in\mathcal{L}_{D}^{n}$ of $A$ and exact row-mean vector $\rho\in\mathcal{L}_{D}^{n}$ of $B$ using dyadic division:
-$\alpha_{j}$ $:=$ $\frac{1}{n}\sum_{i=1}^{n}A_{ij},$
-$\rho_{i}$ $:=$ $\frac{1}{n}\sum_{j=1}^{n}B_{ij}.$
-Set $A_{0}:=A-\mathbf{1}\alpha^{\top}$ and $B_{0}:=B-\rho\mathbf{1}^{\top}$.
-(2) Dyadic $\ell_{2}$ normalize (sqrt-free). For each row $i$ of $A_{0}$, compute $k_{i}:=k(A_{0,i,:})$ via Definitions 6.16.2, and scale:
-$A^{\prime}_{i,:}~:=~2^{-k_{i}}A_{0,i,:}.$
-For each column $j$ of $B_{0}$, compute $\ell_{j}:=k(B_{0,:,j})$ and scale:
-$B^{\prime}_{:,j}~:=~2^{-\ell_{j}}B_{0,:,j}.$
+  return $AB$.
+end if
+(1) Center.
+For $j=1,\dots,n$, set $\alpha_{j}\leftarrow\frac{1}{n}\sum_{i=1}^{n}A_{ij}$ (exact dyadic division).
+For $i=1,\dots,n$, set $\rho_{i}\leftarrow\frac{1}{n}\sum_{j=1}^{n}B_{ij}$ (exact dyadic division).
+$A_{0}\leftarrow A-\mathbf{1}\alpha^{\top}$ and $B_{0}\leftarrow B-\rho\mathbf{1}^{\top}$.
+(2) Dyadic $\ell_{2}$ normalize (sqrt-free).
+For each row $i$ of $A_{0}$, compute $k_{i}\leftarrow k(A_{0,i,:})$ via Definitions 6.16.2, and scale $A^{\prime}_{i,:}\leftarrow 2^{-k_{i}}A_{0,i,:}$.
+For each column $j$ of $B_{0}$, compute $\ell_{j}\leftarrow k(B_{0,:,j})$ and scale $B^{\prime}_{:,j}\leftarrow 2^{-\ell_{j}}B_{0,:,j}$.
 Store the exponents $\{k_{i}\}_{i=1}^{n}$ and $\{\ell_{j}\}_{j=1}^{n}$ for the undo step.
-(2b) Compute the node-local lattice exponent $\sigma$ (no extra passes). Using exponent offsets $\exp(\cdot)$ in the representation $x=(w/D)2^{e}$ (Appendix B.1), compute:
-$a_{r}$ $:=$ $\min_{i}\exp(A^{\prime}_{ir}),~b_{r}:=\min_{j}\exp(B^{\prime}_{rj}),$ (69)
-$\mathcal{R}$ $:=$ $\{\,r:a_{r}<\infty~\text{and}~b_{r}<\infty\,\}.$ (70)
-If $\mathcal{R}=\emptyset$, set $\sigma:=0$; otherwise set $g_{\min}:=\min_{r\in\mathcal{R}}(a_{r}+b_{r})$ and $\sigma:=\max\{0,-g_{\min}\}$.
-This is exactly Definition 3.2 / Eq.
-266, and can be accumulated during Step (2).
-(3) Pack, recurse, and extract quadrants. Partition $A^{\prime},B^{\prime}$ into $2\times 2$ blocks of size $(n/2)\times(n/2)$.
-/* Certified base selection (no tuning knobs; at most one upgrade). */
-Compute the candidate exponent $b_{0}$ (i.e., $\beta_{0}=2^{b_{0}}$) from Eq. 270. Set $b\leftarrow b_{0}$ and $\beta\leftarrow 2^{b}$
-Evaluate the exact integer certificate Eq. 271 (and, in a reference implementation, assert the normalization bounds from Step (2))
+(2b) Compute the node-local lattice exponent $\sigma$ (no extra passes).
+Using exponent offsets $\exp(\cdot)$ in the representation $x=(w/D)2^{e}$ (Appendix B.1), compute for each $r=1,\dots,n$:
+  $a_{r}\leftarrow\min_{i}\exp(A^{\prime}_{ir})$ and $b_{r}\leftarrow\min_{j}\exp(B^{\prime}_{rj})$.
+Define the index sets (cf. Definition 3.2):
+  $\mathcal{R}_{T}\leftarrow\{r\in\{1,\dots,n\}:a_{r}<\infty~\text{and}~b_{r}<\infty\}$,
+  $\mathcal{R}_{U}\leftarrow\{r\in\{1,\dots,n/2\}:a_{r}<\infty~\text{and}~b_{r+n/2}<\infty\}$.
+if $\mathcal{R}_{T}=\emptyset$ and $\mathcal{R}_{U}=\emptyset$ then
+  $\sigma\leftarrow 0$.
+else
+  $g_{T}\leftarrow\min_{r\in\mathcal{R}_{T}}(a_{r}+b_{r})$.
+  $g_{U}\leftarrow\min_{r\in\mathcal{R}_{U}}(a_{r}+b_{r+n/2})$.
+  $g_{\min}\leftarrow\min\{g_{T},g_{U}\}$ and $\sigma\leftarrow\max\{0,-g_{\min}\}$.
+end if
+This is exactly Definition 3.2 / Eq. 289 and can be accumulated during Step (2).
+(3) Pack, recurse, and extract quadrants (folded across the $j$-index).
+Partition $A^{\prime},B^{\prime}$ into $2\times 2$ blocks of size $(n/2)\times(n/2)$.
+$\triangleright$ Certified base selection (unchanged; still certifies the inner $\beta$-sieve).
+Compute the candidate exponent $b_{0}$ (i.e., $\beta_{0}=2^{b_{0}}$) from Eq. 295. Set $b\leftarrow b_{0}$ and $\beta\leftarrow 2^{b}$.
+Evaluate the exact integer certificate Eq. 296 (and, in a reference implementation, assert the normalization bounds from Step (2)).
 if the certificate fails at $b=b_{0}$ then
- $b\leftarrow b_{0}+1$$\beta\leftarrow 2^{b}$
- // one-step upgrade, Eq. 272
-   Re-pack using the upgraded $\beta$ and re-evaluate the certificate (in exact integer arithmetic this second check must succeed; if it does not, treat it as overflow/truncation)
+  $b\leftarrow b_{0}+1$; $\beta\leftarrow 2^{b}$ $\triangleright$ one-step upgrade, Eq. 297
+  Re-pack using the upgraded $\beta$ and re-evaluate the certificate (in exact integer arithmetic this second check must succeed; if it does not, treat it as overflow/truncation).
   if the certificate still fails then
-    return ClassicalMM(A,B)(A,B)
-    // exact fallback; Appendix B.5
-foreach $(i,j)\in\{1,2\}^{2}$ do
-   Form packed blocks (Definition 5.6):
-$X_{ij}$ $:=$ $A^{\prime}_{i1}+\beta^{-1}A^{\prime}_{i2},$
-$Y_{ij}$ $:=$ $B^{\prime}_{1j}+\beta B^{\prime}_{2j}.$
-$Z_{ij}\leftarrow\textsc{KernelMM}(X_{ij},Y_{ij})$ $T_{ij}\leftarrow\mathcal{S}_{\beta,\sigma}(Z_{ij})$
- // Entrywise sieve (Definition 5.1)
+   return ClassicalMM(A,B)(A,B) $\triangleright$ exact fallback; Appendix B.5
+  end if
+end if
+$\triangleright$ Folded column packing: introduce a second dyadic base $\gamma$ to carry both block-columns at once.
+Set $\gamma\leftarrow\beta^{2}$. $\triangleright$ any certified $\gamma\gg\beta$ works; $\beta^{2}$ is a simple constant-overhead choice
+$\triangleright$ Single folded right embedding (both block-columns packed into $\{0,1\}$ in the $\gamma$-grade).
+$Y\leftarrow\big(B^{\prime}_{11}+\beta B^{\prime}_{21}\big)~+~\gamma\big(B^{\prime}_{12}+\beta B^{\prime}_{22}\big)$.
+for all $i\in\{1,2\}$ do
+$\triangleright$ Left embedding (one per block-row), same as before but now independent of $j$.
+  $X_{i}\leftarrow A^{\prime}_{i1}+\beta^{-1}A^{\prime}_{i2}$.
+  $Z_{i}\leftarrow\textsc{KernelMM}(X_{i},Y)$. $\triangleright$ two recursive calls total
+$\triangleright$ Project to the two $\gamma$-coefficients: $Z_{i}=Z_{i}^{(0)}+\gamma Z_{i}^{(1)}$ (no $\gamma^{-1}$ band).
+  $Z_{i}^{(0)}\leftarrow\mathcal{S}_{\gamma,\sigma}(Z_{i})$.
+  $Z_{i}^{(1)}\leftarrow(Z_{i}-Z_{i}^{(0)})/\gamma$. $\triangleright$ exact exponent shift
+$\triangleright$ Inner extraction (the original $\beta$-sieve) applied to each $\gamma$-bucket.
+  $T_{i1}\leftarrow\mathcal{S}_{\beta,\sigma}(Z_{i}^{(0)})$.
+  $T_{i2}\leftarrow\mathcal{S}_{\beta,\sigma}(Z_{i}^{(1)})$.
+end for
 Assemble $T$ from the four blocks $T_{ij}$.
-(4) Undo normalization and un-center. Undo the dyadic scalings to recover $A_{0}B_{0}$ from $T$:
-$(A_{0}B_{0})_{ij}~=~2^{k_{i}+\ell_{j}}T_{ij}~~~\text{(entrywise)},$
-and then recover $C=AB$ using the bilinear centering identity (Lemma 8.1):
-$C~:=~A_{0}B_{0}+\mathbf{1}(\alpha^{\top}B_{0})+(A_{0}\rho)\mathbf{1}^{\top}+(\alpha^{\top}\rho)\mathbf{1}\mathbf{1}^{\top}.$
-return $C$
-Algorithm 1 KernelMM(A,B)(A,B) (dyadic lattice semantics)
+(4) Undo normalization and un-center.
+Undo the dyadic scalings to recover $A_{0}B_{0}$ from $T$ entrywise:
+  $(A_{0}B_{0})_{ij}\leftarrow 2^{k_{i}+\ell_{j}}T_{ij}$.
+Recover $C=AB$ using the bilinear centering identity (Lemma 8.1):
+$C\leftarrow A_{0}B_{0}+\mathbf{1}(\alpha^{\top}B_{0})+(A_{0}\rho)\mathbf{1}^{\top}+(\alpha^{\top}\rho)\mathbf{1}\mathbf{1}^{\top}$.
+return $C$.
 Remark 8.4 (Order of operations).
@@
 Therefore, by the coefficient-lattice alignment hypothesis and Lemma 5.2, extraction applies: entrywise and gives:
-    $\mathcal{S}_{\beta}(Z_{ij})~=~\mathcal{S}_{\beta}(X_{ij}Y_{ij})=T_{ij},$ (73)
+    $\mathcal{S}_{\beta,\sigma}(Z_{ij})~=~\mathcal{S}_{\beta,\sigma}(X_{ij}Y_{ij})=T_{ij},$ (69)
     so the extracted blocks assembled by the algorithm equal the exact quadrants of $A^{\prime}B^{\prime}$, so the assembled matrix $T$ equals $A^{\prime}B^{\prime}$ exactly.
@@
 Remark 9.1 (Certified regime and staged retry).
-In exact dyadic lattice semantics, Appendix  B.5 proves that the certified base-selection rule (Eq.
+In exact dyadic lattice semantics, Appendix B.5 proves that the certified base-selection rule (Eq.
 296 with the one-step upgrade Eq.
-272) succeeds after at most one upgrade at every node, provided coefficient-lattice alignment is preserved (i.e., no mantissa bits are dropped below the local $\mathcal{L}_{\mathrm{coeff}}(\sigma)$ resolution).
-Therefore, within the declared exact-execution regime, the ClassicalMM branch in Algorithm  1 is unreachable and serves only as a correct-on-return guard for overflow/truncation (or implementation bugs).
+297) succeeds after at most one upgrade at every node, provided coefficient-lattice alignment is preserved (i.e., no mantissa bits are dropped below the local $\mathcal{L}_{\mathrm{coeff}}(\sigma)$ resolution).
+Therefore, within the declared exact-execution regime, the ClassicalMM branch in Algorithm 1 is unreachable and serves only as a correct-on-return guard for overflow/truncation (or implementation bugs).
 On fixed-width hardware, if a run ever leaves this regime, one may either take the conservative ClassicalMM fallback, or increase the staged-precision budget / word width and retry.
-Algorithm  2 formalizes a simple retry-with-more-staging wrapper, and Theorem  9.2 shows that it preserves the soft-quadratic bit/word complexity when the required precision $b(n)$ is polylogarithmic.
+Algorithm 2 formalizes a simple retry-with-more-staging wrapper, and Theorem 9.2 shows that it preserves the soft-quadratic bit/word complexity when the required precision $b(n)$ is polylogarithmic.
 9.1 Retry with more staging (optional)
-The ClassicalMM branch in Algorithm  1 is a correct-on-return backstop for implementations that leave the declared exact-execution regime (Appendix  B.5).
+The ClassicalMM branch in Algorithm 1 is a correct-on-return backstop for implementations that leave the declared exact-execution regime (Appendix B.5).
 In staged-precision fixed-word mode, a more efficient response is often to increase the staged mantissa budget and rerun the computation.
@@
 Attempt primitive.
-Let KernelMM_Attempt(A,B;K)(A,B;K) denote Algorithm  1 executed in digit-band staged mode with $K$ words per scalar, except that the failure-policy fallback of Appendix  B.5 is replaced by returning a failure flag and aborting the attempt (i.e., no extracted value computed under a failed certificate is ever returned).
-Input: $A,B\in\mathcal{L}_{D}^{n\times n}$ with $n$ a power of two; initial staging $K_{0}\geq 1$; maximum staging budget $K_{\max}$
-Output: $C=AB$ exactly (or exact fallback if the staging limit is exhausted)
-$K\leftarrow K_{0}$
+Let KernelMM_Attempt(A,B;K)(A,B;K) denote Algorithm 1 executed in digit-band staged mode with $K$ words per scalar, except that the failure-policy fallback of Appendix B.5 is replaced by returning a failure flag and aborting the attempt (i.e., no extracted value computed under a failed certificate is ever returned).
+[…]
+$A,B\in\mathcal{L}_{D}^{n\times n}$ with $n$ a power of two; initial staging $K_{0}\geq 1$; maximum staging budget $K_{\max}$.
+$C=AB$ exactly (or exact fallback if the staging limit is exhausted).
+$K\leftarrow K_{0}$.
 while $K\leq K_{\max}$ do
- $(C,\mathrm{ok})\leftarrow\textsc{KernelMM\_Attempt}(A,B;K)$
+  $(C,\mathrm{ok})\leftarrow\textsc{KernelMM\_Attempt}(A,B;K)$.
   if $\mathrm{ok}$ then
-    return $C$
- $K\leftarrow 2K$
-return ClassicalMM(A,B)(A,B)
-// correct-on-return backstop
-[…]
+   return $C$.
+  end if
+  $K\leftarrow 2K$.
+end while
+return ClassicalMM(A,B)(A,B) $\triangleright$ correct-on-return backstop.
 Theorem 9.2 (Retry wrapper: correctness and complexity).
 Fix a base word width $w$ and consider digit-band (slice) staging with $K$ words per scalar.
-Assume that for a full run of Algorithm  1 under $K$-word staging in the certified (no-failure) regime, the fixed-word cost satisfies $B(n,K)=\Theta(n^{2}\log n\cdot K)$ (each scalar primitive costs $\Theta(K)$ word operations).
+Assume that for a full run of Algorithm 1 under $K$-word staging in the certified (no-failure) regime, the fixed-word cost satisfies $B(n,K)=\Theta(n^{2}\log n\cdot K)$ (each scalar primitive costs $\Theta(K)$ word operations).
 […]
@@
 Correctness.
-On the successful iteration, no node triggers the failure policy, so the attempted execution follows exactly the same computation path as Algorithm  1 and satisfies the hypotheses of Theorem  8.5, so it returns $C=AB$ exactly.
-If instead the wrapper exhausts the staging limit ($K>K_{\max}$), it returns ClassicalMM(A,B)(A,B), which is exact by Algorithm  3.
+On the successful iteration, no node triggers the failure policy, so the attempted execution follows exactly the same computation path as Algorithm 1 and satisfies the hypotheses of Theorem 8.5, so it returns $C=AB$ exactly.
+If instead the wrapper exhausts the staging limit ($K>K_{\max}$), it returns ClassicalMM(A,B)(A,B), which is exact by Algorithm 4.
 Cost.
 For each failed attempt with budget $K_{i}$, the wrapper performs at most the work of a full staged run at the same $K_{i}$. Therefore the total cost is bounded by:
-$\sum_{i=0}^{s}B(n,K_{i})~=~\Theta(n^{2}\log n)\sum_{i=0}^{s}K_{i}~=~\Theta(n^{2}\log n)\,(2K_{s}-K_{0})~=~O(n^{2}\log n\cdot K_{s}).$ (74)
+$\sum_{i=0}^{s}B(n,K_{i})$ $=$ $\Theta(n^{2}\log n)\sum_{i=0}^{s}K_{i}$ (70)
+$=$ $\Theta(n^{2}\log n)(2K_{s}-K_{0})$ (71)
+$=$ $O(n^{2}\log n\cdot K_{s}).$ (72)
 Because $K_{s}<2K^{\star}$, this is $O(n^{2}\log n\cdot K^{\star})=O(B(n,K^{\star}))$, and likewise $O(B(n,2K^{\star}))$.
@@
   5. 5.
-    applying the sieve entrywise to extract the middle bands,
+    Applying the sieve entrywise to extract the middle bands, and
   6. 6.
-    and performing the rank-one centering correction.
+    Performing the rank-one centering correction.
 Each of these steps takes $O(m^{2})$ scalar primitive operations (add/subtract/multiply by constants, and rounding calls).
@@
 Theorem 9.4 (Arithmetic complexity).
-Let $T(n)$ denote the arithmetic time of KernelMM on $n\times n$ inputs, measured as the number of scalar ring operations (additions, multiplications by fixed constants such as powers of two, and nearest-integer rounding calls), assuming $n$ is a power of two and that Appendix  B.5’s certificate-and-upgrade rule succeeds at every recursion node (so the algorithm never takes the ClassicalMM fallback), then:
+Let $T(n)$ denote the arithmetic time of KernelMM on $n\times n$ inputs, measured as the number of scalar ring operations (additions, multiplications by fixed constants such as powers of two, and nearest-integer rounding calls), assuming $n$ is a power of two and that Appendix B.5’s certificate-and-upgrade rule succeeds at every recursion node (so the algorithm never takes the ClassicalMM fallback), then:
 $T(n)~=~4T(n/2)+O(n^{2}),$ (73)
@@
 Proof.
-We now consider a single call of KernelMM on size $n$ under the assumption that the certificate-and-upgrade rule of Appendix  B.5 succeeds at this node and recursively at all descendants (so no fallback occurs).
+We now consider a single call of KernelMM on size $n$ under the assumption that the certificate-and-upgrade rule of Appendix B.5 succeeds at this node and recursively at all descendants (so no fallback occurs).
 Recursive calls:
@@
 Fix a word width $w$ for mantissa storage, and assume exponent offsets and loop indices fit in fixed-width machine integers.
-Assume additionally that Appendix  B.5’s certificate-and-upgrade rule succeeds at every recursion node (so the algorithm never takes the ClassicalMM fallback).
+Assume additionally that Appendix B.5’s certificate-and-upgrade rule succeeds at every recursion node (so the algorithm never takes the ClassicalMM fallback).
 Let $b(n)$ denote the required effective mantissa precision (in bits) needed to represent packed temporaries and preserve the extractor slack $\delta$ at all decision sites for dimension $n$ under the chosen execution discipline.
@@
 Thus the middle band $T:=f_{0}^{\prime}\ast g_{0}^{\prime}+f_{1}^{\prime}\ast g_{1}^{\prime}$ and the tail band $L:=f_{1}^{\prime}\ast g_{0}^{\prime}$ satisfy $|T|\leq 2$ and $|L|\leq 1$ coefficientwise.
-Therefore, choosing $\beta$ so that the master gap inequality holds coefficientwise (Lemma 5.3) the master gap inequality holds coefficientwise (Lemma 5.3) gives exact middle-band recovery coefficientwise by Lemma 5.2.
-For example, any $\beta$ satisfying:
+Therefore, choosing $\beta$ so that the master gap inequality holds coefficientwise (Lemma 5.3) gives exact middle-band recovery coefficientwise by Lemma 5.2. For example, any $\beta$ satisfying:
 $D^{2}\left(\frac{|L|}{\beta}+\frac{|T|}{\beta}+\frac{|L|}{\beta^{2}}\right)\leq\frac{1}{2}-\delta$ (88)
@@
 10 Prior related work and what is new here
-The core algebraic trick behind our quadrant packing is not new: embed multiple quantities into a single multiplication by introducing a large parameter (a base $\beta$, or a formal variable $\lambda$), and then decodes a desired coefficient/band by an extractor (rounding, digit extraction, interpolation, CRT, etc.).
+The core algebraic trick behind our quadrant packing is not new: embed multiple quantities into a single multiplication by introducing a large parameter (a base $\beta$, or a formal variable $\lambda$), and then decode a desired coefficient/band by an extractor (rounding, digit extraction, interpolation, CRT, etc.).
 Variants of this idea appear across several communities.
@@
 The GPR analysis is unchanged, but only $\Lambda_{m}$ improves.
-Lemma 13.9 (Sieve stability under additive drift).
-Fix $\beta\geq 2$ and consider a packed scalar:
-$z^{\star}~=~\beta U+T+L/\beta$ (123)
-with $U,T\in D^{-2}\mathbb{Z}$. Let $\delta\in(0,1/2]$, and suppose the ideal master gap holds with slack $\delta$, i.e.:
-$\lvert D^{2}L/\beta\rvert+\lvert D^{2}T/\beta\rvert+\lvert D^{2}L/\beta^{2}\rvert~\leq~\frac{1}{2}-\delta.$ (124)
-Let the implemented scalar be $z=z^{\star}+\eta$ with additive drift $\eta\in\mathbb{R}$, then if:
-$\lvert D^{2}\eta\rvert~\leq~\delta,$ (125)
-the sieve extraction is invariant, i.e., $\mathcal{S}_{\beta}(z)=\mathcal{S}_{\beta}(z^{\star})=T$.
+Lemma 13.9 (Scaled sieve stability under additive drift).
+Fix $\beta\geq 2$ and let $\sigma\geq 0$ be the node-local coefficient-lattice exponent (Definition 3.2). Consider a packed scalar:
+$z^{\star}~=~\beta U+T+L/\beta,$ (121)
+then with $L\in\mathbb{R}$:
+$U,T\in\mathcal{L}_{\mathrm{coeff}}(\sigma)~=~D^{-2}2^{-\sigma}\mathbb{Z}.$ (122)
+Let $\delta\in(0,1/2]$, and suppose the ideal scaled master gap holds with slack $\delta$, i.e.:
+$\lvert D^{2}2^{\sigma}L/\beta\rvert+\lvert D^{2}2^{\sigma}T/\beta\rvert+\lvert D^{2}2^{\sigma}L/\beta^{2}\rvert~\leq~\frac{1}{2}-\delta.$ (123)
+Let the implemented scalar be $z=z^{\star}+\eta$ with additive drift $\eta\in\mathbb{R}$. If:
+$\lvert D^{2}2^{\sigma}\eta\rvert~\leq~\delta,$ (124)
+then the scaled sieve extraction is invariant:
+$\mathcal{S}_{\beta,\sigma}(z)=\mathcal{S}_{\beta,\sigma}(z^{\star})=T.$ (125)
 Proof.
-The sieve uses only the two rounded quantities $\left\lfloor D^{2}z\right\rceil$ and $\left\lfloor D^{2}z/\beta\right\rceil$.
-Under the ideal slack $\delta$, both rounding arguments are at distance at least $\delta$ from their respective half-integer boundaries.
-The additive perturbation $\eta$ shifts each rounded argument by at most $\delta$ in the lifted units, so Lemma 12.10 (or Lemma 4.1 applied to the lifted arguments) implies the rounding outcomes do not change.
+Let $M:=D^{2}2^{\sigma}$ be the lifted scaling at this decision site.
+The scaled sieve $\mathcal{S}_{\beta,\sigma}$ depends only on the two nearest-integer roundings $\left\lfloor Mz\right\rceil$ and $\left\lfloor Mz/\beta\right\rceil$.
+Under the ideal slack $\delta$, both lifted rounding arguments are at distance at least $\delta$ from their respective half-integer boundaries.
+The additive perturbation $\eta$ shifts each lifted argument by at most $\lvert M\eta\rvert\leq\delta$, so Lemma 12.10 (or Lemma 4.1 applied to the lifted arguments) implies the rounding outcomes do not change.
 Therefore the sieve output does not change.
@@
   3. 3.
-    Recovering middle-band quantities via sieve-style extraction decisions (the same discrete decisions as in Part I), but with all primitive operations executed in format $\mathsf{F}$.
+    Recovering middle-band quantities via the scaled sieve extractor $\mathcal{S}_{\beta,\sigma}$ (Definition 5.1), where the node-local lattice exponent $\sigma$ is computed from the (float) normalized operands as in Definition 3.2; all primitive operations are executed in format $\mathsf{F}$.
 Assume the following hypotheses hold at every recursion node and for every scalar extraction site:
@@
     Ledger-bounded drift: The total drift between the implemented and ideal decision statistics is bounded by $\varepsilon_{\mathrm{tot}}\leq\delta$ (Definition 12.9).
+For scaled-sieve sites, the relevant decision statistics are the lifted rounding arguments MzMz and $Mz/\beta$ with $M=D^{2}2^{\sigma}$ (Lemma 13.9).
   4. 4.
@@
 (Offset-bucket invariant) At each recursion node, each block $X$ is represented as:
-$X=\Delta_{X}\mathbf{1}\mathbf{1}^{\top}+R_{X},$ (179)
+$X~=~\Delta_{X}\mathbf{1}\mathbf{1}^{\top}+R_{X},$ (179)
 where $\Delta_{X}$ is a block-dependent scalar offset and $R_{X}$ is a residual matrix whose finite entries lie in a bounded window (a fixed bucket range).
@@
 If Equation 189 holds, then this is at most $1/2-\delta$, so Lemma 5.3 applies entrywise.
-Therefore the two sieve gap inequalities hold and Lemma 5.2 gives $\mathcal{S}_{\beta}(Z_{\ell})=T_{\ell}$ exactly for that entry.
+Therefore the two sieve gap inequalities hold and Lemma 5.2 gives $\mathcal{S}_{\beta,\sigma_{\ell}}(Z_{\ell})=T_{\ell}$ exactly for that entry.
 Because the argument is uniform in $\ell$ and in the entry, extraction is exact at every depth.
@@
 Remark 21.9 (Examples of extractors).
-  • Rounding-gap extractor (Part I): Here $\mathsf{Packed}=\mathbb{R}$ (or a coefficient lattice), $\mathrm{Extract}=\mathcal{S}_{\beta}$, and $\mathsf{Cond}$ is the master gap inequality ensuring band separation.
+  • Rounding-gap extractor (Part I): Here $\mathsf{Packed}=\mathbb{R}$ (or a coefficient lattice), $\mathrm{Extract}=\mathcal{S}_{\beta,\sigma}$ (with the node-local lattice exponent $\sigma$), and $\mathsf{Cond}$ is the scaled master gap inequality ensuring band separation.
   • CRT extractor: Here $\mathsf{Packed}$ is a tuple of residues $(Z\bmod p_{1},\dots,Z\bmod p_{r})$ and $\mathrm{Extract}$ is CRT reconstruction.
@@
 ∎
+21.5 C) Worked example: a folded $(\beta,\gamma)$ packing for $2\times 2$ block matrix multiplication
+This subsection gives a “template package” for a folded multigraded packing of the $2\times 2$ block recursion for matrix multiplication (or any bilinear kernel).
+The point of the fold is that a single packed kernel evaluation returns two target coefficients (via a constant-size multi-degree band), reducing the node fan-out from four packed products (Algorithm refalg:KernelMM, Step (3)) to two.
+Definition 21.11 (Folded multigraded embedding, target band, and coefficient-projection extractor).
+Let $M,N$ be abelian groups (or modules) and let $\mathcal{B}:M\times M\to N$ be a bilinear kernel (Definition 18.1).
+For blocked matrix multiplication on $(m\times m)$ blocks, one may take $M=N=\mathbb{R}^{m\times m}$ and $\mathcal{B}(X,Y)=XY$.
+Fix a node of a $2\times 2$ block recursion, and write the node-local operands as:
+$A$ $=$ $\begin{pmatrix}A_{11}&A_{12}\\ A_{21}&A_{22}\end{pmatrix}$, (252)
+$B$ $=$ $\begin{pmatrix}B_{11}&B_{12}\\ B_{21}&B_{22}\end{pmatrix}$, (253)
+where each block lies in $M$. The desired block outputs are:
+$C_{ij}~=~\mathcal{B}(A_{i1},B_{1j})+\mathcal{B}(A_{i2},B_{2j})\in N,~~~i,j\in\{1,2\}.$ (254)
+The setting is the bivariate graded extension $M[\beta^{\pm 1},\gamma]$ and $N[\beta^{\pm 1},\gamma]$ (a $t=2$ instance of Definition 21.5 with bases $(\beta_{1},\beta_{2})=(\beta,\gamma)$).
+Folded embeddings.
+Define two left embeddings (one per block-row), each using the $\beta$-window $\{0,-1\}$:
+$X_{i}(\beta,\gamma)~:=~A_{i1}+\beta^{-1}A_{i2}\in M[\beta^{\pm 1},\gamma],~~~i\in\{1,2\}.$ (255)
+Define a single folded right embedding that packs both block-columns into the $\gamma$-window $\{0,1\}$:
+$Y(\beta,\gamma)~:=~(B_{11}+\beta B_{21})+\gamma(B_{12}+\beta B_{22})\in M[\beta^{\pm 1},\gamma].$ (256)
+Define the two packed kernel outputs:
+$Z_{i}(\beta,\gamma)~:=~\mathcal{B}\!\left(X_{i}(\beta,\gamma),Y(\beta,\gamma)\right)\in N[\beta^{\pm 1},\gamma],~~~i\in\{1,2\}.$ (257)
+Target band and coefficient projection extractor.
+Let the target multi-degree band be:
+$E_{0}~:=~\{(0,0),(0,1)\}\subset\mathbb{Z}^{2},$ (258)
+corresponding to degrees $\beta^{0}\gamma^{0}$ and $\beta^{0}\gamma^{1}$. For a packed value:
+$Z(\beta,\gamma)~=~\sum_{p,q}\beta^{p}\gamma^{q}Z^{[(p,q)]},$ (259)
+write $Z^{[(p,q)]}$ for the coefficient of $\beta^{p}\gamma^{q}$. Define the extractor (Definition 21.8) by exact coefficient projection:
+$\mathrm{Extract}(Z)~:=~\left(Z^{[(0,0)]},~Z^{[(0,1)]}\right)\in N\times N.$ (260)
+Lemma 21.12 (Folded coefficient identity (two outputs per packed product)).
+For each $i\in\{1,2\}$, the target-band coefficients of $Z_{i}=\mathcal{B}(X_{i},Y)$ satisfy:
+$Z_{i}^{[(0,0)]}$ $=$ $\mathcal{B}(A_{i1},B_{11})+\mathcal{B}(A_{i2},B_{21})~=~C_{i1},$ (261)
+$Z_{i}^{[(0,1)]}$ $=$ $\mathcal{B}(A_{i1},B_{12})+\mathcal{B}(A_{i2},B_{22})~=~C_{i2},$ (262)
+then $\mathrm{Extract}(Z_{i})=(C_{i1},C_{i2})$.
+Proof.
+Expand the embeddings by explicit $(\beta,\gamma)$ degrees:
+$X_{i}$ $=$ $\beta^{0}\gamma^{0}A_{i1}+\beta^{-1}\gamma^{0}A_{i2},$ (263)
+$Y$ $=$ $\beta^{0}\gamma^{0}B_{11}+\beta^{1}\gamma^{0}B_{21}+\beta^{0}\gamma^{1}B_{12}+\beta^{1}\gamma^{1}B_{22}.$ (264)
+By bilinearity and the convolution identity for graded bilinear embeddings (Lemma 20.1) applied coefficientwise in the bivariate grading, the coefficient at degree $(p,q)$ is the sum of $\mathcal{B}(\cdot,\cdot)$ over pairs of input degrees adding to $(p,q)$.
+For $(p,q)=(0,0)$, the contributing pairs are $(0,0)+(0,0)$ and (1,0)+(1,0)(-1,0)+(1,0), giving $\mathcal{B}(A_{i1},B_{11})+\mathcal{B}(A_{i2},B_{21})$.
+For $(p,q)=(0,1)$, the contributing pairs are $(0,0)+(0,1)$ and (1,0)+(1,1)(-1,0)+(1,1), giving $\mathcal{B}(A_{i1},B_{12})+\mathcal{B}(A_{i2},B_{22})$. ∎
+Algorithm 3 FoldedGPR(A,B)(A,B) (folded $(\beta,\gamma)$ node)
+Node-local operands $A,B$ for a bilinear kernel $\mathcal{B}:M\times M\to N$; recursion depth $\ell$ implicit.
+$C=\mathcal{B}(A,B)$ in $N$.
+(1) Recursive invariant amplification.
+[…]
+if base case then
+[…]
+end if
+(2) Folded multigraded embedding.
+Partition $\widehat{A},\widehat{B}$ into $2\times 2$ blocks $(\widehat{A}_{ij})$, $(\widehat{B}_{ij})$.
+for $i\leftarrow 1$ to 22 do
+  $X_{i}\leftarrow\widehat{A}_{i1}+\beta^{-1}\widehat{A}_{i2}$.
+end for
+$Y\leftarrow(\widehat{B}_{11}+\beta\widehat{B}_{21})+\gamma(\widehat{B}_{12}+\beta\widehat{B}_{22})$.
+(3) Two recursive kernel calls and multi-band extraction.
+for $i\leftarrow 1$ to 22 do
+  $Z_{i}\leftarrow\textsc{FoldedGPR}(X_{i},Y)$.
+  $(C_{i1},C_{i2})\leftarrow\mathrm{Extract}(Z_{i})=(Z_{i}^{[(0,0)]},Z_{i}^{[(0,1)]})$.
+end for
+Assemble $\widehat{C}\leftarrow\begin{pmatrix}C_{11}&C_{12}\\ C_{21}&C_{22}\end{pmatrix}$.
+Undo the projection on $\widehat{C}$ and return $C$.
+Proposition 21.13 (BWBM/master-condition view, extractor correctness, and fan-out reduction).
+Assume the node-local invariant amplification enforces the usual depth-$\ell$ classical bound $S_{\ell}$ for atomic kernel evaluations (Definition 18.2).
+Then the folded construction in Definition 21.11 is GPR-admissible for the target band $E_{0}=\{(0,0),(0,1)\}$ (Definition 21.6) with a depth-independent constant, and the coefficient-projection extractor is correct.
+Moreover, the folded node replaces the four packed products of the unfurled $2\times 2$ quadrant rule by two packed products per node.
+Proof.
+(A) BWBM certificate in the $\beta$-grade (slice-by-$\gamma$ view).
+Write $Y(\beta,\gamma)=Y_{0}(\beta)+\gamma Y_{1}(\beta)$ where:
+$Y_{0}(\beta)$ $=$ $B_{11}+\beta B_{21},$ (265)
+$Y_{1}(\beta)$ $=$ $B_{12}+\beta B_{22}.$ (266)
+For each fixed $\gamma$-slice $q\in\{0,1\}$, the pair:
+$X_{i}(\beta)$ $=$ $A_{i1}+\beta^{-1}A_{i2},$ (267)
+$Y_{q}(\beta)$ $=$ $B_{1,q+1}+\beta B_{2,q+1}$ (268)
+has exponent windows $D=\{0,-1\}$ and $D^{\prime}=\{0,1\}$ (width $\Delta=1$), no mixing ($\Lambda=1$), and bounded convolution width: for the target $\beta$-degree $e_{0}=0$ there are exactly two contributing pairs $(0,0)$ and (1,1)(-1,1) so $k_{T}=2$, while for the lower degree $e_{0}-1=-1$ the only contributing pair is (1,0)(-1,0) so $k_{L}=1$.
+Thus BWBM (Definition 20.2) holds (per slice) with:
+$(\Delta,\Lambda,k_{T},k_{L})~=~(1,1,2,1),$ (269)
+and Lemma 20.3 yields the induced (univariate) master constants:
+$C_{T}$ $=$ $k_{T}\Lambda^{2}=2,$ (270)
+$C_{L}$ $=$ $k_{L}\Lambda^{2}=1.$ (271)
+(B) Multigraded master condition for the folded target band $E_{0}$.
+By Lemma 21.12, each target coefficient $Z_{i}^{[(0,0)]}$ and $Z_{i}^{[(0,1)]}$ is a sum of exactly two atomic bilinear evaluations at depth $\ell$. Under the depth-$\ell$ classical bound, this gives:
+$\|Z_{i}^{[(0,0)]}\|_{N}$ $\leq$ $2S_{\ell},$ (272)
+$\|Z_{i}^{[(0,1)]}\|_{N}$ $\leq$ $2S_{\ell}.$ (273)
+Therefore, the entire target band $E_{0}$ obeys a depth-uniform multigraded master condition (Definition 21.6) with a constant factor $C=2$ under max-aggregation across degrees, and $C=4$ under sum-aggregation across degrees.
+(C) Extractor correctness.
+Here $\mathrm{Extract}$ is literal coefficient projection onto distinct degrees in the formal graded ring, so its sufficient condition (Definition 21.8) is purely representational (“degrees do not collide”) and holds identically in the symbolic model.
+Under any numeric realization (e.g., mixed-radix in one base), the depth-uniform master bound from (B) is exactly the hypothesis needed to choose global packing parameters so that the sufficient condition holds, and Theorem 21.10 then gives correctness at every depth.
+(D) Why the fold gives two children.
+In the unfurled quadrant rule used in Algorithm 1 (Step (3)), the packed pairs $(X_{ij},Y_{ij})$ are formed for all $(i,j)\in\{1,2\}^{2}$, requiring four recursive kernel calls to recover the four blocks $C_{ij}$.
+In the folded scheme, a single packed product $Z_{i}=\mathcal{B}(X_{i},Y)$ returns two target coefficients, $\beta^{0}\gamma^{0}$ and $\beta^{0}\gamma^{1}$, which are exactly $(C_{i1},C_{i2})$ by Lemma 21.12.
+Therefore only two recursive kernel calls (for $i=1,2$) are needed per node.
+∎
+Remark 21.14 (Finite-window model (no formal symbols)).
+The same construction can be expressed without explicit $\beta,\gamma$ symbols by working over the constant support window:
+$W~=~\{(-1,0),(0,0),(1,0),(-1,1),(0,1),(1,1)\}$ (274)
+and representing packed values as coefficient tables $\{Z^{[(p,q)]}\}_{(p,q)\in W}$.
+Multiplication is then truncated degree-convolution (Lemma 20.1), and the extractor is coordinate projection onto $E_{0}=\{(0,0),(0,1)\}$.
+Remark 21.15 (Algebraic view (no explicit “bands”)).
+Note that explicit “bands” can be abstracted away by working in a finite support graded representation and treating packed values as coefficient tables (degrees -11,0,++11) or the fixed multigraded window in the folded construction).
+From this perspective, the three-band identity is just degree-convolution, and “extraction” is the linear projection onto the target degree(s).
+The dyadic sieve implements this projection numerically on the dyadic lattice by enforcing a rounding gap.
+This clarifies that the kernel is algebraic (constant-width extension with a fixed projection), so classical MM-based algebraic reductions transfer at the arithmetic-operation level, while our earlier GPR reductions are needed to preserve bounded representations in the bit/word model.
 21.6 Alternative numeric domains
@@
 The framework is strongest when (1) the host algorithm is GEMM-dominant in a blocked/recursive form, and (2) a projection discipline exists that keeps the relevant bands uniformly bounded.
-The main improvements ahead are sharpening constants, enlarging feasible regimes under more structure assumptions, and exploring alternative extractor mechanisms with better headroom tradeoffs.
+Although the algorithmic structure presented here is sufficient to achieve the principal objective of rigorous and honest near-quadratic scaling, there is need to explore refactorizations of the algorithm to maximize practical performance.
+This includes relatively obvious space and computation reuse among the four recursion, and extends to restructured “foldings” to optimize the mapping of computations to hardware (serial and/or parallel).
 Appendix A Parallel bit-complexity and communication overhead
@@
 Appendix B Exact dyadic-lattice extractor specification (integral mode)
-This appendix gives a concrete, implementable specification of the dyadic-lattice rounding operator $\frac{1}{D^{2}}\left\lfloor D^{2}(\right\rceil\cdot)$ and the sieve extractor $\mathcal{S}_{\beta}(\cdot)$ used in Part I.
+This appendix gives a concrete, implementable specification of the dyadic-lattice rounding operator $\frac{1}{D^{2}}\left\lfloor D^{2}(\right\rceil\cdot)$ and the scaled sieve extractor $\mathcal{S}_{\beta,\sigma}(\cdot)$ used in Part I.
 The purpose is to eliminate any ambiguity about “exactness” in the integral semantics: all operations below are exact on the declared success set (master gap holds with slack), and any failure of the gap condition can be detected.
@@
 B.2 Exact nearest-integer rounding of dyadic rationals
-We use a deterministic nearest-integer rounding $\left\lfloor\cdot\right\rceil:\mathbb{R}\to\mathbb{Z}$ with a fixed tie-breaking rule.
-Because the paper’s extraction guarantees use strict slack, tie-breaking is irrelevant on the success set, but we specify it anyway for completeness (ties-to-even is acceptable).
-Definition B.2 (Exact rounding of an integer divided by a power of two).
+We use a deterministic nearest-integer rounding $\left\lfloor\cdot\right\rceil:\mathbb{R}\to\mathbb{Z}$ with ties-to-even tie-breaking.
+Because the paper’s extraction guarantees use strict slack, tie-breaking is irrelevant on the success set, but we specify it anyway for completeness.
+Definition B.2 (Exact rounding of an integer divided by a power of two (ties-to-even)).
 Let $N\in\mathbb{Z}$ and $s\in\mathbb{Z}_{\geq 0}$.
-Define $\mathrm{RoundPow2}(N,s):=\left\lfloor N/2^{s}\right\rceil$ computed with $N=\mathrm{sgn}\cdot A$, where $\mathrm{sgn}\in\{+1,-1\}$ and $A=\lvert N\rvert\in\mathbb{Z}_{\geq 0}$:
+Define $\mathrm{RoundPow2}(N,s):=\left\lfloor N/2^{s}\right\rceil$ using ties-to-even.
+Write $N=\mathrm{sgn}\cdot A$, where $\mathrm{sgn}\in\{+1,-1\}$ and $A=\lvert N\rvert\in\mathbb{Z}_{\geq 0}$.
+If $s=0$, return $\mathrm{RoundPow2}(N,0):=N$.
+Otherwise ($s\geq 1$), define:
 $q~:=~\left\lfloor\frac{A}{2^{s}}\right\rfloor,~r~:=~A-q\cdot 2^{s},$ (278)
-so $0\leq r<2^{s}$. Let $h:=2^{s-1}$ for $s\geq 1$ (and $h:=0$ for $s=0$). Define:
-$\tilde{q}~:=~\begin{cases}q,&r<h,\\ q+1,&r>h,\\ q~\text{or}~q+1~\text{(tie-breaking)},&r=h.\end{cases}$ (256)
+so $0\leq r<2^{s}$, and let $h:=2^{s-1}$. Define:
+$\tilde{q}~:=~\begin{cases}q,&r<h,\\ q+1,&r>h,\\ q,&r=h~~\text{and}~~q\equiv 0~(\mathrm{mod}~2),\\ q+1,&r=h~~\text{and}~~q\equiv 1~(\mathrm{mod}~2).\end{cases}$ (279)
 Return $\mathrm{RoundPow2}(N,s):=\mathrm{sgn}\cdot\tilde{q}$.
 Remark B.3.
-In code, this is implemented by bit shifts and bit masks on the absolute value (using a wider fixed-width type such as __int128 as needed), with a deterministic rule for the $r=h$ tie case.
+In code, this is implemented by bit shifts and bit masks on the absolute value (using a wider fixed-width type such as __int128 as needed), with a deterministic ties-to-even rule for the $r=h$ tie case.
 B.3 Exact coefficient rounding $\frac{1}{D^{2}}\left\lfloor D^{2}(\right\rceil\cdot)$
@@
 This is immediate from the representation $D^{2}z=(Dw)2^{e}$ and exact rounding of dyadic rationals by Definition B.2. ∎
-B.4 Exact sieve extractor $\mathcal{S}_{\beta}(\cdot)$
+B.4 Exact scaled sieve extractor $\mathcal{S}_{\beta,\sigma}(\cdot)$
 We now specialize to bases $\beta$ that are powers of two, which makes division by $\beta$ an exponent-offset update.
@@
 Let $\sigma\geq 0$ be the node-local coefficient-lattice exponent at the current recursion node (Definition 3.2; Eq.
-266).
+289).
 For $z\in\mathcal{L}_{D}$, define:
@@
 All steps are exact in $\mathcal{L}_{D}$ (i.e., dividing by $\beta=2^{b}$ is an exponent-offset update) and $\frac{1}{D^{2}}\left\lfloor D^{2}(\right\rceil\cdot)$ is exact by the preceding lemma.
-Therefore the implemented $\mathcal{S}_{\beta}$ satisfies the mathematical sieve definition.
+Therefore the implemented $\mathcal{S}_{\beta,\sigma}$ satisfies the mathematical sieve definition.
 ∎
@@
 $\delta_{0}~:=~2^{-3}=1/8,$ (285)
-and we do not treat $\delta$ as a tuning parameter. At every recursion node we require the sufficient-base inequality from Theorem  7.1 with $\delta=\delta_{0}$:
+and we do not treat $\delta$ as a tuning parameter. At every recursion node we require the sufficient-base inequality from Theorem 7.1 with $\delta=\delta_{0}$:
 $D^{2}2^{\sigma}\left(\frac{2}{\beta}+\frac{1}{\beta^{2}}\right)~\leq~\frac{1}{2}-\delta_{0},~~\beta=2^{b},~b\in\mathbb{Z}_{\geq 1}.$ (286)
 When Eq.
-263 holds and the normalization bounds $\|A^{\prime}_{i,:}\|_{2}\leq 1$, $\|B^{\prime}_{:,j}\|_{2}\leq 1$ are satisfied, extraction is exact because the master gap in Lemma  5.3 is open by at least $\delta_{0}$.
+286 holds and the normalization bounds $\|A^{\prime}_{i,:}\|_{2}\leq 1$, $\|B^{\prime}_{:,j}\|_{2}\leq 1$ are satisfied, extraction is exact because the master gap in Lemma 5.3 is open by at least $\delta_{0}$.
 Tight node-local coefficient lattice exponent $\sigma$ (no over-estimation).
@@
 and let:
-$g_{\min}$ $:=$ $\min_{r:a_{r}<\infty,~b_{r}<\infty}(a_{r}+b_{r}),$ (266)
+$g_{T}$ $:=$ $\min_{r:a_{r}<\infty,~b_{r}<\infty}(a_{r}+b_{r}),$ (289)
+$g_{U}$ $:=$ $\min_{r\in\{1,\dots,m/2\}:a_{r}<\infty,~b_{r+m/2}<\infty}(a_{r}+b_{r+m/2}),$ (290)
+$g_{\min}$ $:=$ $\min\{g_{T},~g_{U}\},$ (291)
 $\sigma$ $:=$ $\max\{0,~-g_{\min}\}.$ (292)
-with the convention that if the set $\{r:a_{r}<\infty,~b_{r}<\infty\}$ is empty (i.e., every dot-product term is zero) then $\sigma:=0$.
-This ensures that every dot-product term $A^{\prime}_{ir}B^{\prime}_{rj}$ lies in the scaled coefficient lattice:
+with the conventions that if an index set is empty we interpret the corresponding minimum as $\infty$, and if both minima are $\infty$ (i.e., all dot-product terms contributing to the middle band $T$ and the upper band $U$ vanish) then we set $\sigma:=0$.
+This ensures that every dot-product term contributing to $T$ and $U$ under the local packing (i.e., $A^{\prime}_{ir}B^{\prime}_{rj}$ and $A^{\prime}_{ir}B^{\prime}_{r+m/2,j}$) lies in the scaled coefficient lattice:
 $\mathcal{L}_{\mathrm{coeff}}(\sigma)~=~D^{-2}2^{-\sigma}\mathbb{Z}$ (293)
@@
 296 fails even after the one-step upgrade), then the implementation must not apply the sieve at that node.
-Instead it must treat this as leaving the declared exact-execution regime (e.g., due to overflow, truncation below lattice resolution, or a bug) and must fall back to a conventional exact multiplication on the original subproblem, i.e., call ClassicalMM(A,B)(A,B) (Algorithm  3) and return its result.
-(An optimized implementation may instead increase the staged-precision budget / word width and restart the node; see Algorithm  2 and Theorem  9.2.) With this policy, the algorithm never returns an incorrect extracted value, i.e., failure affects only performance, not correctness.
+Instead it must treat this as leaving the declared exact-execution regime (e.g., due to overflow, truncation below lattice resolution, or a bug) and must fall back to a conventional exact multiplication on the original subproblem, i.e., call ClassicalMM(A,B)(A,B) (Algorithm 4) and return its result.
+(An optimized implementation may instead increase the staged-precision budget / word width and restart the node; see Algorithm 2 and Theorem 9.2.) With this policy, the algorithm never returns an incorrect extracted value, i.e., failure affects only performance, not correctness.
 Explicit fallback routine (ClassicalMM).
@@
 A reference implementation may use a wider accumulator (or arbitrary-precision integers for the mantissas) for the inner sum, because this path is only a correctness backstop and is outside the near-quadratic performance model.
-Input: $A,B\in\mathcal{L}_{D}^{m\times m}$
-Output: $C=AB$ exactly
+Algorithm 4 ClassicalMM(A,B)(A,B) (exact dyadic fallback)
+$A,B\in\mathcal{L}_{D}^{m\times m}$.
+$C=AB$ exactly.
 for $i\leftarrow 1$ to $m$ do
   for $j\leftarrow 1$ to $m$ do
-    $s\leftarrow 0$
+   $s\leftarrow 0$.
    for $r\leftarrow 1$ to $m$ do
-       $s\leftarrow s+A_{ir}B_{rj}$
-    $C_{ij}\leftarrow s$
-return $C$
-Algorithm 3 ClassicalMM(A,B)(A,B) (exact dyadic fallback)
+     $s\leftarrow s+A_{ir}B_{rj}$.
+   end for
+   $C_{ij}\leftarrow s$.
+  end for
+end for
+return $C$.
 Mandatory call-site (Algorithm 1).